### v1

# Turán number of complete multipartite graphs in multipartite graphs

Abstract.
In this paper we study a multi-partite version of the Erdős–Stone theorem. Given integers $r<k$ and $t\geq 1$, let $\operatorname{ex}_{k}(n,K_{r+1}(t))$ be the maximum number of edges of $K_{r+1}(t)$-free $k$-partite graphs with $n$ vertices in each part, where $K_{r+1}(t)$ is the $t$-blowup of $K_{r+1}$. An easy consequence of the supersaturaion result gives that $\operatorname{ex}_{k}(n,K_{r+1}(t))=\operatorname{ex}_{k}(n,K_{r+1})+o(n^{2})$. Similar to a result of Erdős and Simonovits for the non-partite case, we find that the error term is closely related to the (multi-partite) Zarankiewicz problem. Using such Zarankiewicz numbers, for $t=2,3$, we determine the error term up to an additive linear term; using some natural assumptions on such Zarankiewicz numbers, we determine the error term up to an additive constant depending on $k$, $r$ and $t$. We actually obtain exact results in many cases, for example, when $k\equiv 0,1\pmod{r}$. Our proof uses the stability method and starts by proving a stability result for $K_{r+1}$-free multi-partite graphs.
Jie Han is partially supported by a Simons Collaboration Grant 630884 and National Natural Science Foundation of China (12371341). Yi Zhao is partially supported by NSF grant DMS 2300346 and Simons Collaboration Grant 710094.
1. Introduction
Generalizing Mantel’s theorem from 1907 [Mantel], Turán’s theorem from 1941 [MR0018405] started the systemetic study of Extremal Graph Theory. Given a graph $F$, let $\operatorname{ex}(n,F)$ denote the largest number of edges in a graph not containing $F$ as a subgraph (called $F$-free). Let $K_{r}$ denote the complete graph on $r$ vertices and $T_{r}(n)$ denote the complete $r$-partite graph on $n$ vertices with $\lfloor n/r\rfloor$ or $\lceil n/r\rceil$ in each part (known as the Turán graph); and $t_{r}(n)$ be the size of $T_{r}(n)$. Turán’s theorem [MR0018405] states that $\operatorname{ex}(n,K_{r+1})=t_{r}(n)$ for all $n\geq r\geq 1$ and in addition, $T_{r}(n)$ is the unique extremal graph.
Let $K_{t_{1},\dots,t_{r}}$ denote the complete $r$-partite graph with parts of size $t_{1},\dots,t_{r}$ and write $K_{r}(t)=K_{t,\dots,t}$ with $r$ parts. A celebrated result of Erdős and Stone [MR0018807] determines $\operatorname{ex}(n,K_{r+1}(t))$ asymptotically:
$\operatorname{ex}(n,K_{r+1}(t))=t_{r}(n)+o(n^{2})=\left(1-\frac{1}{r}\right)\frac{n^{2}}{2}+o(n^{2}).$
Erdős [MR0232703] and Simonovits [MR0233735] independently improved the error term above to $O(n^{2-1/t})$. Simonovits [MR0233735] also showed that any extremal graph for $K_{r+1}(t)$ can be obtained from $T_{r}(n)$ by adding or removing $O(n^{2-1/t})$ edges. Later Erdős and Simonovits [MR0292707] determined the structure of extremal graphs for $K_{r+1}(t)$ for $t\leq 3$ as follows.
Theorem 1.
[MR0292707] For $t\leq 3$, every extremal graph $G$ for $K_{r+1}(t)$ has a vertex partition $U_{1},\dots,U_{r}$ such that
  • $G[U_{i},U_{j}]$ is complete for all $i\neq j$,
  • $G[U_{i}]=n/r+o(n)$,
  • $G[U_{1}]$ is extremal for $K_{t,t}$, and
  • $G[U_{2}],\dots,G[U_{r}]$ are extremal for $K_{1,t}$.
The restriction $t\leq 3$ in Theorem 1 comes from our knowledge on $\operatorname{ex}(n,K_{t,t})$. A well-known open problem in Extremal Graph Theory is proving $\operatorname{ex}(n,K_{t,t})=\Omega(n^{2-1/t})$ and this is only known for $t\leq 3$.
Extremal problems whose host graphs are multipartite graphs have been studied since 1951, when Zarankiewicz proposed the study of the largest number of edges in a bipartite graph not containing a copy of $K_{s,t}$. Let $\mathcal{G}(n_{1},\dots,n_{k})$ denote the family of $k$-partite graphs with $n_{1},\dots,n_{k}$ vertices in its parts and write $\mathcal{G}_{k}(n)=\mathcal{G}(n,\dots,n)$ with $k$ parts. Given a graph $F$, define $\operatorname{ex}(n_{1},\dots,n_{k};F)$ as the largest number of edges in $F$-free graphs from $\mathcal{G}(n_{1},\dots,n_{k})$, and let $\operatorname{ex}_{k}(n,F)=\operatorname{ex}(n,\dots,n;F)$ (with $k$ parts). (Trivially $\operatorname{ex}_{k}(n,F)=\binom{k}{2}n^{2}$ if the chromatic number $\chi(F)>k$.) In 1975 Bollobás, Erdős, and Szemerédi [MR0389639] investigated several Turán-type problem for multipartite graphs. Applying a simple counting argument, they showed that
$\operatorname{ex}_{k}(n,K_{r+1})=t_{r}(k)n^{2}$ (1.1)
for any $n,k,r\in\mathbb{N}$ with $k\geq r$. The main results of [MR0389639] concern the minimum degree version of this problem, which has been intensively studied, see [MR2195582, MR1860440, MR1208805, MR4496024, MR2246152].
In this paper we study $\operatorname{ex}_{k}(n,K_{r+1}(t))$, the multi-partite versions of the Erdős–Stone theorem and Theorem 1.
We first prove a stability theorem for $\operatorname{ex}_{k}(n,K_{r+1})$. This result was independently obtained by Chen, Lu, and Yuan [CLY23]. Given $r,k\in\mathbb{N}$ with $k\geq r$, write $k=ar+b$ for $10\leq b\leq r-$ 1. By Turán’s theorem, the Turán graph $T_{r}(k)=K_{a,\dots,a,a+1,\dots,a+1}$ (with $b$ parts of size $a+1$ and $r-b$ parts of size $a$) is the unique largest $K_{r+1}$-free graph on $k$ vertices. We now describe extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$. Let ${\mathcal{T}}_{r,k}(n)$ be the collection of $k$-partite graphs with parts $V_{1},\dots,V_{k}$ of size $n$ defined as follows. If $b>0$, we arbitrarily divide $V_{ar+1},\dots,V_{k}$ into $r$ sets $W_{1},\dots,W_{r}$ (some of them may be empty) such that each $W_{i}$ is a subset of $V_{j}$ for some $j$; if $b=0$, then let $W_{1},\dots,W_{r}$ be empty sets. Now let $T$ be the $r$-partite graph with parts $U_{1},\dots,U_{r}$ such that
$U_{i}=W_{i}\cup Z_{i},\text{ where }Z_{i}:=V_{(i-1)a+1}\cup\cdots\cup V_{ia},$
obtained from the complete $r$-partite graph $K(U_{1},\dots,U_{r})$ by removing edges between $W_{i}$ and $W_{i^{\prime}}$, $i\neq i^{\prime}$, whenever $W_{i},W_{i^{\prime}}\subseteq V_{j}$ for some $j$ (in other words, $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$). Since $T$ is $r$-partite, it is $K_{r+1}$-free. Let $U=\bigcup_{i\in[r]}U_{i}$ and $W=\bigcup_{i\in[r]}W_{i}$. Note that $e_{T}(U)=\binom{r}{2}a^{2}n^{2}$ and $e_{T}(W,T)=|W|(k-a-1)n=b(k-a-1)n^{2}$. Since $t_{r}(k)=\binom{r}{2}a^{2}+b(k-a-1)$, it follows that $e(T)=t_{r}(k)n^{2}$. By (1.1), $T$ is an extremal graph for $K_{r+1}$. Let ${\mathcal{T}}_{r,k}(n)$ be the collection of all such $T$.
Given two graphs $G,H\in\mathcal{G}_{k}(n)$ on the same parts $V_{1},\dots,V_{k}$, we say that $G$ and $H$ are $\gamma$-close if $|E(G)\triangle E(H)|\leq\gamma n^{2}$.
Theorem 2.
For any positive integers $r\leq k$ and any $\gamma>0$, there exist $\varepsilon>0$ and $n_{0}$ such that the following holds for every integer $n\geq n_{0}$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Then $G$ is $\gamma$-close to a member of ${\mathcal{T}}_{r,k}(n)$.
We now consider $\operatorname{ex}_{k}(n,K_{r+1}(t))$, We assume that $k\geq r+1$ because otherwise $\text{ex}_{k}(n,K_{r+1}(t))=\binom{k}{2}n^{2}$ trivially. Applying (1.1) and either the Regularity Lemma or the Graph Removal Lemma, one can easily derive the following Erdős–Stone theorem for multipartite graphs:
$\text{ex}_{k}(n,K_{r+1}(t))=t_{r}(k)n^{2}+o(n^{2}).$
Applying Theorem 2 and the Graph Removal Lemma, we derive the following stability result for $\operatorname{ex}_{k}(n,K_{r+1}(t))$, which handles the non-extremal case for $\operatorname{ex}_{k}(n,K_{r+1}(t))$.
Theorem 3.
For any $k,r,t\in\mathbb{N}$ and any $\gamma>0$, there exist $\varepsilon>0$ and $n_{0}\in\mathbb{N}$ such that the following holds for every integer $n\geq n_{0}$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}(t)$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Then $G$ is $\gamma$-close to a member of ${\mathcal{T}}_{r,k}(n)$. In particular, we have $e(G)\leq(t_{r}(k)+\gamma)n^{2}$.
To give the precise value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$, we need the following definitions. Given $a,t,n_{1},\dots,n_{a}\in\mathbb{N}$, let $z_{t}^{(a)}(n_{1},\dots,n_{a})$ be the $a$-partite Zarankiewicz number for $K_{t,t}$, namely, the maximum number of edges in a $K_{t,t}$-free $a$-partite graph with part sizes $n_{1},\dots,n_{a}$. We also write $z_{t}^{(a)}(n)$ for $z_{t}^{(a)}(n,\dots,n)$. Given $r,k,t,n\in\mathbb{N}$ with $k\geq r+1$, assume $k=ar+b$ for some $10\leq b\leq r-$ 1, in particular, $b\geq 1$ if $a=1$. Let
$t_{1}:=\left\lfloor\frac{t-1}{4a^{2}+6a}\right\rfloor,\quad\text{and}\quad b^{\prime}:=\min\left\{b-1,\left\lfloor\frac{r-b}{2}\right\rfloor\right\}.$
Note that $b^{\prime}=b$ if and only if $b\leq(r+2)/3$. The following two functions will appear in our lower and upper bounds for $\operatorname{ex}_{k}(n,K_{r+1}(t))$, respectively.
$h_{1}(n,r,k,t):=\begin{cases}\lfloor\frac{t-1}{2}(k-a-1)n\rfloor&\text{ if $a\geq 2$, $b\geq 2$ and $t_{1}=0$},\\ \lfloor\frac{t-1}{2}(k-a-1)n\rfloor+\min\{b-1,r-b\}\frac{(2a^{2}+3a)t_{1}^{2}}{2}-\frac{r-1}{2}&\text{ if $a\geq 2$, $b\geq 2$ and $t_{1}>0$},\\ (r-1)\lfloor\frac{t-1}{2}an\rfloor&\text{ if $a\geq 2$ and $b=0,1$},\\ (t-1)(b-1)n+b^{\prime}\lfloor\frac{(t-1)^{2}}{4}\rfloor&\text{ if $a=1$ and $b\geq 1$},\\ \end{cases}$
and
$h_{2}(n,r,k,t):=\begin{cases}\lfloor\frac{t-1}{2}(k-a-1)n\rfloor+(r-b)\frac{(t-1)^{2}}{16(a-1)}&\text{ if }a\geq 2,t\geq a,\\ \lfloor\frac{t-1}{2}(k-a-1)n\rfloor&\text{ if }a\geq 2,t<a,\\ (r-1)\lfloor\frac{t-1}{2}an\rfloor&\text{ if $a\geq 2$ and $b=0,1$},\\ (t-1)(b-1)n+(b-1)\lfloor\frac{(t-1)^{2}}{4}\rfloor&\text{ if $a=1$ and $b\geq 1$}.\end{cases}$
Note that $h_{1}(n,r,k,t)=h_{2}(n,r,k,t)$ in the following cases:
$(1)\ a\geq 2\text{ and }t<a,\quad(2)\ a\geq 2\text{ and }b=0,1,\quad(3)\ a=1\text{ and }1\leq b\leq(r+2)/3.$ (1.2)
For $i=1,2$, we accordingly define
$g_{i}(n,r,k,t):=t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+h_{i}(n,r,k,t),$
and note that $\lceil k/r\rceil$ is equal to $a+1$ if $b>0$ and $a$ otherwise.
We have the following lower bound for $\text{ex}_{k}(n,K_{r+1}(t))$.
Theorem 4.
Given $r,k,t,n\in\mathbb{N}$, we have $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq g_{1}(n,r,k,t)$.
The second terms in the second and the fourth cases of the definition of $h_{1}(n,r,k,t)$ show that $\operatorname{ex}_{k}(n,K_{r+1}(t))>t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+\lfloor\frac{t-1}{2}(k-a-1)n\rfloor$ when $a,b\geq 2$ and $t_{1}>0$, and $\operatorname{ex}_{k}(n,K_{r+1}(t))>t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+(t-1)(b-1)n$ when $b\geq a=1$, and we shall elaborate on this at the end of this section.
Although we can use $z_{t}^{(a)}(n)$ without knowing its precise value in the lower bound, our proofs of the upper bounds need several estimates on it. Kővári, Sós, Turán [MR0065617] showed that $z_{t}^{(2)}(n)=O(n^{2-1/t})$ for $t\geq 2$ and proving a matching lower bound is a well-known open problem:
  1. ($Z$)
    $z_{t}^{(2)}(n)=\Omega(n^{2-1/t})$ for $t\geq 2$.
Note that this is known for $t=2,3$ [MR0200182, MR0223262]. In addition, we will need the following properties.
  1. ($E1$)
    There exists $\delta>0$ such that for large $n$, $z_{t}^{(a+1)}(n)-z_{t}^{(a)}(n)\geq{\delta n^{2-1/t}}.$
  2. ($E2$)
    for any $\varepsilon\in(0,1]$ and integer $a\geq 2$, there exists $\delta>0$ such that for large $n$,
    $z_{t}^{(a)}(n)-z_{t}^{(a)}((1-\varepsilon)n,n,\dots,n)\geq{\delta n^{2-1/t}}.$
  3. ($E3$)
    $z_{t}^{(a)}(n_{1},\dots,n_{a})-z_{t}^{(a)}(n_{1}-1,\dots,n_{a})\gg 1$. That is, for any constant $C^{*}$, there exists $n_{0}\in\mathbb{N}$ such that the $z_{t}^{(a)}(n_{1},\dots,n_{a})-z_{t}^{(a)}(n_{1}-1,\dots,n_{a})\geq C^{*}$ whenever $n_{1},\dots,n_{a}\geq n_{0}$.
($E1$) is a special case of ($E2$) with $\varepsilon=1$. However, we distinguish them because we can indeed derive ($E1$) from ($Z$).
Theorem 5.
Suppose ($Z$) holds. For any $r,k,t\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ is an integer and $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}(t)$-free, then $e(G)=t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+O(n)$.
Our loose estimate on $\operatorname{ex}_{k}(n,K_{r+1}(t))$ from Theorem 5 only assumes ($Z$) and thus holds for $t=2,3$.
Theorem 6.
Suppose ($Z$)($E2$) and ($E3$) hold. For any $r,k,t\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ is an integer and $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}(t)$-free, then $e(G)\leq g_{2}(n,r,k,t)$.
Since $h_{1}(n,r,k,t)=h_{2}(n,r,k,t)$ and $g_{1}(n,r,k,t)=g_{2}(n,r,k,t)$ hold under (1.2), Theorems 34 and 6 together determine $\text{ex}_{k}(n,K_{r+1}(t))$ in the cases of (1.2) for large $n$ and under the hypotheses ($Z$)($E2$) and ($E3$). In all other cases, the difference between our upper and lower bounds, $h_{2}(n,r,k,t)-h_{1}(n,r,k,t)$, is $O_{r,k}(t^{2})$.
Remark. Theorem 1 says that any extremal graph for $K_{r+1}(t)$ is the join of an extremal graph for $K_{t,t}$ and $r-1$ extremal graphs for $K_{1,t}$. Inspired by by this result, a natural guess of an extremal graph for our $r$-partite problem is starting from an $n$-blowup of $T_{r}(k)$, add a $K_{t,t}$-free $(a+1)$-partite graph to one of its parts with $a+1$ partition sets (assuming $b>0$), and $K_{1,t}$-free graphs to other $r-1$ parts. This graph has
$t_{r}(k)n^{2}+z_{t}^{a+1}(n)+\left\lfloor\frac{(k-a-1)(t-1)}{2}\right\rfloor$
edges. However, Theorem 4 says that this is not always extremal. Indeed, by moving vertices around, we obtain graphs with more edges and such graphs do not admit a partition similar to the ones in Theorem 1 (see the proof of Theorem 4, Section 3). This shows that our problem is different and more challenging than the non-partite version considered by Erdős and Simonovits [MR0292707].
Nevertheless, when $b=0$ this complexity does not exist, and we give an analog of Theorem 1.
Theorem 7.
For $r,k\in\mathbb{N}$ with $r\mid k$ and $t=2,3$, there exist $C_{0},n_{0}\in\mathbb{N}$ such that the following holds for $n\geq n_{0}$. Let $G$ be a $K_{r+1}(t)$-free $k$-partite graph with $n$ vertices in each part and $\operatorname{ex}_{k}(n,K_{r+1}(t))$ edges. Then there is a partition of the $k$ vertex clusters of $G$ into $r$ groups $U_{1},\dots,U_{r}$, each with $k/r$ clusters, and a vertex set $Z\subseteq V(G)$ with $|Z|\leq C_{0}$ such that
  • $G[U_{i}\setminus Z,U_{j}\setminus Z]$ is almost complete for all $i\neq j$,
  • $G[U_{1}\setminus Z]$ is $K_{t,t}$-free, and
  • for $i\in[2,r]$, $G[U_{i}\setminus Z]$ is $K_{1,t}$-free.
Assuming ($E2$), we can show that $Z=\emptyset$ in the theorem above. However, we choose to present a result that resembles Theorem 1 and requires no additional condition.
Organization. In Section 2 we prove the stability theorems (Theorems 2 and 3), that is, on $K_{r}$-free and $K_{r}(t)$-free multi-partite graphs. Then we prove the lower bound, Theorem 4, by giving the corresponding constructions in Section 3. Finally we prove the upper bounds, Theorems 56 and 7, together in Section 4.
Notation. We omit floors and ceilings unless they are crucial, e.g., we may choose a set of $\varepsilon n$ vertices even if our assumption does not guarantee that $\varepsilon n$ is an integer.
When $X,Y\subseteq V(G)$ intersect, $E_{G}(X,Y)$ is defined as the collection of ordered pairs in $(x,y)\in X\times Y$ such that $\{x,y\}\in E(G)$. Write $e_{G}(X,Y)=|E_{G}(X,Y)|$. For a vertex $v$ in $G$, let $N(v,X)=N(v)\cap X$ and $d(v,X)=|N(v,X)|$. Moreover, given $X\subseteq V(G)$, let $e(X,G)$ be the number of edges of $G$ incident to the vertices of $X$. Given two graphs $G$ and $H$ on a common vertex set $V$, $G\cap H$ denotes a graph on $V$ with $E(G\cap H)=E(G)\cap E(H)$. Given a $k$-partition $\{V_{1},V_{2},\dots,V_{k}\}$, a set $S$ is called crossing if $|S\cap V_{i}|\leq 1$, $i\in[k]$.
When we choose constants $x,y>0$, $x\ll y$ means that for any $y>0$ there exists $x_{0}>0$ such that for any $x<x_{0}$ the subsequent statement holds. Hierarchies of other lengths are defined similarly. Furthermore, all constants in the hierarchy are positive and for a constant appearing in the form $1/s$, we always mean to choose $s$ as an integer.
2. Proof of the Stability Theorems (Theorems 2 and 3)
The extremal problem $\operatorname{ex}_{k}(n,K_{r+1})$ (instead of its stability version) has a simple probabilistic proof. Indeed, let $G\in\mathcal{G}_{k}(n)$ with $e(G)>t_{r}(k)n^{2}$. Denote the parts of $V(G)$ by $V_{1},\dots,V_{k}$ and the densities of the bipartite graphs $G[V_{i},V_{j}]$ as $d_{ij}$, $i,j\in[k]$. For each $i\in[r]$, uniformly choose a random vertex $v_{i}\in V_{i}$ independent of other choices. Thus, the probability for $v_{i}v_{j}\in E(G)$ is precisely $d_{ij}$. Let $X$ be the number of edges spanned on $\{v_{1},\dots,v_{k}\}$, and note that $\mathbb{E}(X)=\sum d_{ij}=e(G)/n^{2}>t_{r}(k)$. This implies that there exists a choice of the $k$-set $\{v_{1},\dots,v_{k}\}$ which spans more than $t_{r}(k)$ edges. By the definition of $t_{r}(k)$, this $k$-set contains a copy of $K_{r+1}$.
Our proof starts with this simple argument – we obtain $\mathbb{E}(X)\geq t_{r}(k)-\varepsilon$ from $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Since $G$ is $K_{r+1}$-free, we know deterministically that $X\leq t_{r}(k)$, and thus by Markov’s inequality we obtain that almost all crossing $k$-sets span exactly $t_{r}(k)$ edges and thus are isomorphic to the Turán graph $T_{r}(k)$.
Proposition 2.1.
For any $r,k\in\mathbb{N}$ and $\varepsilon>0$, there exist $\gamma>0$ and $n_{0}\in\mathbb{N}$ such that the following holds for every $n\geq n_{0}$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Then for every $k^{\prime}\in[1,k]$, all but at most $\gamma n^{k^{\prime}}$ crossing $k^{\prime}$-sets $K$ in $G$ satisfy that all but at most $\gamma n^{k-k^{\prime}}$ crossing $k$-sets containing $K$ are isomorphic to $T_{r}(k)$.
Proof.
Denote the parts of $V(G)$ by $V_{1},\dots,V_{k}$ and the densities of the bipartite graphs $G[V_{i},V_{j}]$ as $d_{ij}$, $i,j\in[k]$. For $i\in[k]$, let $v_{i}$ be a uniformly random vertex in $V_{i}$. Then let $X$ be the number of edges spanned on $\{v_{1},...,v_{k}\}$ and note that $\mathbb{E}(X)=\sum d_{ij}=e(G)/n^{2}\geq t_{r}(k)-\varepsilon$. Since $G[\{v_{1},...,v_{k}\}]$ is $K_{r+1}$-free, we have $X\leq t_{r}(k)$ deterministically. Let $Y=t_{r}(k)-X$. Then $Y$ is a nonnegative random integer with $\mathbb{E}(Y)=t_{r}(k)-\mathbb{E}(X)\leq\varepsilon$. By Markov’s inequality, $\mathbb{P}(Y>0)\leq\mathbb{E}(Y)\leq\varepsilon$, which implies that $\mathbb{P}[X=t_{r}(k)]=\mathbb{P}(Y=0)\geq 1-\varepsilon$. Therefore, all but at most $\varepsilon n^{k}$ crossing $k$-sets in $G$ span a copy of $T_{r}(k)$. The proposition follows by counting and choosing $\gamma^{2}\geq\binom{k}{k^{\prime}}\varepsilon$. ∎
We also need the celebrated Graph Removal Lemma for cliques due to Erdős, Frankl and Rödl [EFR86].
Lemma 2.2 (Graph Removal Lemma).
For any $r\in\mathbb{N}$ and any $\varepsilon>0$, there exist $\beta>0$ and integer $n_{0}\in\mathbb{N}$ such that the following holds for every integer $n\geq n_{0}$. If $G$ is a graph with at most $\beta n^{r}$ copies of $K_{r}$, then it can be made $K_{r}$-free by removing at most $\varepsilon n^{2}$ edges.
A crossing copy of $T_{r}(k)$ in the $k$-partite graph $G$ gives us a partition of $[k]$ into $r$ parts of size $a$ or $a+1$, where each part corresponds to a color class of $T_{r}(k)$. Throughout the rest of this section, a partition of $[k]$ refers to such a partition.
Now we give our proof.
Proof of Theorem 2.
Write $k=ar+b$ with $10\leq b\leq r-$ 1. Without loss of generality, assume $\gamma\ll 1/r$. Choose constants
$1/n\ll\varepsilon:=\varepsilon_{k}\ll\cdots\ll\varepsilon_{ar}\ll\gamma^{\prime}\ll\gamma_{ar}\ll\cdots\ll\gamma_{k}=:\gamma\ll 1/r,1/k.$
We use induction on both $r\geq 1$ and $b\geq 0$. Let $G$ be a $k$-partite $K_{r+1}$-free graph with $e(G)\geq(t_{r}(k)-\varepsilon_{k})n^{2}$. The base case $r=1$ is trivial, and thus we assume $r\geq 2$.
Here is an outline of the proof. The base case $b=0$ (i.e., $k=ar$) is simple. By Proposition 2.1 and averaging, there is a partition of $[k]$ such that many crossing $k$-sets of $V(G)$ induce a copy of $K_{r}(a)$ under this partition. We call the partition sets under this partition rows. For every $i\in[r]$, we find a crossing $a(r-1)$-set $S_{i}$ intersecting all but the $i$-th row such that if we extend $S_{i}$ to crossing $ar$-sets, then almost all these $ar$-sets are isomorphic to $K_{r}(a)$. This implies that the $i$-th row (and thus every row) has $o(n^{2})$ edges. Together with $e(G)\geq(t_{r}(k)-\varepsilon_{k})n^{2}$, we conclude that $G$ is close to a blow-up of $K_{r}(a)$.
For the $b>0$ case, we first show that we may assume that every $V_{i}$ is incident to many edges (Claim 2.3). Indeed, if say, $V_{k}$ is incident to too few edges, then we obtain the structure of $G-V_{k}$ by induction hypothesis. Next we add the vertices of $V_{k}$ to the partition of $G-V_{k}$ joining $W_{i}$’s. A key observation is that if we tend to add a set $V_{k}^{i}\subseteq V_{k}$ of vertices to the $i$-th row, then (current) $|W_{i}|=o(n)$ and thus there are only $o(n^{2})$ edges in $G[W_{i},V_{k}^{i}]$ (as a result, adding the vertices of $V_{k}^{i}$ will not create too many edges inconsistent with the new partition). Finally we need to move things around to make sure that after the operation all parts have at most $n$ vertices.
Now we may assume that all parts are incident to a good amount of edges as in (2.1). Then for $a\geq 2$, by (2.1) and Proposition 2.1, we show that every part $V_{i}$ is in a group of $a$ parts which only induces $o(n^{2})$ edges as in ($\dagger$)2. Repeated applications of ($\dagger$)2 give a contradiction with  (2.1). The case $a=1$ is nevertheless more complicated. We first find a pair of parts, say $(V_{1},V_{2})$, which is not too dense (because $G$ is $K_{r+1}$-free). Together with Proposition 2.1, this implies that almost all crossing $(k-2)$-sets in $\prod_{3\leq i\leq k}V_{i}$ form a copy of $T^{\prime}:=K_{1,\dots,1,2,\dots,2}$, which has $b-1$ clusters of size 2 and $r-b$ clusters of size 1, see ($\ddagger$)2. Then we can show that $G^{\prime}:=G-(V_{1}\cup V_{2})$ is close to a member of ${\mathcal{T}}_{r-1,k-2}(n)$, due to the fact that $G^{\prime}$ has few copies of $K_{r}$ and $(t_{r-1}(k-2)-o(1))n^{2}$ edges. Indeed, by the removal lemma, $G^{\prime}$ has a subgraph $G^{\prime\prime}$ which is $K_{r}$-free and still has $(t_{r-1}(k-2)-o(1))n^{2}$ edges, which is close to a member of ${\mathcal{T}}_{r-1,k-2}(n)$ by induction, and so is $G^{\prime}$. The rest of the proof is to find an edge $uv\in V_{1}\times V_{2}$ such that $u$ and $v$ have large common neighborhood in each part $V_{i}$, $i\in[3,k]$, which leads to a copy of $K_{r-1}$ in its common neighborhood, and thus $K_{r+1}\subseteq G$, a contradiction. We find such $uv$ by dedicate counting and the fact that $G[V_{1},V_{2}]=pn^{2}$ where $p$ is bounded away from 1.
The base case. We first prove the base case $b=0$, that is, $k=ar$. In this case $t_{r}(k)=\binom{r}{2}a^{2}$ and ${\mathcal{T}}_{r,k}(n)$ contains a unique member $T_{r,k}(n)$ which is the complete $r$-partite graph with exactly $a$ parts in each color classes. If $a=1$, then $G\in\mathcal{G}_{r}(n)$ is a subgraph of $T_{r,k}(n)$ and $G$ is $\varepsilon_{k}$-close to $T_{r,k}(n)$ because $e(G)\geq(t_{r}(k)-\varepsilon_{k})n^{2}$.
Now assume $a\geq 2$. Since $e(G)\geq(t_{r}(k)-\varepsilon_{k})n^{2}$, by Proposition 2.1, almost all crossing $k$-sets induce a copy of $T^{*}:=K_{r}(a)$. Since there are $\frac{k!}{(a!)^{r}r!}$ partitions of $[k]$, by averaging, there is one partition such that at least $n^{k}/(2k!)$ crossing $k$-sets of $G$ induces $T^{*}$ under this partition. Without loss of generality, denote this partition as $[1,a]\cup[a+1,2a]\cup\cdots\cup[(r-1)a+1,ar]$. Now, for each $i\in[r]$, we shall choose a crossing $(a(r-1))$-set $S_{i}$ from
$Y_{i}:=\prod_{j\neq i}\prod_{j^{\prime}\in[a]}V_{(j-1)a+j^{\prime}}$
such that all but at most $\gamma^{\prime}n^{a}$ crossing $k$-sets containing $S_{i}$ induce $T^{*}$ (in particular, the $a$ new vertices would form an independent set). Indeed, this is possible as there are at least $(n^{k}/(2k!))/n^{a}=n^{a(r-1)}/(2k!)$ choices for crossing sets $S_{i}$ in $Y_{i}$ so that $G[S_{i}]$ is isomorphic to $K_{r-1}(a)$, and among them, by Proposition 2.1, at most $\gamma^{\prime}n^{a(r-1)}$ crossing sets violate the extension property.
For $i\in[r]$, let $U_{i}:=\bigcup_{j^{\prime}\in[a]}V_{(i-1)a+j^{\prime}}$ and we shall show that $G$ is $\gamma_{k}$-close to $K(U_{1},\dots,U_{r})$. We claim that $G[U_{i}]$ has at most $\binom{a}{2}\gamma^{\prime}n^{2}$ edges. Indeed, every edge of $G[U_{i}]$ together with $S_{i}$ can be extended to $n^{a-2}$ crossing $k$-sets not isomorphic to $T^{*}$. If $G[U_{i}]$ contains more than $\binom{a}{2}\gamma^{\prime}n^{2}$ edges, then there are more than $\binom{a}{2}\gamma^{\prime}n^{2}\cdot n^{a-2}/\binom{a}{2}=\gamma^{\prime}n^{a}$ crossing $k$-sets containing $S_{i}$ not isomorphic to $T^{*}$ (a crossing $k$-set contains at most $\binom{a}{2}$ edges in $G[U_{i}]$ and is thus counted at most $\binom{a}{2}$ times), contradicting the definition of $S_{i}$. Therefore, $\bigcup_{i\in[r]}G[U_{i}]$ contains at most $r\binom{a}{2}\gamma^{\prime}n^{2}$ edges. Furthermore, since $e(G)\geq(t_{r}(k)-\varepsilon_{k})n^{2}$, $G[U_{1},\dots,U_{r}]$ (as an $r$-partite graph) has at most
$\varepsilon_{k}n^{2}+r\binom{a}{2}\gamma^{\prime}n^{2}$
crossing non-edges. Overall, this yields that one can obtain $T_{r,k}(n)$ by altering at most $\varepsilon_{k}n^{2}+2r\binom{a}{2}\gamma^{\prime}n^{2}\leq\gamma_{ar}n^{2}$ edges from $G$ and we are done.
The general case. Now suppose $b>0$ and the result holds for $k-1$ with $(\varepsilon_{k-1},\gamma_{k-1})$ in place of $(\varepsilon,\gamma)$ and we will verify it for $k$. We first prove the following claim. Let $\varepsilon^{\prime}:=\varepsilon_{k-1}/2$.
Claim 2.3.
If $e(V_{i},G)\leq(k-1-a+\varepsilon^{\prime})n^{2}$ for some $i\in[k]$, then $G$ is $\gamma$-close to some member of ${\mathcal{T}}_{r,k}(n)$.
Proof.
Without loss of generality, assume $i=k$. By the assumption on $e(G)$, we have
$e(G-V_{k})=e(G)-e(V_{k},G)\geq(t_{r}(k)-\varepsilon)n^{2}-(k-1-a+\varepsilon^{\prime})n^{2}\geq(t_{r}(k-1)-\varepsilon_{k-1})n^{2}.$
Then by the induction hypothesis, $G-V_{k}$ is $\gamma_{k-1}$-close to a member $T\in{\mathcal{T}}_{r,k-1}(n)$. Since $G-V_{k}$ is $K_{r+1}$-free, $e(G-V_{k})\leq t_{r}(k-1)n^{2}$, and thus $e(V_{k},G)\geq(k-1-a-\varepsilon)n^{2}$. Recalling the definition of ${\mathcal{T}}_{r,k-1}(n)$, let us denote the color classes of $T$ by $U_{1},\dots,U_{r}$ and for $i\in[r]$ let $U_{i}=W_{i}\cup Z_{i}$, such that each $Z_{i}$ is a union of $a$ clusters of $V_{1},\dots,V_{k-1}$.
We first show the following claim.
  1. ($enumi$)
    Given any vertex $v\in V_{k}$, if there exists $i_{0}\in[r]$ such that $|N(v)\cap Z_{i}|\geq\gamma^{4}n$ for every $i\in[r]\setminus\{i_{0}\}$, then $|N(v)\cap U_{i_{0}}|<\gamma^{4}n$.
Indeed, otherwise let $X_{i_{0}}\subseteq N(v)\cap U_{i_{0}}$ of size $\gamma^{4}n$ and $X_{i}\subseteq N(v)\cap Z_{i}$ of size $\gamma^{4}n$ for every $i\in[r]\setminus\{i_{0}\}$. Note that $T[X_{i},X_{j}]$ for distinct $i,j\in[r]$ is a complete bipartite graph. Since $G-V_{k}$ is $\gamma_{k-1}$-close to $T$, it follows that $e(G[X_{i},X_{j}])\geq(\gamma^{4}n)^{2}-\gamma_{k-1}n^{2}$ for any distinct $i,j\in[r]$. As $\gamma_{k-1}\ll\gamma$, by the Turán result, we find a copy of $K_{r}$ in $\bigcup_{i\in[r]}X_{i}$, which forms a copy of $K_{r+1}$ together with $v$, a contradiction. In particular, ($enumi$) implies that $|N(v)\cap Z_{i_{0}}|\leq\gamma^{4}n$ for some $i_{0}\in[r]$. Since $v$ misses at least $(a-\gamma^{4})n$ vertices in $Z_{i_{0}}$ and $n-1$ vertices in $V_{k}$, we have $d_{G}(v)\leq(k-1-a+\gamma^{4})n$ for any $v\in V_{k}$.
Let $V_{k}^{\prime}$ be the set of vertices $v\in V_{k}$ such that $d_{G}(v)<(k-1-a-\gamma^{2})n$, then we have
$(k-1-a-\varepsilon)n^{2}\leq e(V_{k},G)<|V_{k}^{\prime}|(k-1-a-\gamma^{2})n+(n-|V_{k}^{\prime}|)(k-1-a+\gamma^{4})n$
and thus $(\gamma^{2}+\gamma^{4})n(n-|V_{k}^{\prime}|)\geq(\gamma^{2}-\varepsilon)n^{2}$, giving that
$|V_{k}^{\prime}|\leq\left(1-\frac{\gamma^{2}-\varepsilon}{\gamma^{2}+\gamma^{4}}\right)n=\frac{\gamma^{4}+\varepsilon}{\gamma^{2}+\gamma^{4}}n\leq\gamma^{2}n.$
For each $v\in V_{k}\setminus V_{k}^{\prime}$, since $d_{G}(v)\geq(k-1-a-\gamma^{2})n$, we have that $|N(v)\cap Z_{i}|\geq\gamma^{4}n$ for all but possibly one $i\in[r]$. However, by ($enumi$), there must be one such $i\in[r]$, that is, for $v$ there exists $i_{0}\in[r]$ such that $|N(v)\cap Z_{i}|\geq\gamma^{4}n$ for all $i\in[r]\setminus\{i_{0}\}$ and $|N(v)\cap U_{i_{0}}|\leq\gamma^{4}n$. This gives a partition of $V_{k}\setminus V_{k}^{\prime}$ to $r$ (possibly empty) sets $V_{k}^{1},\dots,V_{k}^{r}$. Furthermore, observe that if $|U_{i}|\geq(a+2\gamma^{2})n$ (that is, $|W_{i}|\geq 2\gamma^{2}n$), then $V_{k}^{i}=\emptyset$, because $v$ can not have $|U_{i}|-\gamma^{4}n>(a+\gamma^{2})n$ non-neighbors in $U_{i}$.
Now we define a new partition $U_{1}^{\prime},\dots,U_{r}^{\prime}$ as follows. If $|W_{i}|<2\gamma^{2}n$, then we find a $j\in[r]$ such that $W_{j}$ and $W_{i}$ are from the same cluster of $G$, add $W_{i}$ to $W_{j}$ and redefine $W_{i}:=V_{k}^{i}$. Denote the resulting partition as $U_{1}^{\prime},\dots,U_{r}^{\prime}$, where $U_{i}^{\prime}=Z_{i}\cup W_{i}$ and we add the vertices of $V_{k}^{\prime}$ to any $U_{i}^{\prime}$ so that it does not exceed $(a+1)n$ vertices. Let $T^{\prime}=K(U_{1}^{\prime},\dots,U_{r}^{\prime})\cap K(V_{1},\dots,V_{k})$ and we shall show that $T^{\prime}$ and $G$ are $\gamma$-close. Indeed, recall that $G-V_{k}$ is $\gamma_{k-1}$-close to $T$, and for every $v\in V_{k}^{i}$, $v$ is incident to at most $\gamma^{4}n$ edges within $Z_{i}$ and thus at least $(k-1-a-\gamma^{2}-\gamma^{4})n$ edges outside $Z_{i}$. Moreover, for the set $V_{k}^{\prime}$ and all $W_{i}$’s that have been moved, the union of them contains at most $r\cdot 2\gamma^{2}n$ vertices, and thus for these vertices we may need to edit at most $2r\gamma^{2}n\cdot kn=2kr\gamma^{2}n^{2}$ edges. Therefore, the distance of $T^{\prime}$ and $G$ is at most
$\gamma_{k-1}n^{2}+n\cdot(\gamma^{2}+2\gamma^{4})n+2kr\gamma^{2}n^{2}<\gamma n^{2},$
where the second term is on (the changes brought by) the vertices of $V_{k}\setminus V_{k}^{\prime}$. We are done. ∎
Choose a new constant $\varepsilon\ll\gamma^{\prime\prime}\ll\varepsilon^{\prime}$. By Claim 2.3, we may assume that for every $i\in[k]$,
$e(V_{i},G)\geq(k-1-a+\varepsilon^{\prime})n^{2}.$ (2.1)
Next we first deal with the easier case $a\geq 2$.
Case 1. $a\geq 2$. By (2.1), for every $i_{0}\in[k]$, take $v\in V_{i_{0}}$ such that $d(v)\geq(k-1-a+\varepsilon^{\prime})n$. Note that there exist $k-a$ sets $V_{i}$ such that $|N(v)\cap V_{i}|\geq(\varepsilon^{\prime}/a)n$ – otherwise there exist $a$ sets $V_{i}$ such that $|N(v)\cap V_{i}|<(\varepsilon^{\prime}/a)n$, yielding $d(v)<a(\varepsilon^{\prime}/a)n+(k-a-1)n$, a contradiction. Therefore there exists $I\subseteq[k]$ of size $k-a$ such that for any $j\in I$, $|N(v)\cap V_{j}|\geq(\varepsilon^{\prime}/a)n$. As $\varepsilon\ll\gamma^{\prime\prime}\ll\varepsilon^{\prime}$, by Proposition 2.1, there exists a crossing $(k-a)$-set $S$ in $\prod_{j\in I}(N(v)\cap V_{j})$ such that all but at most $\gamma^{\prime\prime}n^{a}$ crossing $k$-sets containing $S$ are isomorphic to $T^{*}:=K_{a,\dots,a,a+1,\dots,a+1}$. Since $S\subseteq N(v)$ and $G$ is $K_{r+1}$-free, $S$ must be $K_{r}$-free. Moreover, $S$ is an induced subgraph of $T^{*}$ with $k-a$ vertices, and we infer that $G[S]$ must be isomorphic to the graph obtained from $T^{*}$ with a color class of size $a$ removed. This yields that all but at most $\gamma^{\prime\prime}n^{a}$ crossing $a$-sets from $\prod_{j\notin I}V_{j}$ form independent sets. Thus, we conclude that for any $j,j^{\prime}\notin I$, $e_{G}(V_{j},V_{j}^{\prime})\leq\gamma^{\prime\prime}n^{2}$. Note that $i_{0}\notin I$, and thus we obtain
  1. ($\dagger$)
    for any $i\in[k]$, there exists $I^{\prime}\subseteq[k]$ of size $a$ such that $i\in I^{\prime}$ and for any $j,j^{\prime}\in I^{\prime}$, $e_{G}(V_{j},V_{j}^{\prime})\leq\gamma^{\prime\prime}n^{2}$.
By Proposition 2.1, there is a partition of $[k]$ that induces at least $n^{k}/(2k!)$ labelled copies of $T^{*}$. For two clusters $V_{i}$ and $V_{j}$ that are in different color classes, we have $e_{G}(V_{i},V_{j})\geq n^{2}/(2k!)$ because each edge in $V_{i}\times V_{j}$ is contained in at most $n^{k-2}$ labelled copies of $T^{*}$. Now fix a color class of size $a+1$. Without loss of generality, denote this color class by $V_{1},\dots,V_{a+1}$ and consider $V_{1}$. By ($\dagger$)2, there exists $I^{\prime}\subseteq[k]$ of size $a$ such that $1\in I^{\prime}$ and for any $j,j^{\prime}\in I^{\prime}$, $e_{G}(V_{j},V_{j}^{\prime})\leq\gamma^{\prime\prime}n^{2}$. Since clusters from different color classes have densities at least $1/(2k!)>\gamma^{\prime\prime}$, we conclude that $I^{\prime}\subseteq[a+1]$. Without loss of generality, suppose $[a+1]\setminus I^{\prime}=\{2\}$. Applying ($\dagger$)2 again, we obtain a set $I^{\prime\prime}\subseteq[a+1]$ such that $2\in I^{\prime\prime}$ and for any $j,j^{\prime}\in I^{\prime\prime}$, $e_{G}(V_{j},V_{j}^{\prime})\leq\gamma^{\prime\prime}n^{2}$. Note that any element $i\in I^{\prime}\cap I^{\prime\prime}$ satisfies that $e_{G}(V_{i},V_{j})\leq\gamma^{\prime\prime}n^{2}$ for every $j\in[a+1]$. However, this implies $e(V_{i},G)\leq(k-1-a+a\gamma^{\prime\prime})n^{2}$, contradicting (2.1) as $\gamma^{\prime\prime}\ll\varepsilon^{\prime}$.
Case 2. $a=1$. Choose new constants $\gamma^{\prime\prime}\ll\varepsilon^{\prime\prime}\ll\beta\ll\varepsilon^{\prime}\leq 1-t_{r}(k)/\binom{k}{2}$. Since $G$ is $K_{r+1}$-free, we have $\sum_{i\neq j}d(V_{i},V_{j})/\binom{k}{2}\leq t_{r}(k)/\binom{k}{2}\leq 1-\varepsilon^{\prime}$. Fix a pair of clusters, say, $(V_{1},V_{2})$ with density $p\leq 1-\varepsilon^{\prime}$. Since there are more than $\gamma^{\prime\prime}n^{2}$ non-edges in $V_{1}\times V_{2}$, by Proposition 2.1, there exists a non-edge in $V_{1}\times V_{2}$ such that all but at most $\gamma^{\prime\prime}n^{k-2}$ crossing $k$-sets containing this pair form a copy of $T^{*}:=K_{1,\dots,1,2,\dots,2}$. That is,
  1. ($\ddagger$)
    all but at most $\gamma^{\prime\prime}n^{k-2}$ crossing $(k-2)$-sets in $\prod_{3\leq i\leq k}V_{i}$ form a copy of $T^{\prime}:=K_{1,\dots,1,2,\dots,2}$, which has $b-1$ clusters of size 2 and $r-b$ clusters of size 1.
We now show that $G^{\prime}:=G[\bigcup_{3\leq i\leq k}V_{i}]$ is close to a member of ${\mathcal{T}}_{r-1,k-2}(n)$. We may assume $k\geq 4$. Note the $e(G^{\prime})n^{k-4}=\sum_{X}e(X)$, where the sum is over all crossing $(k-2)$-sets $X$ in $\prod_{3\leq i\leq k}V_{i}$. By ($\ddagger$)2, this sum is at least $(n^{k-2}-\gamma^{\prime\prime}n^{k-2})e(T^{\prime})=(1-\gamma^{\prime\prime})n^{k-2}t_{r-1}(k-2)$. Hence,
$e(G^{\prime})\geq t_{r-1}(k-2)(1-\gamma^{\prime\prime})n^{2}.$ (2.2)
Moreover, since $T^{\prime}$ is $K_{r}$-free, ($\ddagger$)2 implies that the number of copies of $K_{r}$ in $G^{\prime}$ is at most $\binom{k-2}{r}\gamma^{\prime\prime}n^{r}$ because every copy of $K_{r}$ in $G^{\prime}$ is contained in $n^{k-2-r}$ crossing $(k-2)$-sets not isomorphic to $T^{\prime}$ and every crossing $(k-2)$-set not isomorphic to $T^{\prime}$ contains at most $\binom{k-2}{r}$ copies of $K_{r}$. The graph removal lemma (Lemma 2.2) implies that there exists $G^{\prime\prime}\subseteq G^{\prime}$ on $V(G^{\prime})$ such that $G^{\prime\prime}$ is $K_{r}$-free and $e(G^{\prime})-e(G^{\prime\prime})\leq\varepsilon^{\prime\prime}n^{2}$. By (2.2) and the assumption $\gamma^{\prime\prime}\ll\varepsilon^{\prime\prime}$, we have
$e(G^{\prime\prime})\geq e(G^{\prime})-\varepsilon^{\prime\prime}n^{2}\geq t_{r-1}(k-2)(1-\gamma^{\prime\prime})n^{2}-\varepsilon^{\prime\prime}n^{2}\geq(1-\varepsilon^{\prime\prime})t_{r-1}(k-2)n^{2}.$
By the induction hypothesis, we derive that $G^{\prime\prime}$ is $\beta$-close to some graph $T_{0}^{\prime}\in{\mathcal{T}}_{r-1,k-2}(n)$. This implies that $G^{\prime}$ is $(2\beta)$-close to this $T_{0}^{\prime}$.
Note that $e(V_{1}\cup V_{2},G)=e(G)-e(G^{\prime})$ and $e(G^{\prime})\leq e(G^{\prime\prime})+\varepsilon^{\prime\prime}n^{2}\leq(t_{r-1}(k-2)+\varepsilon^{\prime\prime})n^{2}$. Thus, we have
$e(V_{1}\cup V_{2},G)\geq(t_{r}(k)-\varepsilon)n^{2}-(t_{r-1}(k-2)+\varepsilon^{\prime\prime})n^{2}\geq(2(k-2)-2\varepsilon^{\prime\prime})n^{2}.$
For $i\in[k]$, let $c_{i}:=e(V_{i},G)/n^{2}$ and recall that $p=e(V_{1},V_{2})/n^{2}$. Then we have $2(t_{r}(k)-\varepsilon)\leq\sum_{i\in[k]}c_{i}\leq 2t_{r}(k)$ and $c_{1}+c_{2}-p\geq 2(k-2)-2\varepsilon^{\prime\prime}$. Our ultimate goal is finding an edge $uv\in V_{1}\times V_{2}$ such that $|N(u)\cap N(v)\cap V_{i}|\geq 2\sqrt{\beta}n$ for every $i\in[3,k]$, which allows us to find a copy of $K_{r-1}$ in $N(u)\cap N(v)$, contradicting that $G$ is $K_{r+1}$-free. Indeed, if there is an edge $uv\in V_{1}\times V_{2}$ such that
$d(u,V_{3}\cup\cdots\cup V_{k})+d(v,V_{3}\cup\cdots\cup V_{k})\geq(c_{1}+c_{2}-p-1+3\sqrt{\beta})n\geq(2k-5+2\sqrt{\beta})n,$ (2.3)
then for every $i\in[3,k]$, there exists $B_{i}\subseteq N(u)\cap N(v)\cap V_{i}$ of size $2\sqrt{\beta}n$. Recall that $G^{\prime}$ is $(2\beta)$-close to a member $T_{0}^{\prime}$ of ${\mathcal{T}}_{r-1,k-2}(n)$. In particular, there exist $r-1$ $B_{i}$’s such that $T_{0}^{\prime}$ induced on the union of these sets forms a copy of $K_{r-1}{(2\sqrt{\beta}n)}$. Thus, to destroy all these copies of $K_{r-1}$, one needs to remove at least $(2\sqrt{\beta}n)^{2}=4\beta n^{2}$ edges from $T_{0}^{\prime}$, while $e(T_{0}^{\prime})-e(G^{\prime})\leq 2\beta n^{2}$. Therefore $G[B_{3}\cup\cdots\cup B_{k}]$ contains a copy of $K_{r-1}$, which together with $uv$ forms a copy of $K_{r+1}$, a contradiction.
To find such $uv$, we first show that most vertices in $V_{i}$, $i\in[k]$ have large degrees. Fix $i\in[k]$ and let $V_{i}^{\prime}$ be the number of vertices $v\in V_{i}$ with $d(v)<c_{i}n-2k\gamma^{\prime\prime}n$. We claim that $|V_{i}^{\prime}|\leq\gamma^{\prime\prime}n/k$. Indeed, for any $j\in[k]$, since $e(V_{j},G)=c_{j}n^{2}$, there exists a set $V_{j}^{*}$ of size $\gamma^{\prime\prime}n/k$ such that $d(v)\leq(1+2\gamma^{\prime\prime}/k)c_{j}n$ for every $v\in V_{j}^{*}$. Now consider the subgraph $G^{*}:=G-(V_{i}^{\prime\prime}\cup\bigcup_{j\neq i}V_{j}^{*})$, where $V_{i}^{\prime\prime}$ is a subset of $V_{i}^{\prime}$ of size $\gamma^{\prime\prime}n/k$. Since $G^{*}$ is $K_{r+1}$-free, we have $e(G^{*})\leq t_{r}(k)(1-\gamma^{\prime\prime}/k)^{2}n^{2}$. By $c_{i}\leq k-1$, $j\in[k]$ and $\sum_{i\in[k]}c_{i}\leq 2t_{r}(k)$, we get
$e(G)$ $\leq t_{r}(k)(1-\gamma^{\prime\prime}/k)^{2}n^{2}+(\gamma^{\prime\prime}n/k)\left(c_{i}n-2k\gamma^{\prime\prime}n+\sum_{j\neq i}(1+2\gamma^{\prime\prime}/k)c_{j}n\right)$
$\leq t_{r}(k)(1-\gamma^{\prime\prime}/k)^{2}n^{2}+(\gamma^{\prime\prime}/k)2t_{r}(k)n^{2}-2\gamma^{\prime\prime}n^{2}\leq t_{r}(k)n^{2}-\gamma^{\prime\prime}n^{2},$
contradicting that $e(G)>t_{r}(k)n^{2}-\varepsilon n^{2}$.
Pick a vertex $u\in V_{1}\setminus V_{1}^{\prime}$ such that $d(u,V_{2})\leq(1+\gamma^{\prime\prime})pn$. Such a vertex exists – otherwise by $|V_{1}^{\prime}|\leq\gamma^{\prime\prime}n/k$ we have $e_{G}(V_{1},V_{2})\geq(1-\gamma^{\prime\prime}/k)n\cdot(1+\gamma^{\prime\prime})pn>pn^{2}$ contradicting our assumption. Consider $R:=N(u,V_{2}\setminus V_{2}^{\prime})$ and an arbitrary vertex $v\in R$. By the definition of $V_{i}^{\prime}$, we have that $d(u)+d(v)\geq c_{1}n+c_{2}n-4k\gamma^{\prime\prime}n$. We may assume that (2.3) fails (otherwise we are done). Together with $d(u)+d(v)\geq c_{1}n+c_{2}n-4k\gamma^{\prime\prime}n$, this implies that
$d(u,V_{2})+d(v,V_{1})\geq(1+p-4k\gamma^{\prime\prime}-3\sqrt{\beta})n.$
In particular, we obtain that $|R|=d(u,V_{2}\setminus V_{2}^{\prime})\geq(p-4k\gamma^{\prime\prime}-3\sqrt{\beta}-\gamma^{\prime\prime}/k)n\geq(p-4\sqrt{\beta})n$ and $d(v,V_{1})\geq(1-p\gamma^{\prime\prime}-4k\gamma^{\prime\prime}-3\sqrt{\beta})n>(1-4\sqrt{\beta})n$ (for every $v\in R$). This gives at least
$(p-4\sqrt{\beta})(1-4\sqrt{\beta})n^{2}\geq(p-8\sqrt{\beta})n^{2}$
edges in $G[V_{1},R]$. On the other hand, by (2.1), we have $c_{2}\geq k-2+\varepsilon^{\prime}$, and thus for every vertex $w\in V_{2}\setminus V_{2}^{\prime}$, we have
$d(w,V_{1})\geq c_{2}n-2k\gamma^{\prime\prime}n-(k-2)n\geq\varepsilon^{\prime}n/2.$
By $|R|\leq(1+\gamma^{\prime\prime})pn$ and $p<1-\varepsilon^{\prime}$, we get $|V_{2}\setminus(V_{2}^{\prime}\cup R)|\geq n-(1+\gamma^{\prime\prime})pn-\gamma^{\prime\prime}n/k\geq\varepsilon^{\prime}n/2$. Therefore, we have $e(G[V_{1},V_{2}\setminus(V_{2}^{\prime}\cup R)])\geq(\varepsilon^{\prime}n/2)^{2}$. However, combining these two estimates and recalling $\beta\ll\varepsilon^{\prime}$ we see that
$e(G[V_{1},V_{2}])\geq(p-8\sqrt{\beta})n^{2}+(\varepsilon^{\prime}/2)^{2}n^{2}>pn^{2},$
a contradiction. ∎
We now prove Theorem 3 by using the Graph Removal Lemma and a well-known result on supersaturation.
Lemma 2.4 (Supersaturation).
For any $r,t\in\mathbb{N}$ and any $\beta>0$, there exists $n_{0}\in\mathbb{N}$ such that the following holds for every integer $n\geq n_{0}$. If $G$ is a graph with $\beta n^{r}$ copies of $K_{r}$, then it contains a copy of $K_{r}(t)$.
Proof of Theorem 3.
Choose constants $1/n\ll\beta\ll\varepsilon\ll\gamma,1/k,1/r,1/t$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}(t)$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. We further assume that $G$ contains fewer than $\beta n^{r+1}$ copies of $K_{r+1}$ – otherwise Lemma 2.4 implies that $G$ contains a copy of $K_{r+1}(t)$, a contradiction. Thus, by Lemma 2.2, $G$ contains a $K_{r+1}$-free spanning subgraph $G^{\prime}$ with at least $e(G)-\varepsilon n^{2}\geq(t_{r}(k)-2\varepsilon)n^{2}$ edges. By Theorem 2, $G^{\prime}$ is $\gamma/2$-close to some $T\in{\mathcal{T}}_{r,k}(n)$. Consequently, $G$ is $\gamma$-close to $T$, as desired. ∎
3. Lower bound – proof of Theorem 4
In this section we prove Theorem 4. We start with the following proposition on $K_{2,2}$-free bipartite graphs.
Proposition 3.1.
Let $t\geq 1$ and $n>5t^{4}$. Suppose $G=(A,B;E)$ is a $K_{2,2}$-free bipartite graph with $n$ vertices in each part and with $\Delta(G)\leq t-1$. Then there is a perfect matching $M$ in the complement of $G$ such that $G\cup M$ is still $K_{2,2}$-free. In particular, there exists a $t$-regular $K_{2,2}$-free bipartite graph with $n$ vertices in each part.
Proof.
When we construct $M$, there are two types of $K_{2,2}$ that we need to avoid in $G\cup M$ – in the first type $M$ has an edge $uv$ where $G$ contains a path of length three with ends $u$ and $v$, and the second type consists of two edges of $M$ and two edges of $G$ on the same four vertices. The first one is easy to avoid as for any vertex $v$ in $G$, the number of vertices that can be reached from $v$ via a walk of length three (equivalently, a path of length one or three) is at most $(\Delta(G))^{3}\leq(t-1)^{3}$. Therefore, when matching a vertex $v\in A\cup B$ in $M$, we need to avoid at most $(t-1)^{3}$ vertices that do not depend on $M$. However, this is not the case for the second type of $K_{2,2}$ because the matching of a vertex in $M$ affects the matching (more precisely, the vertices that need to be avoided) of another vertex. To overcome this, we proceed in three phases.
Choose a set $S$ of arbitrary $2t^{3}$ vertices from $A$, and note that $N=\bigcup_{v\in S}N_{G}(v)$ has size at most $|S|t\leq 2t^{4}$. For the first phase of our process, we greedily find a matching $M_{0}$ that covers the vertices of $N$ and avoids the vertices of $S$. To achieve this, each time we match a vertex $v\in N$, we need to avoid the vertices of $S$, the vertices of $A$ that are already matched in $M_{0}$, and those vertices that have distance 1 or 3 from $v$ in the graph $G\cup M_{0}$. Since $\Delta(G\cup M_{0})\leq t$, we need to avoid at most $|S|+(|N|-1)+t^{3}<n$ vertices and thus this is possible. For the second phase, we greedily match all unmatched vertices $v\in A\setminus S$ and denote the resulting matching by $M_{1}$ (containing $M_{0}$). Similarly, each time we need to avoid the vertices that are already matched and the vertices that have distance 1 or 3 from $v$ in the graph $G\cup M_{1}$. This is possible as there are at most $(n-|S|-1)+t^{3}<n$ such vertices.
For the third phase, we are left with sets of unmatched vertices $S\subseteq A$ and $T\subseteq B$ such that $T\cap N(S)=\emptyset$. The key point is that $G[S,T]$ is an empty graph so no second type of $K_{2,2}$ will be created on $S\cup T$. For each vertex $v\in S\cup T$, we need to avoid the vertices that have distance 1 or 3 from $v$ in the graph $G\cup M_{1}$, and there are at most $(t-1)^{2}t<t^{3}=|S|/2$ such vertices. Therefore, we can choose a perfect matching on $S\cup T$ by Hall’s Marriage Theorem. This gives the desired perfect matching $M$.
The “in particular” part follows by starting with an empty bipartite graph and iteratively adding perfect matchings. ∎
Now we prove our lower bound on $\text{ex}_{k}(n,K_{r+1}(t))$.
Proof of Theorem 4.
We first deal with the case $a=1$. Let $V_{i,j}$, $(i,j)\in[r]\times[2]$ be vertex sets, where each of $V_{b+1,2},\dots,V_{r,2}$ are empty sets and all other sets have size $n$. Let $G:=K(V_{1,1}\cup V_{1,2},\dots,V_{r,1}\cup V_{r,2})$ be the blowup of the Turán graph $T_{r}(k)$. Thus $e(G)=t_{r}(k)n^{2}$.
We first revise the partition as follows. Let $t^{\prime}:=\lceil(t-1)/2\rceil$ and $b^{\prime}:=\min\{b-1,\lfloor(r-b)/2\rfloor\}$. Let $\{V_{i,j}^{\prime},(i,j)\in[r]\times[2]\}$ be obtained from $\bigcup V_{i,j}$ by moving $t^{\prime}$ vertices from $V_{i,1}$ to $V_{i+b-1,1}$, and moving $t^{\prime}$ vertices from $V_{i,2}$ to $V_{i+b+b^{\prime}-1,2}$, for every $i\in[2,b^{\prime}+1]$. For $i\in[r]$, let $U_{i}:=V_{i,1}^{\prime}\cup V_{i,2}^{\prime}$ and $H:=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$. Let $H^{\prime}$ be obtained from $H$ by adding
  • a $K_{t,t}$-free bipartite graph on $U_{1}$ of size $Z_{t}^{(2)}(n)$, and
  • a maximum $\{K_{1,t},K_{2,2}\}$-free bipartite graph on $U_{i}$ for $2\leq i\leq b+2b^{\prime}$.
Note that, for $i\in[2,b]$, Proposition 3.1 implies that each $H^{\prime}[U_{i}]$ is $(t-1)$-regular and thus for $i\in[2,b^{\prime}+1]$, $H^{\prime}[U_{i}]$ has $(n-t^{\prime})(t-1)$ edges, and for $i\in[b^{\prime}+2,b]$ it has $n(t-1)$ edges; for $i\in[b+1,b+2b^{\prime}]$, each $H^{\prime}[U_{i}]$ consists of $t^{\prime}$ vertex-disjoint stars $K_{1,t-1}$ centered at the $t^{\prime}$ new vertices.
We first observe that $H$ contains $t_{r}(k)n^{2}-b^{\prime}(t^{\prime})^{2}$ edges. Indeed, for every $i\in[2,b^{\prime}+1]$, the $t^{\prime}$ vertices moved from $V_{i,1}$ to $V_{i+b-1,1}$ lose $nt^{\prime}$ edges to $V_{i+b-1,1}$ and gain $(n-t^{\prime})t^{\prime}$ edges to $V_{i,2}$, thus having a net loss $(t^{\prime})^{2}$ edges between $U_{i}$ and $U_{i+b-1}$. The same holds for the $t^{\prime}$ vertices moved from $V_{i,2}$ to $V_{i+b+b^{\prime}-1,2}$. On the other hand, the $t^{\prime}$ new vertices in $V^{\prime}_{i+b-1,1}$ and $t^{\prime}$ new vertices in $V^{\prime}_{i+b+b^{\prime}-1,2}$ are all joined, giving additional $(t^{\prime})^{2}$ edges. Thus, the net loss of changing $G$ to $H$ is $b^{\prime}(t^{\prime})^{2}$ edges.
Furthermore, the vertices in $V_{2,2}\cup\cdots\cup V_{b,2}$ and the $b^{\prime}t^{\prime}$ vertices moved from $V_{2,1}\cup\cdots\cup V_{b^{\prime}+1,1}$ each have degree $t-1$ in $\bigcup_{2\leq i\leq r}H^{\prime}[U_{i}]$ and thus the number of edges in $H^{\prime}\setminus H$ is
$z_{t}^{(2)}(n)+(t-1)(b-1)n+(t-1)b^{\prime}t^{\prime}.$
Thus, we have $e(H^{\prime})=t_{r}(k)n^{2}+z_{t}^{(2)}(n)+(t-1)(b-1)n+b^{\prime}\lfloor\frac{(t-1)^{2}}{4}\rfloor=g_{1}(n,r,k,t)$.
At last, we observe that $H^{\prime}$ is $K_{r+1}(t)$-free. Indeed, by construction, every $U_{i}$ is triangle-free, $U_{1}$ is $K_{t,t}$-free and other $U_{i}$’s are $\{K_{1,t},K_{2,2}\}$-free. Therefore, a copy of $K_{r+1}(t)$ can contain at most $12t-$ 1 vertices from $U_{1}$ and at most $t$ vertices from other $U_{i}$’s, which is impossible.
Now we assume $a\geq 2$. Let $V_{i,j}$, $(i,j)\in[r]\times[a+1]$ be disjoint vertex sets, where $V_{i,a+1}=\emptyset$ for $b<i\leq r$, and all other sets have size $n$. Let $H$ be a graph defined as follows. Let
${t_{1}:=\left\lfloor\frac{t-1}{4a^{2}+6a}\right\rfloor\,\,\text{ and }\,\,t^{\prime}\in[t-at_{1}-1,t-at_{1}]\text{ be even}.}$
Let $\{V_{i,j}^{\prime},(i,j)\in[r]\times[a+1]\}$ be a partition of $\bigcup_{(i,j)\in[r]\times[a+1]}V_{i,j}$ obtained from moving $t_{1}$ vertices from each of $V_{i,j}$, $2\leq i\leq\min\{b,r+1-b\}$, $11\leq j\leq a+$ 1 to $V_{i+b-1,j}$. Denote by $Z_{i+b-1,j}$ the set of vertices moved from $V_{i,j}$ and let $Z_{i+b-1}:=\bigcup_{j=1}^{a+1}Z_{i+b-1,j}$. For $i\in[r]$, let $U_{i}:=\bigcup_{j\in[a+1]}V_{i,j}^{\prime}$. Let $H$ be obtained from $K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$ by adding graphs in each $U_{i}$ as follows.
  1. (S1)
    Add a maximum $K_{t,t}$-free $(a+1)$-partite graph in $U_{1}$.
  2. (S2)
    for $i\in[2,b]\cup[2b,r]$, [^1] When $2b>r$, the interval $[2b,r]$ is an empty set. add on $U_{i}$ a $\{K_{1,t},K_{2,2}\}$-free bipartite graph with $\lfloor|U_{i}|\frac{t-1}{2}\rfloor$ edges.
  3. (S3)
    for $i\in[b+1,\min\{2b-1,r\}]$, we add a triangle-free graph $H_{i}$ on $U_{i}$ such that $e(H_{i})\geq\lfloor|U_{i}|\frac{t-1}{2}\rfloor+\frac{t_{1}t^{\prime}}{2}$, and for any $j\in[a+1]$, $H_{i}-Z_{i,j}$ contains no biclique of order larger than $t$.
We first show that 1 is indeed possible. For $i\in[2,b]\cup[2b,r]$, each $U_{i}$ consists of $a$ or $a+1$ parts of equal size $n^{\prime}$, where $n^{\prime}=n$ or $n^{\prime}=n-t_{1}$. We first show that there exist disjoint sets $X_{1},\dots,X_{s},Y_{1},\dots,Y_{s}\subseteq U_{i}$, such that for each $j\in[s]$, $X_{j},Y_{j}$ belong to distinct parts of $U_{i}$, $|X_{j}|=|Y_{j}|\geq(n-t_{1})/2$ and $\bigcup_{i=1}^{s}(X_{i}\cup Y_{i})$ is either $U_{i}$ (if $|U_{i}|$ is even) or $U_{i}\setminus\{v\}$ for some $v\in V_{i,1}^{\prime}$ (if $|U_{i}|$ is odd). Indeed, this is trivial if $U_{i}$ has an even number of parts. If $U_{i}$ has an odd number of parts, since $a\geq 2$, there are at least three parts. We arbitrarily pair up all but the first three parts. Now if $n^{\prime}$ is even, then $|U_{i}|$ is even, and let $V_{i,1}^{\prime}=X_{1}\cup X_{2}$, $V_{i,2}^{\prime}=Y_{1}\cup X_{3}$ and $V_{i,3}^{\prime}=Y_{2}\cup Y_{3}$, where each small part has size $n^{\prime}/2$ and we are done. Otherwise both $n^{\prime}$ and $|U_{i}|$ are odd. In this case we take any $v\in V_{i,1}^{\prime}$, and let $V_{i,1}^{\prime}\setminus\{v\}=X_{1}\cup X_{2}$, $V_{i,2}^{\prime}=Y_{1}\cup X_{3}$ and $V_{i,3}^{\prime}=Y_{2}\cup Y_{3}$, where $|X_{1}|=|Y_{1}|=|X_{2}|=|Y_{2}|=(n^{\prime}-1)/2$ and $|X_{3}|=|Y_{3}|=(n^{\prime}+1)/2$.
Now we define the desired bipartite graph on $U_{i}$. If $|U_{i}|$ is even, then apply Proposition 3.1 to each pair $(X_{j},Y_{j})$. Since the resulting graph is $(t-1)$-regular and $K_{2,2}$-free, it satisfies the desired properties. Otherwise $|U_{i}|$ is odd, and we apply Proposition 3.1 and add $K_{2,2}$-free $(t-1)$-regular bipartite graphs on pairs of sets except $(X_{3},Y_{3})$ (as defined above). For $(X_{3},Y_{3})$ and $v\in V_{i,1}^{\prime}$, we connect $\lfloor(t-1)/2\rfloor$ edges from $v$ to each of $X_{3}$ and $Y_{3}$ and denote the set of these neighbors by $X^{*}$. Then by Proposition 3.1, we add a $(t-2)$-regular $K_{2,2}$-free bipartite graph on $(X_{3},Y_{3})$, and then add a matching of size $|X_{3}|-\lfloor(t-1)/2\rfloor$ such that the resulting graph is still $K_{2,2}$-free, all the vertices in $X_{3}\cup Y_{3}$ have degree $t-1$, and $X^{*}$ remains an independent set. Indeed, this is possible by applying Proposition 3.1 to the current graph with a complete bipartite graph added to $X^{*}$. By construction, the resulting graph is $K_{2,2}$-free, and all vertices except $v$ have degree $t-1$, while $v$ has degree $t-1$ or $t-2$. So it follows that the resulting graph has $\lfloor|U_{i}|\frac{t-1}{2}\rfloor$ edges as desired.
We now verify (S3). Recall that $t^{\prime}$ is even. Choose $A_{i,1}\subseteq V_{i,1}$ and $A_{i,2}\subseteq V_{i,2}$, each of size $t^{\prime}/2$. Add in $H_{i}$ a complete bipartite graph with parts $Z_{i}$ and $A_{i,1}\cup A_{i,2}$ (which is a copy of $K_{t^{\prime},(a+1)t_{1}}$). Next, if $a$ is even, then we add a $(t-1)$-regular $K_{2,2}$-free bipartite graph on each of $(V_{i,3},V_{i,4}),(V_{i,5},V_{i,6}),\dots,(V_{i,a-1},V_{i,a})$; if $a$ is odd, then we add a $(t-1)$-regular $K_{2,2}$-free bipartite graph on each of $(V_{i,4},V_{i,5}),(V_{i,6},V_{i,7}),\dots,(V_{i,a-1},V_{i,a})$. Thus, we are left with $V_{i,1},V_{i,2}$ and possibly $V_{i,3}$ when $a$ is odd. If $a$ is odd, then for $j=1,2$, take arbitrary $B_{i,j}\subseteq V_{i,j}\setminus A_{i,j}$ of size $\lceil n/3\rceil-t^{\prime}/2$; if $a$ is even, then let $B_{i,j}=V_{i,j}\setminus A_{i,j}$. We first connect vertices of $Z_{i}$ to $B_{i,1}\cup B_{i,2}$ such that every vertex of $Z_{i}$ has $t-1-t^{\prime}$ neighbors in $B_{i,1}\cup B_{i,2}$ and all these neighbors are distinct and distributed as evenly as possible in the two sets $B_{i,1},B_{i,2}$. In particular, if one of $B_{i,1},B_{i,2}$ receives one more edge than the other, then we remove one arbitrary such edge and denote the vertex of this edge in $Z_{i}$ by $w$. Denote by $B_{i,1}^{*}$ and $B_{i,2}^{*}$ the neighbors we have just obtained. So as for now for any vertex $v\in Z_{i}$, $N_{H_{i}}(v)\subseteq A_{i,1}\cup A_{i,2}\cup B_{i,1}^{*}\cup B_{i,2}^{*}$ and $d_{H_{i}}(v)\leq t-1$. Using Proposition 3.1, we add a $(t-1-at_{1})$-regular $K_{2,2}$-free bipartite graph $G^{\prime}$ on $(A_{i,1}\cup B_{i,1},A_{i,2}\cup B_{i,2})$ such that
  1. (A1)
    $G^{\prime}[A_{i,1}\cup B_{i,1}^{*},A_{i,2}\cup B_{i,2}^{*}]$ contains no edge and
  2. (A2)
    no two vertices in $A_{i,1}\cup A_{i,2}\cup B_{i,1}^{*}\cup B_{i,2}^{*}$ have any common neighbors in $U_{i}\setminus Z_{i}$.
Indeed, this can be done by applying Proposition 3.1 (repeatedly) on $(A_{i,1}\cup B_{i,1},A_{i,2}\cup B_{i,2})$ with the initial graph $K(A_{i,1}\cup B_{i,1}^{*},A_{i,2}\cup B_{i,2}^{*})$. Note that vertices in $A_{i,1}\cup A_{i,2}$ have degree exactly $t-1-at_{1}+(a+1)t_{1}=t-1+t_{1}$. Furthermore, we add a perfect matching on $(B_{i,1}\setminus B_{i,1}^{*},B_{i,2}\setminus B_{i,2}^{*})$ and $at_{1}-1$ edge-disjoint perfect matchings on $(B_{i,1},B_{i,2})$ such that i) they are edge-disjoint from $G^{\prime}$, ii) the resulting graph on $A_{i,1}\cup B_{i,1}\cup A_{i,2}\cup B_{i,2}$ is $K_{2,2}$-free and iii) (A1) and (A2) hold. This is possible by Proposition 3.1. By now the construction is completed if $a$ is even. If $a$ is odd, then for $j=1,2$, let $X_{i,j}\subseteq V_{i,j}\setminus(A_{i,j}\cup B_{i,j})$ of size $\lfloor n/2\rfloor$, and $X_{i,j}^{\prime}=V_{i,j}\setminus(A_{i,j}\cup B_{i,j}\cup X_{i,j})$. So $|X_{i,j}^{\prime}|\geq n/7$ is large. We split $V_{i,3}$ to $Y_{i,1},Y_{i,2}$, each of size $\lfloor n/2\rfloor$, and possibly one vertex $u$ (only when $n$ is odd). By Proposition 3.1, we add $(t-1)$-regular $K_{2,2}$-free bipartite graphs on each of $(X_{i,1},Y_{i,1})$ and $(X_{i,2},Y_{i,2})$. If $n$ is even, then by Proposition 3.1, we add a $(t-1)$-regular $K_{2,2}$-free bipartite graph on $(X_{i,1}^{\prime},X_{i,2}^{\prime})$ and the construction is finished. If $n$ is odd, then we connect $\lfloor(t-1)/2\rfloor$ edges from $u$ to each of $X_{i,1}^{\prime}$ and $X_{i,2}^{\prime}$ and connect $u$ and $w$ (if $w$ exists). Denote by $X^{**}$ the set of these neighbors of $u$ in $X_{i,1}^{\prime}$ and $X_{i,2}^{\prime}$. Next, similarly to above, by Proposition 3.1 we can add a $(t-2)$-regular $K_{2,2}$-free graph on $(X_{i,1}^{\prime},X_{i,2}^{\prime})$, and then add a matching of size $|X_{i,1}^{\prime}|-\lfloor(t-1)/2\rfloor$ to $(X_{i,1}^{\prime},X_{i,2}^{\prime})$ so that all vertices in $X_{i,1}^{\prime}\cup X_{i,2}^{\prime}$ have degree $t-1$, the resulting graph is $K_{2,2}$-free, and $X^{**}$ remains an independent set. The construction is completed. By construction,
  1. (1)
    vertices in $A_{i,1}\cup A_{i,2}$ have degree $t-1+t_{1}$ and all other vertices have degree $t-1$ with the exception of at most one vertex ($u$ or $w$ above), which has degree $t-2$.;
  2. (2)
    for any $j\in[a+1]$, $H_{i}-Z_{i,j}$ has maximum degree $t-1$ and thus is $K_{1,t}$-free;
  3. (3)
    $H_{i}$ is triangle-free;
  4. (4)
    $H_{i}-Z_{i}$ is $K_{2,2}$-free.
Now we verify the last assertion of (S3). Let $K$ be a biclique of order larger than $t$ in $H_{i}-Z_{i,j}$ for some $j\in[a+1]$. By (2) and (4), $V(K)$ must intersect $Z_{i}$. Then as $Z_{i}$ forms an independent set, $V(K)\cap Z_{i}$ is in one part of $K$, denoted by $P_{1}$, and by construction, the other part, denoted by $P_{2}$, is a subset of $A_{i,1}\cup A_{i,2}\cup B_{i,1}^{*}\cup B_{i,2}^{*}\cup\{u\}$. Since by (2) $H_{i}-Z_{i,j}$ is $K_{1,t}$-free, we have $|P_{1}|\geq 2$ and $|P_{2}|\geq 2$. By construction, for any $x\in A_{i,1}\cup A_{i,2}\cup B_{i,1}^{*}\cup B_{i,2}^{*}$, $N(x)\cap N(u)\cap U_{i}=\emptyset$. Therefore, together with (A2), the vertices of $P_{2}$ have no common neighbor outside $Z_{i}$, yielding $P_{1}\subseteq Z_{i}\setminus Z_{i,j}$ and $|P_{1}|\leq at_{1}$. Now, $P_{1}\subseteq Z_{i}$ and $|P_{1}|\geq 2$ imply that $P_{2}$ are the common neighbors of at least two vertices from $Z_{i}$, which gives that $P_{2}\subseteq A_{i,1}\cup A_{i,2}$ and thus $|P_{2}|\leq t^{\prime}$. Together we obtain $|V(K)|=|P_{1}|+|P_{2}|\leq at_{1}+t^{\prime}\leq t$, a contradiction. Therefore, $H_{i}-Z_{i,j}$ does not contain any bicliques on more than $t$ vertices.
At last, we compute $e(H_{i})$. Note that only vertices in $A_{i,1}\cup A_{i,2}$ have degree more than $t-1$ – they have degree $t-1+t_{1}$. All other vertices have degree $t-1$ with the exception of at most one vertex ($u$ or $w$ above), which has degree $t-2$. Since $t^{\prime}$ is even, $t_{1}t^{\prime}/2\in\mathbb{N}$. Thus, we have
$e(H_{i})=\left\lfloor\frac{|U_{i}|(t-1)+t_{1}t^{\prime}}{2}\right\rfloor=\left\lfloor|U_{i}|\frac{t-1}{2}\right\rfloor+\frac{t_{1}t^{\prime}}{2}.$
So (S3) is verified.
Claim 3.2.
The graph $H$ is $K_{r+1}(t)$-free and $e(H)\geq g_{1}(n,r,k,t)$.
Unlike the case $a=1$, we need some arguments to prove that $H$ is $K_{r+1}(t)$-free. The difference is that $H[U_{1}]$ might not be triangle-free (which is trivial for $a=1$, as $H[U_{1}]$ is bipartite), and thus it may contribute e.g. a copy of $K_{3}(t-1)$ towards a copy of $K_{r+1}(t)$.
Proof.
We first show that $H$ is $K_{r+1}(t)$-free. Suppose instead, that $H$ contains a copy of $K_{r+1}(t)$ on a vertex set $K$. We first claim the following.
  1. ($\heartsuit$)
    For every $i\in[2,r]$, $V(K)\cap U_{i}$ is a biclique or an independent set and $|V(K)\cap U_{i}|\leq 2t$. If $|V(K)\cap U_{i}|>t$, then $i\in[b+1,\min\{2b-1,r\}]$ and $V(K)\cap U_{i-(b-1)}=\emptyset$.
Indeed, for $i\in[2,r]$, since $H[U_{i}]$ is triangle-free, $V(K)\cap U_{i}$ is a biclique or an independent set. By (S3), if $|V(K)\cap U_{i}|>t$, then $i\in[b+1,\min\{2b-1,r\}]$ and $V(K)$ intersects all $Z_{i,j}$ for $j\in[a+1]$, which imply that $V(K)\cap U_{i-(b-1)}=\emptyset$. Furthermore, we know $|V(K)\cap(U_{i}\setminus Z_{i,1})|\leq t$ and thus $|V(K)\cap U_{i}|\leq t+|Z_{i,1}|<2t$. So ($\heartsuit$)3 is proved.
Let $K_{1}$ be the subgraph of $K$ induced on its vertices in $U_{1}$. By ($\heartsuit$)3, for each $i\in[2,r]$, we have either $|V(K)\cap U_{i}|\leq t$ or $|V(K)\cap(U_{i}\cup U_{i-b+1})|\leq 2t$. So we have $|V(K)\setminus V(K_{1})|\leq(r-1)t$ and thus $|V(K_{1})|\geq 2t$ (which is now possible as commented above).
Let $r^{\prime}$ be the number of members of $\mathcal{U}:=\{U_{i},i\in[2,r]\}$ that intersect $V(K)$. Since the number of members of $\mathcal{U}$ not intersecting $V(K)$ is $r-1-r^{\prime}$, by ($\heartsuit$)3, there are at most $r-1-r^{\prime}$ parts $U_{i}$, $i>1$ such that $|V(K)\cap U_{i}|>t$. Now we define an auxiliary graph $G^{*}$ on $C_{1},\dots,C_{r+1}$, the color classes of $K$. Indeed, for distinct $i,j\in[r+1]$, $C_{i}C_{j}\in E(G^{*})$ if and only if there exists $l\in[2,r]$ such that $C_{i}\cap U_{l}\neq\emptyset$ and $C_{j}\cap U_{l}\neq\emptyset$. Note that by ($\heartsuit$)3, each $U_{l}$ defines at most one edge of $G^{*}$. Let $\mathcal{C}_{1},\dots,\mathcal{C}_{s}$ be the connected components of $G^{*}$. Since $K_{1}$ is $K_{t,t}$-free, there is at most one $i\in[s]$ such that $V(\mathcal{C}_{i}):=\bigcup_{C_{j}\in\mathcal{C}_{i}}C_{j}$ contains at least $t$ vertices of $V(K_{1})$: two such components each containing $t$ vertices of $K_{1}$ will induce a copy of $K_{t,t}$ in $K_{1}$. Without loss of generality, assume that $|V(\mathcal{C}_{i})\cap V(K_{1})|<t$ for all $2\leq i\leq s$. For each $i\in[s]$, let $R_{i}$ be the set of indices $l\in[2,r]$ such that $|U_{l}\cap V(\mathcal{C}_{i})|>0$ and $R_{i}^{\prime}\subseteq R_{i}$ be the set of $l$ such that $|U_{l}\cap V(\mathcal{C}_{i})|>t$. So we have
$\sum_{i=1}^{s}|R_{i}|=r^{\prime}\quad\text{ and }\quad\sum_{i=1}^{s}|R_{i}^{\prime}|\leq r-1-r^{\prime}.$
Moreover, by the definition of $\mathcal{C}_{1},\dots,\mathcal{C}_{s}$, the $R_{i}$’s are pairwise disjoint. We claim that for $i\in[2,s]$, $|R_{i}|\geq|\mathcal{C}_{i}|-|R_{i}^{\prime}|$. Indeed, if $|R_{i}|\leq|\mathcal{C}_{i}|-|R_{i}^{\prime}|-1$, then by ($\heartsuit$)3, we have $|V(\mathcal{C}_{i})\cap\bigcup_{l\in R_{i}}U_{l}|\leq|R_{i}^{\prime}|\cdot 2t+(|\mathcal{C}_{i}|-2|R_{i}^{\prime}|-1)t=(|\mathcal{C}_{i}|-1)t$. Since $|V(\mathcal{C}_{i})\cap V(K_{1})|<t$, we obtain $|\mathcal{C}_{i}|t=|V(\mathcal{C}_{i})|<t+|V(\mathcal{C}_{i})\cap\bigcup_{l\in R_{i}}U_{l}|\leq|\mathcal{C}_{i}|t$, a contradiction. So the claim follows. Overall, we get
$\sum_{i=2}^{s}|R_{i}|\geq\sum_{i=2}^{s}(|\mathcal{C}_{i}|-|R_{i}^{\prime}|)\geq r+1-|\mathcal{C}_{1}|-((r-1-r^{\prime})-|R_{1}^{\prime}|)\geq r^{\prime}+2-|\mathcal{C}_{1}|.$
Therefore, $|R_{1}|\leq r^{\prime}-(r^{\prime}+2-|\mathcal{C}_{1}|)=|\mathcal{C}_{1}|-2$ and thus $G^{*}$ has at most $|\mathcal{C}_{1}|-2$ edges on $\mathcal{C}_{1}$. This contradicts with that $\mathcal{C}_{1}$ is connected. So $H$ is $K_{r+1}(t)$-free.
It remains to compute $e(H)$. First suppose $b=0,1$, then for $12\leq i\leq r-$ 1, $|U_{i}|=an$ and thus $H[U_{i}]$ has at least $\lfloor\frac{t-1}{2}an\rfloor$ edges. It follows that $e(H)=g_{1}(n,r,k,t)$. Secondly suppose $b\geq 2$ and let $b^{\prime}:=\min\{b-1,r-b\}$. Note that $e_{H}(U_{1})=z_{t}^{(\lceil k/r\rceil)}(n)$ and $t^{\prime}\geq t-at_{1}-1\geq(4a^{2}+5a)t_{1}$. So
$\sum_{i=2}^{r}e_{H}(U_{i})\geq\sum_{i=2}^{r}\left\lfloor|U_{i}|\frac{t-1}{2}\right\rfloor+b^{\prime}\frac{t_{1}t^{\prime}}{2}\geq\left\lfloor(k-a-1)n\frac{t-1}{2}\right\rfloor+b^{\prime}\frac{(4a^{2}+5a)t_{1}^{2}}{2}-\frac{r-1}{2}.$
Let $H^{\prime}=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$. It remains to bound $e(H^{\prime})$. Note that for each $i\in[b+1,b+b^{\prime}]$, we have moved a set $Z_{i}$ of $(a+1)t_{1}$ vertices from $\bigcup_{j}V_{i-b+1,j}$ to $U_{i}$. Thus, $H^{\prime}$ is obtained from $H^{*}:=K(U_{1}^{*},\dots,U_{r}^{*})\cap K(V_{1},\dots,V_{k})$ where $U_{i}^{*}=\bigcup_{j\in[a+1]}V_{i,j}$ by disconnecting the edges between $Z_{i}$ and $U_{i}^{*}$ and connecting the vertices of $Z_{i}$ to $U_{i-b+1}$ except for pairs of vertices from the same $V_{j}$’s ($Z_{i}$ remains to be an independent set). That is, for each vertex in $Z_{i}$ the net gain is $-an+(an-at_{1})=-at_{1}$, and thus the total net gain is $-a(a+1)t_{1}^{2}$. Since $e(H^{*})=t_{r}(k)n^{2}$, we obtain $e(H^{\prime})=t_{r}(k)n^{2}-b^{\prime}a(a+1)t_{1}^{2}$. Together with the bounds on $e_{H}(U_{i})$ we get
$e(H)\geq t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+\left\lfloor(k-a-1)n\frac{t-1}{2}\right\rfloor+b^{\prime}\frac{(2a^{2}+3a)t_{1}^{2}}{2}-\frac{r-1}{2}.$
So we obtain $e(H)\geq g_{1}(n,r,k,t)$ for the case $t_{1}>0$.
Finally for the case $t_{1}=0$ or $b=0,1$ we give a slight improvement to avoid the loss $-(r-1)/2$. Let $W^{\prime}:=\bigcup_{2\leq i\leq r}V_{i,a+1}$. If $b\geq 2$, starting from the partition $\{V_{i,j}\}$, we move at most $r$ vertices in $W^{\prime}$ to other rows, (e.g. move a vertex from $V_{2,a+1}$ to $V_{r,a+1}$, etc.), so that in the resulting partition $i\in[r]$, $U_{i}:=\bigcup_{j\in[a+1]}V_{i,j}$, all but at most one of the $U_{i}$’s, $2\leq i\leq r$ have even order. If $b\leq 1$, then $W^{\prime}=\emptyset$ and we do nothing in this step. Then as in the previous proof, we i) add a maximum $K_{t,t}$-free $(a+1)$-partite graph in $U_{1}$, and ii) for $i\in[2,r]$, add on $U_{i}$ a $\{K_{1,t},K_{2,2}\}$-free bipartite graph with $\lfloor|U_{i}|\frac{t-1}{2}\rfloor$ edges. Let $H$ be obtained by adding all edges between each two rows, i.e., adding $K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$. For $b\geq 2$, due to our control on the parities of the sizes, we have $\sum_{i\in[r]}e_{H}(U_{i})=z_{t}^{(a+1)}(n)+\left\lfloor(k-a-1)n\frac{t-1}{2}\right\rfloor$. For $b\leq 1$, we have $\sum_{i\in[r]}e_{H}(U_{i})=z_{t}^{(\lceil k/r\rceil)}(n)+(r-1)\left\lfloor\frac{t-1}{2}an\right\rfloor$. That is, we have $\sum_{i\in[r]}e_{H}(U_{i})=z_{t}^{(\lceil k/r\rceil)}(n)+h_{1}(n,r,k,t)$ in both cases. Moreover, note that the movement of the vertices does not change the number of edges in $J:=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$, because i) the induced subgraph on $\bigcup_{i\in[r],j\in[a]}V_{i,j}$ does not change, ii) the degree of every vertex of $\bigcup_{i\in[r]}V_{i,a+1}$ in the graph $J$ does not change. Thus, the graph $K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$ contains exactly $t_{r}(k)n^{2}$ edges and we obtain $e(H)=t_{r}(k)n^{2}+z_{t}^{(\lceil k/r\rceil)}(n)+h_{1}(n,r,k,t)$. ∎
Now the proof of Theorem 4 is completed. ∎
4. Proof of the upper bounds
Below we prove ($E1$).
Proof of ($Z$) $\Rightarrow$ ($E1$).
Given $a+1$ sets $V_{1},\dots,V_{a+1}$ of size $n$, we define an $(a+1)$-partite graph $G$ on $V_{1},\dots,V_{a+1}$ as follows. Let $V_{2}^{\prime}$ be a set of $n$ vertices consisting of $\lfloor n/2\rfloor$ vertices from $V_{1}$ and $\lceil n/2\rceil$ vertices from $V_{2}$. We place an extremal graph $G^{\prime}$ for $z_{t}^{(a)}(n)$ on $V_{2}^{\prime},V_{3},\dots,V_{a+1}$, in other words, $G^{\prime}$ is an $a$-partite $K_{t,t}$-free graph with $z_{t}^{(a)}(n)$ edges. Next we add a maximum bipartite $K_{t,t}$-free graph $G^{\prime\prime}$ on the remaining vertices of $V_{1}$ and $V_{2}$. By ($Z$), $e(G^{\prime\prime})\geq z_{t}^{(2)}(\lfloor n/2\rfloor)\geq\delta n^{2-1/t}$ for some $\delta>0$. Thus $G=G^{\prime}\cup G^{\prime\prime}$ is $K_{t,t}$-free and $e(G)=e(G^{\prime})+e(G^{\prime\prime})\geq z_{t}^{(a)}(n)+\delta n^{2-1/t}$. This gives ($E1$). ∎
Now we prove the upper bounds on $\operatorname{ex}_{k}(n,K_{r}(t))$.
The proof of Theorem 6 is done by a series of claims and estimates, and actually contains the proofs of Theorems 5 and 7. More precisely, the first half of the proof of Theorem 6 only requires ($Z$) and (consequently) ($E1$), but not ($E2$) or ($E3$). The (weaker) estimates obtained in this part of the proof suffice for Theorems 5 and 7. The second part of the proof contains refined estimates obtained from ($E2$) and ($E3$) and completes the proof of Theorem 6.
Outline of the proofs. Now we give an outline of our proofs. Let $G\in\mathcal{G}_{k}(n)$ be $K_{r+1}(t)$-free and has the maximum number of edges. Since $e(G)>t_{r}(k)n^{2}$, we can assume that $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$. Using that $e(G)$ is maximum, we can easily derive a minimum degree condition by some symmetrization arguments.
Next we define atypical vertices. Roughly speaking, there are two types of atypical vertices: the first type of vertices, denoted by $Z^{\prime\prime}\cup W^{\prime\prime}$, are the “wrong” ones that do not exist in ${\mathcal{T}}_{r,k}(n)$; the second type of vertices, denoted by $(W^{\prime}_{i}\setminus W_{i})\cup\bigcup_{j\neq i}Z_{j}^{i}$ for $i\in[r]$, are the vertices that are not in $U_{i}$ but behave like the vertices of $U_{i}$, in other words, they are in the wrong place. In the first half of the proof (i.e., the proofs of Theorem 5 and Theorem 7), we completely ignore the first type of atypical vertices because there are only a constant number of them (see $(P1)$ and $(P3)$) and they contribute only $O(n)$ to $e(G)$. For the second type of atypical vertices, there are only $o(n)$ of them (see $(P1)$ and $(P3)$) and we move them to appropriate rows and redefine our partition as $\tilde{U}_{1},\dots,\tilde{U}_{r}$ (see (4.2)). A key observation is that $Z_{j}^{i}\neq\emptyset$ (namely, there is a vertex in $Z_{j}$ but behaves as a vertex in $Z_{i}$) is possible only if $|W_{j}|\geq(1-o(1))n$.
Now we estimate $e(G)$. We split $E(G)$ into $E_{G}(\tilde{U}_{1}),\dots,E_{G}(\tilde{U}_{r})$, and $E(G^{\prime})$, where $G^{\prime}:=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have a relatively good estimate of $e(G^{\prime})$ (see Claim 4.4) taking into account that the partition is no longer balanced. In contrast, due to the second type of atypical vertices, we can only show that each $G[\tilde{U}_{i}]$ is “almost” $K_{t,t}$-free (see Claim 4.5). Similarly, we show that all but at most one rows are “almost” $K_{1,t}$-free (see Claim 4.7). Assuming that $e_{G}(\tilde{U}_{1})$ is the largest among all $e_{G}(\tilde{U}_{i})$, $i\in[r]$, we can collect the estimates and give an upper bound of $e(G)$. Now we show that $\tilde{U}_{1}$ has no atypical vertices (Claim 4.8), and thus $e_{G}(\tilde{U}_{1})\leq z_{t}^{(a+1)}(n)$. We further refine our estimate on $\tilde{U}_{i}$, $i>1$, and show that each second type atypical vertex contributes at most a constant number of edges to $E(G)\setminus E(G^{\prime})$ (Claims 4.9 and 4.10). In summary, $\tilde{U}_{1}$ is indeed $K_{t,t}$-free, $e_{G}(\tilde{U}_{i})=O(n)$ for $i>1$, and $|Z^{\prime\prime}\cup W^{\prime\prime}|=O(1)$, from which we conclude the proofs of Theorem 5 and Theorem 7.
To prove Theorem 6, we refine earlier estimates as follows. We first show that $|W_{1}^{\prime}|=(1-o(1))n$ and $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, where we use ($E2$). The rest of the proofs are further refinements of our estimates for various cases (see, e.g. the definition of $h_{2}$). In particular, at the very last step, we use ($E3$) to show that all of $Z_{1}^{i}$, $i\in[2,r]$ should be empty and $|W_{1}^{\prime}|$ should be equal to $n$.
We start with the following simple proposition.
Proposition 4.1.
Given $r,t\in\mathbb{N}$ and reals $\gamma,\varepsilon>0$ such that $\varepsilon^{2}>3r^{2}t^{2}\gamma$, and let $n$ be sufficiently large. Suppose $G$ is a $K_{r+1}(t)$-free graph with vertex partition $V=U_{1}\cup\cdots\cup U_{r}$ such that $|U_{i}|\geq n$ for $i\in[r]$ and $d(U_{i},U_{j})\geq 1-\gamma$, $i,j\in[r]$, $i\neq j$. Let $X\subseteq V$ be the set of vertices $v$ satisfies that $d(v,U_{i})\geq\varepsilon|U_{i}|$ for all $i\in[r]$. Then $|X|\leq 2(t-1)\varepsilon^{-rt}$.
Proof.
We call a copy of $K_{r}{(t)}$ in $G$ useful if it consists of exactly $t$ vertices from each of $U_{1},\dots,U_{r}$. We first show that for every $v\in X$, $N(v)$ contains many useful copies of $K_{r}{(t)}$. Indeed, $d(U_{i},U_{j})\geq 1-\gamma$ for every $i,j\in[r]$, $i\neq j$ implies that $G[U_{i},U_{j}]$ has at most $\gamma|U_{i}||U_{j}|$ non-edges. Since $d(v,U_{i})\geq\varepsilon|U_{i}|$ for all $i\in[r]$, take $W_{i}\subseteq N(v)\cap U_{i}$ of size exactly $\varepsilon|U_{i}|$. We can find $\prod_{i\in[r]}\binom{\varepsilon|U_{i}|}{t}$ $rt$-sets which consists of $t$ vertices from each $U_{i}$, amongst which, at most
$\sum_{i,j\in[r],i\neq j}\gamma|U_{i}||U_{j}|\cdot\left(\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}\right)\cdot\frac{t}{\varepsilon|U_{i}|}\frac{t}{\varepsilon|U_{j}|}\leq\frac{r^{2}t^{2}\gamma}{\varepsilon^{2}}\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}$
of them contain crossing non-edges. Therefore, $N(v)$ contains at least
$\left(1-\frac{r^{2}t^{2}\gamma}{\varepsilon^{2}}\right)\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}\geq\frac{\varepsilon^{rt}}{2}\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t}$
useful copies of $K_{r}{(t)}$, where we used that $r^{2}t^{2}\gamma\varepsilon^{-2}<1/3$. Since $G$ is $K_{r+1}(t)$-free, each useful copy $K$ of $K_{r}{(t)}$ is in $N(v)$ for at most $t-1$ choices of $v\in X$. Double counting on the number of pairs $(v,K)$ such that $K\subseteq N(v)$ is useful, we obtain that
$|X|\frac{\varepsilon^{rt}}{2}\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t}\leq(t-1)\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t},$
which gives $|X|\leq 2(t-1)\varepsilon^{-rt}$. ∎
Given integers $1\leq s\leq t$ and sufficiently large $m,n$, Kővári, Sós, Turán [MR0065617] showed that $z(m,n,s,t)\leq Cmn^{1-1/s}$ for some $C=C(t)>0$, that is, a bipartite graph $G$ with parts of size $m$ and $n$ has at most $Cmn^{1-1/s}$ edges if $G$ has no copy of $K_{s,t}$ where the part of size $s$ is in the part of $G$ of size $m$. This easily implies that $\operatorname{ex}(n,K_{s,t})\leq Cn^{2-1/t}$ for sufficiently large $n$.
Proofs of Theorems 5, 6, and 7.
Suppose ($Z$) holds, that is, $z_{t}^{(2)}(n)\geq cn^{2-1/t}$ for some $c>0$. Take $C=C(t)$ as in the Kővári–Sós–Turán result in the previous paragraph. We choose constants
$1/n\ll\gamma\ll\varepsilon\ll\varepsilon^{\prime}\ll 1/k,1/t,c,C.$
Suppose $G$ is $K_{r+1}(t)$-free and has the maximum number of edges, that is, $e(G)=\operatorname{ex}_{k}(n,K_{r+1}(t))$. Suppose further that $e(G)>g_{2}(n,r,k,t)>t_{r}(k)n^{2}$. By Theorem 3, $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$, where the color classes of $T$ are denoted by $U_{1},\dots,U_{r}$ with $U_{i}=W_{i}\cup Z_{i}$ such that $Z_{i}=V_{(i-1)a+1}\cup\cdots\cup V_{ia}$, $W_{i}=\emptyset$ if $b=0$ and $W_{i}$ is a subset of $V_{j}$ for some $j\geq ar+1$ otherwise, and $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$. For $i\in[r]$, if $W_{i}\neq\emptyset$, then let $q_{i}$ be the index such that $W_{i}\subseteq V_{q_{i}}$ and we know $ar<q_{i}\leq k$. For simplicity, we write $z_{t}(n)=z_{t}^{(\lceil k/r\rceil)}(n)$.
The fact that $G$ is $\gamma$-close to $T$ gives the following observation.
  1. (D0)
    for any $i\in[r]$, there exists $B_{i}\subseteq U_{i}$ of size at most $2\sqrt{\gamma}n$ such that for any $v\in U_{i}\setminus B_{i}$ and $A\subseteq\bigcup_{j\in[r]\setminus\{i\}}U_{j}$ satisfying that none of the vertices of $A$ is in the same cluster as $v$ is, we have $\overline{d}(v,A)\leq\sqrt{\gamma}n$.
To see it, fix $i\in[r]$ and write $U^{*}:=\bigcup_{j\neq i}U_{j}$. Since $G$ is $\gamma$-close to $T$, we have
$e_{G}(Z_{i},U^{*})\geq|Z_{i}||U^{*}|-\gamma n^{2},\quad\text{and}\quad e_{G}\left(W_{i},U^{*}\setminus V_{q_{i}}\right)\geq|W_{i}||U^{*}\setminus V_{q_{i}}|-\gamma n^{2}.$
Let $B_{i}^{\prime}\subseteq Z_{i}$ be the set of vertices $v$ such that $\overline{d}(v,U^{*})>\sqrt{\gamma}n$, and $B_{i}^{\prime\prime}\subseteq W_{i}$ be the set of vertices $w$ such that $\overline{d}(w,U^{*}\setminus V_{q_{i}})>\sqrt{\gamma}n$. The displayed line above implies that $|B_{i}^{\prime}|\leq\sqrt{\gamma}n$ and $|B_{i}^{\prime\prime}|\leq\sqrt{\gamma}n$. Now (D0) holds by setting $B_{i}=B_{i}^{\prime}\cup B_{i}^{\prime\prime}$.
Minimum degree
For $i\in[k]$, let $N_{i}:=N_{T}(u_{i})$ for some $u_{i}\in V_{i}$. Note that this is well-defined as the vertices of $V_{i}$ share the same neighborhood in $T$. Using the maximality of $e(G)$, we derive that for every $u\in V_{i}$, $i\in[k]$
$d_{G}(u)\geq d_{T}(u)-2t\gamma n.$
Indeed, since $G$ is $\gamma$-close to $T$, that is, $|E(G)\triangle E(T)|\leq\gamma n^{2}$, for each $i\in[k]$, we can greedily pick distinct $u_{1},\dots,u_{t}\in V_{i}$, such that $|N_{G}(u_{j})\triangle N_{i}|\leq\gamma n^{2}/(n-j)\leq 2\gamma n$, for $j\in[t]$. Let $N_{i}^{\prime}:=\bigcap_{j\in[t]}N_{G}(u_{j})$ and note that $|N_{i}^{\prime}\triangle N_{i}|\leq 2t\gamma n$. In particular, $|N_{i}^{\prime}|\geq|N_{i}|-2t\gamma n$. Now for a contradiction suppose there is $u\in V_{i}$ such that $d_{G}(u)<d_{T}(u)-2t\gamma n=|N_{i}|-2t\gamma n$. Then we replace $N_{G}(u)$ by $N_{i}^{\prime}$, that is, we disconnect all the edges of $u$ in $G$ and connect $u$ to the vertices of $N_{i}^{\prime}$. Thus, we obtain a $k$-partite graph on the same vertex set as $G$ and has more edges than $G$. Therefore, by the maximality of $G$, this new graph contains a copy of $K_{r+1}(t)$, denoted by $K$. Clearly, $K$ must contain the vertex $u$, as $G$ is $K_{r+1}(t)$-free. Moreover, $K$ must miss at least one vertex from $u_{1},\dots,u_{t}$, say $u_{j}$, because the set $\{u,u_{1},\dots,u_{t}\}$ is independent in $G$ and $K$ has independence number $t$. However, as the neighborhood of $u$ $N_{i}^{\prime}$ is a subset of $N_{G}(u_{j})$, we can replace $u$ by $u_{j}$ and still get a copy of $K_{r+1}(t)$, which is in $G$, a contradiction.
Therefore, comparing with the degrees in $T$, we derive that for any vertex $u$,
$d_{G}(u)\geq\begin{cases}(k-a)n-|W_{i}|-2t\gamma n,\,\,\,\,\hfill\text{ if $u\in Z_{i}$ for $i\in[r]$,}\\ (k-1-a)n-2t\gamma n,\,\,\,\,\hfill\text{ if }u\in W.\end{cases}$ (4.1)
Atypical vertices
In this step we identify a set of atypical vertices, that is, those behave differently from the majority of the vertices. Let $W:=\bigcup_{i\in[r]}W_{i}=V_{ar+1}\cup\cdots\cup V_{k}$. We define $W^{\prime\prime}:=\{v\in W:d(v,Z_{j})\geq\varepsilon n,\text{ for all }j\in[r]\}$ and $W_{i}^{\prime}:=\{v\in W:d(v,Z_{i})<\varepsilon n\}$. Then we have $W=W^{\prime\prime}\cup W_{1}^{\prime}\cup\cdots\cup W_{r}^{\prime}$. Next, for $i\in[r]$, let $Z^{\prime\prime}:=\bigcup_{i\in[r]}Z_{i}^{\prime\prime}$, where
$Z_{i}^{\prime\prime}:=\{v\in Z_{i}:d(v,Z_{j})\geq\varepsilon n,\text{ for all }j\in[r]\setminus\{i\}\quad\text{and}\quad d(v,U_{i})\geq\varepsilon n\}.$
Let $Z_{i}^{\prime}:=Z_{i}\setminus Z_{i}^{\prime\prime}$ for all $i$. We write $Z_{i}^{\prime}$ as $\bigcup_{j\in[r]}Z_{i}^{j}$, where $Z_{i}^{j}$, $j\neq i$, consists of the vertices $v$ such that $d(v,Z_{j})<\varepsilon n$, and $Z_{i}^{i}$ consists of the vertices $v$ such that $d(v,U_{i})<\varepsilon n$. Below are some useful properties of these sets.
Claim 4.2.
The following properties hold for all $i\in[r]$.
  1. $(P1)$
    $|W_{i}^{\prime}\setminus W_{i}|\leq 2\gamma n$ and $|W^{\prime\prime}|\leq C_{0}:=2t\varepsilon^{-rt}$.
  2. $(P2)$
    $W=W^{\prime\prime}\cup W_{1}^{\prime}\cup\cdots\cup W_{r}^{\prime}$ is a partition of $W$.
  3. $(P3)$
    $|Z_{i}^{\prime\prime}|\leq C_{0}$, $|Z_{i}^{j}|\leq\sqrt{\gamma}n$ for $j\neq i$, and $|Z_{i}^{i}|\geq(1-\sqrt{\gamma})an$.
  4. $(P4)$
    $\bigcup_{j\in[r]}Z_{i}^{j}$ is a partition of $Z_{i}^{\prime}$.
Proof.
Recall the definition of $W^{\prime\prime}$ and that $d(Z_{i},Z_{j})\geq 1-\gamma$ for distinct $i,j\in[r]$. Applying Proposition 4.1 to the graph $G[W^{\prime\prime}\cup Z]$ with vertex partition $(U_{1},\dots,U_{r})$, we obtain that $|W^{\prime\prime}|\leq C_{0}:=2t\varepsilon^{-rt}$. We next show that $|W_{i}^{\prime}\setminus W_{i}|\leq 2\gamma n$ for each $i\in[r]$. Indeed, because $G$ is $\gamma$-close to $T$, we have $e_{G}(Z_{i},W_{i}^{\prime}\setminus W_{i})\geq an|W_{i}^{\prime}\setminus W_{i}|-\gamma n^{2}$. On the other hand, by definition, $e_{G}(Z_{i},W_{i}^{\prime}\setminus W_{i})<|W_{i}^{\prime}\setminus W_{i}|\cdot\varepsilon n$. Thus, we get $|W_{i}^{\prime}\setminus W_{i}|<\gamma n/(a-\varepsilon)<2\gamma n$, verifying $(P1)$.
To see $(P2)$, suppose there is a vertex $v\in W_{i}^{\prime}\cap W_{j}^{\prime}$. By definition, $d(v)\leq(k-1)n-2(a-\varepsilon)n<(k-1-a)n-\sqrt{\gamma}n$, contradicting (4.1).
Next we show $(P3)$. Fix $i\in[r]$. Since $G$ is $\gamma$-close to $T$, we have $d(Z_{j},Z_{j^{\prime}})\geq 1-\gamma$ and $d(U_{i},Z_{j})\geq 1-\gamma$ for distinct $j,j^{\prime}\in[r]\setminus\{i\}$. Thus, we can apply Proposition 4.1 on $G[U_{i}\cup\bigcup_{j\neq i}Z_{j}]$ (with the obvious $r$-partition) and obtain $|Z_{i}^{\prime\prime}|\leq C_{0}$. Moreover, for $i\neq j$, from $d(Z_{i},Z_{j})\geq 1-\gamma$ we infer $|Z_{i}^{j}|\leq(\gamma/\varepsilon)n\leq\sqrt{\gamma}n$, as $\gamma\ll\varepsilon$. Therefore, we also get $|Z_{i}^{i}|\geq|Z_{i}|-|Z_{i}^{\prime\prime}|-\sum_{j\neq i}|Z_{i}^{j}|\geq an-C_{0}-(r-1)\gamma n/\varepsilon\geq(1-\sqrt{\gamma})an$.
Now we show $(P4)$. By definition, if $v\in Z_{i}^{i}$, then $d(v,U_{i})<\varepsilon n$; if $v\in Z_{i}^{j}$ for $j\neq i$, then $d(v,Z_{j})<\varepsilon n$. Thus, we have $Z_{i}^{\prime}\subseteq\bigcup_{j\in[r]}Z_{i}^{j}$ by definition. A vertex $v\in Z_{i}^{i}\cap Z_{i}^{j}$, $j\neq i$, satisfies that $d(v)<kn-(|U_{i}|-\varepsilon n)-(a-\varepsilon)n\leq(k-a)n-|W_{i}|-(1-2\varepsilon)n$, contradicting (4.1). A vertex $v\in Z_{i}^{j}\cap Z_{i}^{j^{\prime}}$ for distinct $j,j^{\prime}\in[r]\setminus\{i\}$ satisfies that $d(v)<(k-1)n-2(a-\varepsilon)n\leq(k-a-2)n+2\varepsilon n$, contradicting (4.1) as well. Thus, $\bigcup_{j\in[r]}Z_{i}^{j}$ is a partition of $Z_{i}^{\prime}$. ∎
For $i\in[r]$, our refined partition is defined by
$\tilde{U}_{i}:=\tilde{Z}_{i}\cup W_{i}^{\prime},\text{ where }\tilde{Z}_{i}:=\bigcup_{j\in[r]}Z_{j}^{i}.$ (4.2)
Then $V(G)=Z^{\prime\prime}\cup W^{\prime\prime}\cup\bigcup_{i\in[r]}\tilde{U}_{i}$. Note that for any $v\in\tilde{U}_{i}$, we have $d(v,Z_{i}^{i})\leq d(v,Z_{i})\leq\varepsilon n$, and thus $d(v,\tilde{Z}_{i})\leq\varepsilon n+(r-1)\sqrt{\gamma}n$ by $(P3)$.
For every $i\in[r]$, note that $(P1)$ implies that $|W_{i}\setminus W_{i}^{\prime}|\leq C_{0}+(r-1)2\gamma n\leq 2r\gamma n$, and similarly $(P3)$ implies that $|Z_{i}\setminus\tilde{Z}_{i}|\leq C_{0}+(r-1)\sqrt{\gamma}n\leq r\sqrt{\gamma}n$.
We now derive a more handy minimum degree condition. For convenience, define $\overline{d}(v,A)=|A|-d(v,A)$. For $v\in Z_{i}^{i}$, we have $\overline{d}(v,\tilde{U}_{i})\geq\overline{d}(v,U_{i})-|U_{i}\setminus\tilde{U}_{i}|$. Since $\overline{d}(v,U_{i})>an+|W_{i}|-\varepsilon n$ and $|U_{i}\setminus\tilde{U}_{i}|\leq|Z_{i}\setminus\tilde{Z}_{i}|+|W_{i}\setminus W^{\prime}_{i}|\leq\varepsilon n/2$, we have $\overline{d}(v,\tilde{U}_{i})\geq an+|W_{i}|-\varepsilon n-\varepsilon n/2$. By (4.1), $\overline{d}(v)\leq an+|W_{i}|+\sqrt{\gamma}n$. It follows that $\overline{d}(v,V\setminus\tilde{U}_{i})\leq 2\varepsilon n$. Now consider $v\in\tilde{U}_{i}\setminus Z^{i}_{i}$. The definition of $\tilde{U}_{i}$ implies that $d(v,Z_{i})<\varepsilon n$ and $\overline{d}(v,Z_{i})>an-\varepsilon n$. Assume $v\in V_{j}$. Then $V_{j}\cap Z_{i}=\emptyset$ and trivially $\overline{d}(v,V_{j})=n$. It follows that $\overline{d}(v,Z_{i}\cup V_{j})>(a+1)n-\varepsilon n$. Hence $\overline{d}(v,\tilde{Z}_{i}\cup V_{j})\geq\overline{d}(v,Z_{i}\cup V_{j})-|Z_{i}\setminus\tilde{Z}_{i}|>(a+1)n-\frac{3}{2}\varepsilon n$. On the other hand, either case of (4.1) implies that $\overline{d}(v)\leq(a+1)n+\sqrt{\gamma}n$. Consequently, $\overline{d}(v,V\setminus(\tilde{Z}_{i}\cup V_{j}))\leq 2\varepsilon n$. In summary, for $i\in[r]$ and $j\in[k]$,
  1. (Deg)
    If $v\in Z_{i}^{i}$, then $\overline{d}(v,V\setminus\tilde{U}_{i})\leq 2\varepsilon n$; if $v\in(\tilde{U}_{i}\setminus Z^{i}_{i})\cap V_{j}$, then $\overline{d}(v,V\setminus(\tilde{Z}_{i}\cup V_{j}))\leq 2\varepsilon n$.
Next we prove further properties on $Z_{i}^{j}$ and $\tilde{Z}_{j}$.
Claim 4.3.
If $Z_{i}^{j}\neq\emptyset$ for some $i\neq j$, then the following holds.
  1. $(Q1)$
    For $v\in Z_{i}^{j}$ and $A\subseteq V(G)\setminus(Z_{i}\cup Z_{j})$, we have $d(v,A)\geq|A|-\varepsilon n-\sqrt{\gamma}n$.
  2. $(Q2)$
    $|W_{i}|\geq(1-\varepsilon-\sqrt{\gamma})n$.
  3. $(Q3)$
    If $|\tilde{Z}_{j}\setminus Z_{j}|\geq t$, then $|W_{j}|\leq 2t\varepsilon n$.
Proof.
Note that $d(v,Z_{j})\leq\varepsilon n$ and $d(v,Z_{i})\leq(a-1)n$, that is, $v$ has at least $n+(an-\varepsilon n)=(a+1)n-\varepsilon n$ non-neighbors in $Z_{i}\cup Z_{j}$. On the other hand, (4.1) says that $v$ has at most $an+|W_{i}|+\sqrt{\gamma}n$ non-neighbors in $G$. Combining these two we get that $v$ has at most $|W_{i}|-n+\varepsilon n+\sqrt{\gamma}n\leq\varepsilon n+\sqrt{\gamma}n$ non-neighbors outside $Z_{i}\cup Z_{j}$, and thus $(Q1)$ holds. The fact that $|W_{i}|-n+\varepsilon n+\sqrt{\gamma}n\geq 0$ implies $(Q2)$.
For $(Q3)$, suppose to the contrary, $|\tilde{Z}_{j}\setminus Z_{j}|\geq t$ and $|W_{j}|>2t\varepsilon n$. By $(Q1)$ with $A=W_{j}$, arbitrary $t$ vertices in $\tilde{Z}_{j}\setminus Z_{j}$ have at least $|W_{j}|-t(\varepsilon+\sqrt{\gamma})n\geq t$ common neighbors in $W_{j}$. We thus obtain a copy of $K_{t,t}$ with one part in $\tilde{Z}_{j}\setminus Z_{j}$ and the other part in $W_{j}$ – denote its vertex set by $B$. For any $i^{\prime}\in[r]\setminus\{j\}$ such that $B\cap Z_{i^{\prime}}^{j}\neq\emptyset$, we have $|W_{i^{\prime}}|\geq(1-\varepsilon-\sqrt{\gamma})n$ by $(Q2)$. Since $|W_{j}|>2t\varepsilon n$, $W_{i^{\prime}}$ and $W_{j}$ do not belong to the same cluster, and thus no vertex of $B$ is in the same cluster that contains $W_{i^{\prime}}$, which implies that the vertices of $B$ have at least $|W_{i^{\prime}}|-2t(2\varepsilon n)\geq n/2$ common neighbors in $W_{i^{\prime}}$ by (Deg). For any $i^{\prime\prime}\in[r]\setminus\{j\}$ such that $B\cap Z_{i^{\prime\prime}}^{j}=\emptyset$ (and thus $B\cap Z_{i^{\prime\prime}}=\emptyset$), by (Deg) we have that the vertices of $B$ have at least $n/2$ common neighbors in $Z_{i^{\prime\prime}}$. Because $G$ is $\gamma$-close to $T$, these common neighborhoods, each of size at least $n/2$, have densities close to one between each pair, and thus contain a copy of $K_{r-1}{(t)}$. Together with $B$, they form a copy of $K_{r+1}(t)$ in $G$, a contradiction. ∎
In particular, when $b=0$ (and thus $W_{i}=\emptyset$ for all $i$), $(Q2)$ implies that $Z_{i}^{j}=\emptyset$ whenever $i\neq j$. Consequently,
$\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}\quad\text{for all $i\in[r]$ when $b=0$}.$ (4.3)
Let $L\subseteq[r]$ be the set of indices $i$ such that $|W_{i}|\geq(1-\varepsilon-\sqrt{\gamma})n$. $(Q2)$ and $(Q3)$ imply that
  • for $i\in[r]\setminus L$, we have $Z_{i}^{j}=\emptyset$ for $j\neq i$.
  • for $i\in L$, $|\tilde{Z}_{i}\setminus Z_{i}|\leq t-1$ and thus $|\tilde{Z}_{i}|\leq an+t-1$.
Estimate $e(G)$
Let $G^{\prime}=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)$.
Since $G^{\prime}$ is $r$-partite, it is $K_{r+1}$-free. As $G^{\prime}$ is a subgraph of $G\in\mathcal{G}_{k}(n)$, we have $e(G^{\prime})\leq t_{r}(k)n^{2}$ (but this is not good enough when $b>0$). Below we give an upper bound for $e(G^{\prime})$, which will be used throughout the proof. Recall that $T=K(V_{1},\dots,V_{k})\cap K(U_{1},\dots,U_{r})$ has precisely $t_{r}(k)n^{2}$ edges,
Claim 4.4.
We have $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(\beta_{i}-\alpha_{i})$, where
$\beta_{i}:=$ $\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|\right)\text{and }$
$\alpha_{i}:=$ $|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|+e_{T}(W_{i}^{\prime})+e_{T}(\tilde{Z}_{i}\setminus Z_{i}).$
Proof.
We first obtain $G^{(0)}:=K(Z_{1}\cup W_{1}^{\prime},\dots,Z_{r}\cup W_{r}^{\prime})\cap K(V_{1},\dots,V_{k})$ from $T$. During this process, we lose the edges of $T$ between $W_{i}$ and $W_{j}$, $j\neq i$, if both ends of the edges are placed in $W^{\prime}_{i}$. Thus
$e(G^{(0)})=t_{r}(k)n^{2}-\sum_{i\in[r]}e_{T}(W_{i}^{\prime}).$ (4.4)
We imagine a dynamic process of obtaining $G^{\prime}$ from $G^{(0)}$ by recursively moving vertices. To estimate $e(G^{\prime})$, we track the changes of the edges with respect to complete $r$-partite graphs (but also respecting the $k$-partition of $G$). More precisely, for $l>0$, let
$G^{(l)}:=K(Z_{1}^{(l)}\cup W_{1}^{\prime},\dots,Z_{r}^{(l)}\cup W_{r}^{\prime})\cap K(V_{1},\dots,V_{k})$
such that the $r$-partition of $G^{(l)}$ can be obtained by moving exactly one vertex from the partition of $G^{(l-1)}$. The process terminates after $m:=\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}|$ steps and thus $G^{\prime}$ is a subgraph of $G^{(m)}$. Furthermore, throughout the process, we only move vertices from the color classes in $L$ to other color classes. Therefore, we can give a linear ordering to the members of $L$, and for $i\in L$ we move vertices from $Z_{i}$ only after we have moved the vertices in color classes $j$ prior to $i$ (denoted by $j<_{L}i$). Now, in the $l$-th step, suppose we move $v$ from $Z_{j}^{(l-1)}$ to $Z_{i}^{(l-1)}$, namely, $v\in Z_{j}^{i}$, then the change is
$e(G^{(l)})-e(G^{(l-1)})=|Z_{j}^{(l-1)}\setminus V_{p}|+|W_{j}^{\prime}|-|\tilde{Z}_{i}^{(l-1)}|-|W_{i}^{\prime}|,$
where $V_{p}\ni v$ and $\tilde{Z}_{i}^{(l-1)}=Z_{i}^{(l-1)}\setminus V_{p}$.
Note that we have $|Z_{j}^{(l-1)}\setminus V_{p}|\leq(a-1)n+|\tilde{Z}_{j}\setminus Z_{j}|$. Moreover for any $j^{\prime}<_{L}j$, we have $Z_{j^{\prime}}^{i}\subseteq Z_{i}^{(l-1)}$. Therefore, we have $|\tilde{Z}_{i}^{(l-1)}|\geq an-|Z_{i}\setminus\tilde{Z}_{i}|+\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|$. Putting all these together, we get
$e(G^{(l)})-e(G^{(l-1)})\leq|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|-|W_{i}^{\prime}|.$
Recalling that we moved $v$ from $Z_{j}^{(l-1)}$ to $Z_{i}^{(l-1)}$ at the $l$-th step, we obtain
$e(G^{\prime})-e(G^{(0)})\leq\sum_{l=1}^{m}\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|-|W_{i}^{\prime}|\right),$
where $i,j$ depends on $l$. Since $m=\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}|$, we have
$\sum_{l=1}^{m}(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-|W_{i}^{\prime}|)$
$=$ $\sum_{i\in[r]}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-|W_{i}^{\prime}|)$
$=$ $\sum_{i\in[r]}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|)-\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|.$
Moreover, it is not hard to see that
$\sum_{l=1}^{m}\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|=\sum_{i\in[r]}\sum_{\{j_{1},j_{2}\}\in\binom{L\setminus\{i\}}{2}}|Z_{j_{1}}^{i}||Z_{j_{2}}^{i}|=\sum_{i\in[r]}e_{T}(\tilde{Z}_{i}\setminus Z_{i}).$
Now the claim follows by combining these estimates with (4.4). ∎
Our main task is bounding the number of edges in each $\tilde{U}_{i}$. For $i\in[r]$, we have $e(G[\tilde{U}_{i}])=e(Z_{i}^{i},G[\tilde{U}_{i}])+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})$. To bound $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}((\tilde{Z}_{i}\setminus Z_{i})\cup W_{i}^{\prime})$, we note that $e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})\leq|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|$ and $e_{G}(W_{i}^{\prime})\leq e_{T}(W_{i}^{\prime})$. However, we may not have $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$ because each $Z_{j}^{i}$ is an independent set in $T$, but may not be independent in $G$ when $a\geq 2$. Thus, $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+\sum_{j\neq i}e_{G}(Z_{j}^{i})$. Putting these together, for each $i\in[r]$, we have
$e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})+e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq\alpha_{i}+\sum_{j\neq i}e_{G}(Z_{j}^{i}).$ (4.5)
Let $f_{i}:=e(Z_{i}^{i},G[\tilde{U}_{i}])$. Applying Claim 4.4 and (4.5), we derive that
$e(G)$ $=e(G^{\prime})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}f_{i}+e_{G}((\tilde{U}_{i}\setminus Z^{i}_{i})$ (4.6)
$\leq t_{r}(k)n^{2}+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}\left(f_{i}+\beta_{i}+\sum_{j\neq i}e_{G}(Z_{j}^{i})\right)$ (4.7)
The following claim shows that a large portion of $G[\tilde{U}_{i}]$ is $K_{t,t}$-free (though we cannot prove that the entire $G[\tilde{U}_{i}]$ is $K_{t,t}$-free).
Claim 4.5.
The following holds for all $i\in[r]$.
  1. $(K1)$
    Both $G[\tilde{Z}_{i}]$ and $G[Z_{i}^{i}\cup W_{i}^{\prime}]$ are $K_{t,t}$-free.
  2. $(K2)$
    If $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$, then $|W_{i}^{\prime}\setminus V_{q_{i}}|\leq t-1$.
  3. $(K3)$
    If $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$, then $G[\tilde{Z}_{i}\cup(W_{i}^{\prime}\cap V_{q_{i}})]$ is $K_{t,t}$-free.
Proof.
For $(K1)$, suppose there is a copy of $K_{t,t}$ in $\tilde{U}_{i}$, with vertex set denoted by $B$, contained in $\tilde{Z}_{i}$ or in $Z_{i}^{i}\cup W_{i}^{\prime}$. Let $N_{B}$ be the set of common neighbors of these $2t$ vertices of $B$. First assume that $B\subseteq\tilde{Z}_{i}$. Then for any $j\in L\setminus\{i\}$, by (Deg) we have $|N_{B}\cap W_{j}^{\prime}|\geq|W_{j}^{\prime}|-4t\varepsilon n$, and thus by $(P1)$ $|N_{B}\cap W_{j}\cap W_{j}^{\prime}|\geq|W_{j}^{\prime}|-4t\varepsilon n-2\gamma n\geq n/2$. For any $j\notin L\cup\{i\}$, because $B\cap Z_{j}=\emptyset$ by $(Q2)$, we have $|N_{B}\cap Z_{j}|\geq an-4t\varepsilon n\geq n/2$ by (Deg). Note that every set in $\{N_{B}\cap Z_{j}:j\notin L\}\cup\{N_{B}\cap W_{j}\cap W_{j}^{\prime}:j\in L\}$ has size at least $n/2$ and every pair of them has density at least $1-4\gamma$. Therefore we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, which gives rise to a copy of $K_{r+1}(t)$ together with $B$, a contradiction.
Second we assume that $B\subseteq Z_{i}^{i}\cup W_{i}^{\prime}$. In this case we note that for any $j\neq i$, we have $B\cap Z_{j}=\emptyset$ and thus by (Deg), we have $|N_{B}\cap Z_{j}^{j}|\geq(1-\sqrt{\gamma})an-4t\varepsilon n\geq n/2$. Then as these sets have high pairwise densities, as in the previous case, we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, yielding a copy of $K_{r+1}(t)$ together with $B$, a contradiction. Now $(K1)$ is proved.
Now we turn to $(K2)$, and suppose $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$ and thus $|W_{i}\cap W_{i}^{\prime}|>2t\varepsilon n$ by $(P1)$. First, if $W_{i}^{\prime}$ contains at least $t$ vertices which are not from $V_{q_{i}}$ (the cluster containing $W_{i}$), then by (Deg), each of these vertices have at most $2\varepsilon n$ non-neighbors in $W_{i}\cap W_{i}^{\prime}$, and thus we can find a copy of $K_{t,t}$ in $W_{i}^{\prime}$, contradicting $(K1)$. So we have $|W_{i}^{\prime}\setminus V_{q_{i}}|\leq t-1$.
For $(K3)$, suppose there is a copy of $K_{t,t}$ as stated in the claim, whose vertex set is denoted by $B$. As in the previous paragraph, we have $|W_{i}|>2t\varepsilon n$ by $(P1)$. Now observe crucially that if $B\cap Z_{j}^{i}\neq\emptyset$, then by $(Q2)$ $|W_{i}|+|W_{j}|>n$, and thus, $W_{i}$ and $W_{j}$ are not from the same cluster. So by (Deg), for any $j\in[r-1]\setminus\{i\}$, if $B\cap Z_{j}^{i}=\emptyset$, then the vertices of $B$ have large common neighborhoods in $Z_{j}^{j}$; if $B\cap Z_{j}^{i}\neq\emptyset$, then the vertices of $B$ have large common neighborhoods in $W_{j}\cap W_{j}^{\prime}$ (note that $|W_{j}|\geq(1-\varepsilon-\sqrt{\gamma})n$ by $(Q2)$). Since each of these common neighborhoods have size at least $n/2$ and each pair of them has high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, yielding a copy of $K_{r+1}(t)$ together with $B$, a contradiction. ∎
We now derive an upper bound for $e(G)$ from Claims 4.4 and 4.5. For $i\in[r]$, we have $\beta_{i}\leq\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|)$ as $|W_{j}^{\prime}|-n\leq|W_{j}^{\prime}\setminus V_{q_{j}}|$. Fix $j\in L\setminus\{i\}$. Note that $|W_{j}|\geq(1-2\varepsilon)n$. We have $|\tilde{Z}_{j}\setminus Z_{j}|\leq t-1$ by $(Q3)$, and $|W_{j}^{\prime}\setminus V_{q_{j}}|\leq t-1$ by $(K2)$. If $|W_{i}|>n/2$, then $|Z_{j}^{i}|\leq t-1$ by $(Q3)$. Furthermore, since $|Z_{i}\setminus\tilde{Z}_{i}|\leq(r-1)\sqrt{\gamma}n+C_{0}$ by $(P3)$, it follows that
$|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|\right)\leq(t-1)(t-1+t-1+(r-1)\sqrt{\gamma}n+C_{0})\leq(t-1)r\sqrt{\gamma}n.$
Otherwise $|W_{i}|\leq n/2$, and by $(Q2)$, we have $Z_{i}^{i^{\prime}}=\emptyset$ for any $i^{\prime}\neq i$. This implies $|Z_{i}\setminus\tilde{Z}_{i}|=|Z_{i}^{\prime\prime}|\leq C_{0}$. Using $|Z_{j}^{i}|\leq\sqrt{\gamma}n$, $(Q3)$, and $(K2)$, we derive that
$|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|\right)\leq\sqrt{\gamma}n(2(t-1)+C_{0})\leq 2C_{0}\sqrt{\gamma}n.$
Summarizing these two cases for all $j\in L\setminus\{i\}$, we obtain that $\beta_{i}\leq(r-1)2C_{0}\sqrt{\gamma}n$, and consequently,
$\sum_{i\in[r]}\beta_{i}\leq 2(r-1)rC_{0}\sqrt{\gamma}n.$ (4.8)
On the other hand, for all $i\neq j$, the graph $G[Z_{j}^{i}]$ is $K_{t,t}$-free by $(K1)$ and thus, by $(P3)$, $\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})\leq r(r-1)C\left(\sqrt{\gamma}n\right)^{2-1/t}$. Applying this with (4.7), (4.8), and the fact that $e(Z^{\prime\prime}\cup W^{\prime\prime},G)\leq(r+1)C_{0}kn$, we obtain that
$e(G)$ $\leq t_{r}(k)n^{2}+(r+1)C_{0}kn+\sum_{i\in[r]}f_{i}+2(r-1)rC_{0}\sqrt{\gamma}n+\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})$ (4.9)
$\leq t_{r}(k)n^{2}+\sum_{i\in[r]}f_{i}+r^{2}C\sqrt{\gamma}n^{2-1/t}.$
Using the assumption $e(G)\geq g_{2}(n,r,k,t)\geq t_{r}(k)n^{2}+z_{t}(n)$, we infer that
$\sum_{i\in[r]}f_{i}\geq z_{t}(n)-r^{2}C\sqrt{\gamma}n^{2-1/t}\geq\frac{c}{2}n^{2-1/t}$ (4.10)
by using ($Z$), $z_{t}(n)\geq z_{t}^{(2)}(n)\geq cn^{2-1/t}$, and $\gamma\ll 1$.
In the rest of the proof we will derive a contradiction to (4.10). We first study the existence of $K_{1,t}$ in each color class. To do so, we consider a copy of $K_{3}(t)$ in $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ for some $i\neq j$.
Claim 4.6.
For any $i\neq j$, if $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ contains a copy $K$ of $K_{3}{(t)}$, then there exists $l\notin\{i,j\}$ such that $V(K)$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$.
Proof.
We may assume that $r>2$ as otherwise the claim is trivial. Suppose to the contrary that there is a copy $K$ of $K_{3}{(t)}$ in, say, $\tilde{U}_{1}$ and $\tilde{U}_{2}$, such that for every $l\in[3,r]$, there is a cluster in $U_{l}$ which does not intersect $B:=V(K)$. Let $V_{i_{l}}$ be a cluster in $Z_{l}$ such that $B\cap V_{i_{l}}=\emptyset$, and if there is no such cluster in $Z_{l}$, then we choose $V_{i_{l}}=V_{q_{l}}$. Note that in the former case, we have $|\tilde{U}_{l}\cap V_{i_{l}}|=|Z_{l}^{l}\cap V_{i_{l}}|\geq(1-\sqrt{\gamma}a)n$. In the latter case, we have $Z_{l}^{1}\neq\emptyset$ or $Z_{l}^{2}\neq\emptyset$, which implies that $|W_{l}|\geq(1-2\varepsilon)n$ by $(Q2)$, and thus $|\tilde{U}_{l}\cap V_{i_{l}}|=|W_{l}^{\prime}\cap V_{i_{l}}|\geq(1-3\varepsilon)n$. Now, by (Deg), every vertex in $B$ has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{l}\cap V_{i_{l}}$ for each $l\in[3,r]$. Since for every $l$ we have $|\tilde{U}_{l}\cap V_{i_{l}}|\geq 0.9n$, one can find large common neighborhoods (e.g. of size $n/2$) of all vertices of $B$ in each $\tilde{U}_{l}\cap V_{i_{l}}$, and then find a copy of $K_{r-2}{(t)}$ in these sets. Altogether we obtain a copy of $K_{r+1}(t)$, a contradiction.
Therefore, for such a copy $K$ of $K_{3}(t)$, there exists $l\notin\{i,j\}$ such that $K$ must intersect all clusters of $U_{l}$. Since $V(K)\cap Z_{l}\neq\emptyset$, we have $Z_{l}^{i}\neq\emptyset$ or $Z_{l}^{j}\neq\emptyset$. Then by $(Q2)$, $|W_{l}|\geq(1-2\varepsilon)n$ and in particular, $V_{q_{l}}\neq\emptyset$. Therefore $V(K)\cap V_{q_{l}}\neq\emptyset$. ∎
Claim 4.7.
For all but exactly one $j\in[r]$, we have $d(v,Z_{j}^{j})\leq t-1$ for all $v\in\tilde{U}_{j}$.
Proof.
First assume that there exists $j\in[r]$ such that $G[\tilde{U}_{j}]$ contains a copy of $K_{1,t}$, with vertex set denoted by $\{v,u_{1},\dots,u_{t}\}$, $v\in\tilde{U}_{j}$ and $u_{1},\dots,u_{t}\in Z_{j}^{j}$. Fix $i\in[r]\setminus\{j\}$ and let $N^{\prime}$ be the set of common neighbors of $u_{1},\dots,u_{t}$ in $\tilde{U}_{i}\cap U_{i}$. Suppose $v\in V_{p}$ and let $N$ be the set of common neighbors of these $t+1$ vertices in $\tilde{U}_{i}\cap U_{i}$. In particular, $N\subseteq N^{\prime}$ and $N$ is almost equal to the union of $a$ or $a+1$ clusters in $\tilde{U}_{i}$. Suppose there is a copy of $K_{t-1,t}$ with parts $S_{1}$ of size $t-1$ and $S_{2}$ of size $t$ such that $S_{1}\subseteq N^{\prime}$ and $S_{2}\subseteq N$. Then by Claim 4.6, there exists $l\in[r]\setminus\{i,j\}$ such that $B\cap Z_{l}\neq\emptyset$ and $B\cap V_{q_{l}}\neq\emptyset$, where $B$ denotes the vertex set of the copy of $K_{3}{(t)}$. This is impossible since $v$ is the only possible vertex in $B\cap(Z_{l}\cup V_{q_{l}})$ and can not satisfy both. Therefore, $G[N,N^{\prime}]$ is $K_{t-1,t}$-free, implying that $e_{G}(N,N^{\prime})=O(n^{2-1/(t-1)})$.
By $(P1)$$(P3)$ and (Deg), we have $|(\tilde{U}_{i}\setminus V_{p})\setminus N|\leq 3(t+1)\varepsilon n$ and $|\tilde{U}_{i}\setminus N^{\prime}|\leq 3t\varepsilon n$. Let $E^{i}$ be the set of the edges incident to $(\tilde{U}_{i}\setminus V_{p})\setminus N$ or $\tilde{U}_{i}\setminus N^{\prime}$ and counted in $f_{i}$. We split it to $E^{i}\cap E_{G}(Z_{i}^{i})$ and $E^{i}\cap E_{G}(\tilde{U}_{i}\setminus Z_{i}^{i},Z_{i}^{i})$. Note that by $(K1)$, each of the terms can be split further into at most $k$ $K_{t,t}$-free bipartite graphs, each with one part of size at most $3(t+1)\varepsilon n$ and the other part of size at most $(1+(r-2)\sqrt{\gamma})an$. Therefore, we obtain that
$f_{i}=O(\varepsilon n^{2-1/t})+O(n^{2-1/(t-1)})=O(\varepsilon n^{2-1/t}).$ (4.11)
Now assume there exist distinct $j_{1},j_{2}\in[r]$ such that each $G[\tilde{U}_{j_{i}}]$ contains a copy of $K_{1,t}$ whose part of size $t$ is in $Z_{j_{i}}^{j_{i}}$. The arguments above imply that (4.11) holds for all $i\in[r]$, and consequently, $\sum_{i\in[r]}f_{i}=O(\varepsilon n^{2-1/t})$, contradicting (4.10).
On the other hand, if $d(v,Z_{j}^{j})\leq t-1$ for all $j\in[r]$ and all $v\in\tilde{U}_{j}$, then $\sum_{j\in[r]}f_{j}\leq(t-1)kn$, again contradicting (4.10). ∎
By Claim 4.7, without loss of generality, we assume that,
for $\text{for }i\geq 2,\quad d(v,Z_{i}^{i})\leq t-1\ \text{for all}\ v\in\tilde{U}_{i},\quad\text{and thus},\quad f_{i}\leq(t-1)|\tilde{U}_{i}|.$ (4.12)
If $b=0$, then $\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}$ for all $i$ by (4.3). In this case $\tilde{U}_{1}$ is $K_{t,t}$-free by $(K1)$ and $\tilde{U}_{i}$ is $K_{1,t}$-free for all $i\geq 2$ by (4.12). Since $G$ is $\gamma$-close to $K_{r}(an)$, $G[U_{i}\setminus Z^{\prime\prime},U_{j}\setminus Z^{\prime\prime}]$ is almost complete for all $i\neq j$. This completes the proof of Theorem 7 with $Z:=Z^{\prime\prime}$ (note that ($Z$) holds because $t=2,3$).
Furthermore, when $b=0$, together with $e(G^{\prime})\leq t_{r}(k)n^{2}$, we conclude that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)\left\lfloor\frac{t-1}{2}an\right\rfloor+|Z^{\prime\prime}|kn.$ (4.13)
By $(P3)$, $|Z^{\prime\prime}|\leq rC_{0}$, and this proves Theorem 5 for the case $b=0$.
We thus assume $b>0$ for the rest of the proof of Theorem 5. By (4.10) and (4.12), we get
$f_{1}\geq z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}.$ (4.14)
In particular, we claim that
$|W_{1}|>3t\varepsilon n$ (4.15)
(which we will refine a moment later). Indeed, the edges counted in $f_{1}$ can be covered by $G[Z_{1}^{1}]$, $G[Z_{1}^{1},W_{1}\cap W_{1}^{\prime}]$, and at most $k$ $K_{t,t}$-free bipartite graphs, each with a part of size at most $\sqrt{\gamma}n$ and a part of size at most $an$. If $|W_{1}|\leq 3t\varepsilon n$, then $e_{G}(Z_{1}^{1},W_{1}\cap W_{1}^{\prime})=O(\varepsilon n^{2-1/t})$. Together with $e_{G}(Z_{1}^{1})\leq z_{t}^{(a)}(n)$, we have
$f_{1}\leq z_{t}^{(a)}(n)+O(\varepsilon n^{2-1/t})<z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}$
by ($E1$), contradicting (4.14).
Now we can give a much cleaner structure, shown in a series of claims below.
Claim 4.8.
Suppose $b>0$. Then $\tilde{U}_{1}=Z_{1}^{1}\cup W_{1}^{\prime}$ and $W_{1}^{\prime}\subseteq V_{q_{1}}$.
Proof.
Suppose to the contrary, there is a vertex $v$ in $\tilde{U}_{1}\setminus(Z_{1}^{1}\cup W_{1}^{\prime})$ or $W_{1}^{\prime}\setminus V_{q_{1}}$, namely, $v\in Z_{i}^{1}$ for some $2\leq i\leq r$ or $v\in W_{1}^{\prime}\setminus V_{q_{1}}$. Suppose $v\in V_{l}$. Then $l\neq q_{1}$. Morepver, if $i$ is defined, then $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus(\tilde{Z}_{i}\cup V_{l})$; otherwise, $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus V_{l}$. By (Deg), we have $\overline{d}(v,W_{1}^{\prime}\cap V_{q_{1}})\leq 2\varepsilon n$. Let $N:=W_{1}^{\prime}\cap V_{q_{1}}\cap N(v)$. We have $|(W_{1}^{\prime}\cap V_{q_{1}})\setminus N|\leq 2\varepsilon N$. Since $|W_{1}^{\prime}\setminus V_{q_{1}}|\leq|W_{1}^{\prime}\setminus W_{1}|\leq 2\gamma n$, it follows that $|W^{\prime}_{1}\setminus N|\leq 2\varepsilon n+2\gamma n\leq 3\varepsilon n$.
Recall (4.15), $|W_{1}|>3t\varepsilon n$. By $(K3)$ (if $v\in\tilde{Z}_{1}\setminus Z_{1}$) or $(K1)$ (if $v\in W_{1}^{\prime}\setminus V_{q_{1}}$), we know that $G[Z_{1}^{1},N]$ contains no $K_{t-1,t}$ with the part of size $t$ in $N$. This implies that $e_{G}(Z_{1}^{1},N)=O(n^{2-1/(t-1)})$. Furthermore, by $(P3)$ and $(K1)$, $G[\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1}]$ is a $K_{t,t}$-free graph with one part of size at most $(r-1)\sqrt{\gamma}n$ and the other part of size at most $an$. Thus, $e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})\leq C(r-1)\sqrt{\gamma}n(an)^{1-1/t}$. By the similar arguments, we have $e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})\leq C(3\varepsilon n)(an)^{1-1/t}$.
Putting these bounds together with $e_{G}(Z_{1}^{1})\leq z_{t}^{(a)}(n)$, we get
$f_{1}$ $=e_{G}(Z_{1}^{1})+e_{G}(Z_{1}^{1},N)+e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})+e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})$
$\leq z_{t}^{(a)}(n)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})+O(\varepsilon n^{2-1/t}).$
By ($E1$), this implies that $f_{1}<z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}$, contradicting (4.14). ∎
Claim 4.8 shows that $\tilde{U}_{1}$ has no atypical vertices and is thus $K_{t,t}$-free by $(K1)$. Furthermore, since $\tilde{U}_{1}=Z_{1}^{1}\cup W_{1}^{\prime}$ and $W_{1}^{\prime}\subseteq V_{q_{1}}$, it follows that
$\alpha_{1}=\beta_{1}=0,\quad\text{and}\quad e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}^{(a+1)}(|Z_{1}^{1}\cap V_{1}|,\dots,|Z_{1}^{1}\cap V_{a}|,|W^{\prime}_{1}|)\leq z_{t}^{(a+1)}(n).$ (4.16)
Next we study $G[\tilde{U}_{i}]$ for $i\geq 2$. A key observation is that copies of $K_{1,t}$ in $G[\tilde{U}_{i}]$ together with copies of $K_{t-1,t}$ in $\tilde{U}_{1}$ may form copies of $K_{3}{(t)}$, which are restricted by Claim 4.6.
Claim 4.9.
Suppose $b>0$ and $i\in[2,r]$.
  1. (1)
    If there is a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$, then there exists $l\in[r]\setminus\{i\}$ such that the vertex set of $K_{1,t}$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$.
  2. (2)
    $\tilde{Z}_{i}\setminus Z_{1}$ and $Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ are $K_{1,t}$-free.
  3. (3)
    If $t<a$, then $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$ is $K_{1,t}$-free.
Proof.
For Part (1), let $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$. Since $B\cap(Z_{1}\cup V_{q_{1}})=\emptyset$ and $\tilde{U}_{1}\subseteq Z_{1}\cup V_{q_{1}}$, by (Deg), all vertices of $B$ have at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}$. Letting $N:=\tilde{U}_{1}\cap\bigcap_{w\in B}N(w)$, we have $|N|\geq|\tilde{U}_{1}|-(t+1)2\varepsilon n$.
First assume that $N$ is $K_{t-1,t}$-free and thus $e_{G}(N)=O(n^{2-1/(t-1)})$. Note that, since $|\tilde{U}_{1}\setminus N|\leq(t+1)2\varepsilon n$, the edges in $\tilde{U}_{1}$ incident to $\tilde{U}_{1}\setminus N$ can be split into $a+1$ bipartite $K_{t,t}$-free graphs each with one part of size at most $(t+1)2\varepsilon n$ and the other part of size at most $an$. Thus, the number of such edges is $O(\varepsilon n^{2-1/t})$. This gives $f_{1}=O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})$, contradicting (4.14).
We thus assume $N$ contains a copy of $K_{t-1,t}$. Together with $B$, they form a copy of $K_{3}{(t)}$ in $G[\tilde{U}_{1}\cup\tilde{U}_{i}]$ and we denote its vertex set by $B^{\prime}$. By Claim 4.6, there exists $l\notin\{1,i\}$ such that $B^{\prime}$ intersects $V_{q_{l}}$ and every cluster of $Z_{l}$. By Claim 4.8, $\tilde{U}_{1}\cap U_{l}=\emptyset$, so $B^{\prime}\cap Z_{l}=B\cap Z_{l}$ and $B$ indeed intersects every cluster of $Z_{l}$. Since $\tilde{U}_{i}\cap Z_{l}\supseteq B\cap Z_{l}\neq\emptyset$, we infer that $|W_{l}|\geq(1-2\varepsilon)n$ from $(Q3)$, which implies that $q_{l}\neq q_{1}$ because of (4.15). It follows that $W_{1}\cap V_{q_{l}}=\emptyset$ and thus $B\cap V_{q_{l}}=B^{\prime}\cap V_{q_{l}}\neq\emptyset$, as desired.
For Part (2), let $A_{i}:=Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ and $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{Z}_{i}\setminus Z_{1}$ or in $A_{i}$. Then, by the first part of the claim, there exists $l\in[r]\setminus\{i\}$ such that $B$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$. This is impossible if $B\subseteq A_{i}$ because $A_{i}\cap Z_{l}=\emptyset$ for any $l\notin\{1,i\}$, and also impossible if $B\subseteq\tilde{Z}_{i}\setminus Z_{1}$ because in which case $B\cap W=\emptyset$ and thus $B\cap V_{q_{l}}=\emptyset$ for any $l\notin\{1,i\}$.
Part (3) follows from Part (1) immediately. ∎
The next claim bounds $e_{G}(Z_{1}^{i})$ for $i>1$.
Claim 4.10.
For every $i\in[2,r]$, $e_{G}(Z_{1}^{i})\leq\binom{a}{2}(t-1)|Z_{1}^{i}|$.
Proof.
Suppose $|W_{1}|\geq(1-\varepsilon-\sqrt{\gamma})n$ (otherwise $Z_{1}^{i}=\emptyset$ for $i>1$ by $(Q2)$ and there is nothing to prove). Fix $i\in[2,r]$. Suppose $a\geq 2$ (if $a=1$ then $e_{G}(Z_{1}^{i})=0$). We claim that for distinct $i_{1},i_{2}\in[a]$, there can not be two copies of $K_{1,t}$ in the bipartite graph $G[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]$, one centered in $Z_{1}^{i}\cap V_{i_{1}}$ and the other centered in $Z_{1}^{i}\cap V_{i_{2}}$. We show this for $i_{1}=1$ and $i_{2}=2$. Suppose $w\in Z_{1}^{i}\cap V_{1}$ and $u_{1},\dots,u_{t}\in Z_{1}^{i}\cap V_{2}$ form a copy of $K_{1,t}$ while $w^{\prime}\in Z_{1}^{i}\cap V_{2}$, $u_{1}^{\prime},\dots,u_{t}^{\prime}\in Z_{1}^{i}\cap V_{1}$ form the other copy of $K_{1,t}$. By (Deg), $w,u_{1}^{\prime},\dots,u_{t}^{\prime}$ each has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}\setminus V_{1}$ and $w^{\prime},u_{1},\dots,u_{t}$ each has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}\setminus V_{2}$. Let
$N_{1}=\left(\tilde{U}_{1}\setminus(V_{1}\cup V_{2})\right)\cap N(w)\cap\bigcap_{i\in[t]}N(u_{i})\quad\text{ and }\quad N_{2}=\left(\tilde{U}_{1}\setminus V_{2}\right)\cap\bigcap_{i\in[t]}N(u_{i})$
(thus $N_{1}\cap N_{2}\subseteq\tilde{U}_{1}$). By $(P3)$ and $|W_{1}\cap W_{1}^{\prime}|\geq(1-\varepsilon-\sqrt{\gamma})n-2r\gamma n\geq(1-2\varepsilon)n$, we have
$|N_{1}|\geq(1-\sqrt{\gamma})an-2n+(1-2\varepsilon)n-(t+1)2\varepsilon n\geq(a-1)n-(2t+5)\varepsilon n$
and similarly $|N_{2}|\geq an-(2t+3)\varepsilon n$. If $G[N_{1},N_{2}]$ contains a copy of $K_{t-1,t}$ with the part of size $t$ in $N_{1}$, then it together with $\{w,u_{1},\dots,u_{t}\}$ forms a copy of $K_{3}{(t)}$ in $\tilde{U}_{1}\cup\tilde{U}_{i}$. Let $B$ be the vertex set of this $K_{3}{(t)}$ and note that $B\subseteq U_{1}$. Thus $B\cap Z_{l}=\emptyset$ for $l\neq 1$ contradicting Claim 4.6. This implies that $e_{G}(N_{1},N_{2})=O(n^{2-1/(t-1)})$. Now let $N_{1}^{\prime}=(\tilde{U}_{1}\setminus(V_{1}\cup V_{2}))\cap N(w^{\prime})\cap\bigcap_{i\in[t]}N(u_{i}^{\prime})$ and $N_{2}^{\prime}=(\tilde{U}_{1}\cap V_{2})\cap\bigcap_{i\in[t]}N(u_{i}^{\prime})$. By the same argument as above, $G[N_{1}^{\prime},N_{2}^{\prime}]$ contains no copy of $K_{t-1,t}$ with the part of size $t$ in $N_{1}^{\prime}$, which implies that $e_{G}(N_{1}^{\prime},N_{2}^{\prime})=O(n^{2-1/(t-1)})$. Note that all but $O(\varepsilon n^{2-1/t})$ edges of $G[\tilde{U}_{1}]\setminus G[V_{1},V_{2}]$ are in $G[N_{1}^{\prime},N_{2}^{\prime}]$ or $G[N_{1},N_{2}]$. We thus derive that $f_{1}\leq z_{t}^{(2)}(n)+O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})<z_{t}(n)-\varepsilon n^{2-1/t}$ by ($E1$), contradicting (4.14).
The conclusion above implies that $e_{G}[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]\leq(t-1)\max\{|Z_{1}^{i}\cap V_{i_{1}}|,|Z_{1}^{i}\cap V_{i_{2}}|\}$. Therefore,
$e_{G}(Z_{1}^{i})=\sum_{i_{1},i_{2}\in[a]}e_{G}[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]\leq\sum_{i_{1},i_{2}\in[a]}(t-1)|Z_{1}^{i}|\leq\binom{a}{2}(t-1)|Z_{1}^{i}|.\qed$
Claims 4.9 (2) and 4.10 together give $e_{G}(Z_{j}^{i})\leq\max\{(t-1)|Z_{j}^{i}|,\binom{a}{2}(t-1)|Z_{j}^{i}|\}=\binom{a}{2}(t-1)|Z_{j}^{i}|$ whenever $i\neq j$. Thus, by $(P3)$,
$\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})\leq r(r-1)\binom{a}{2}(t-1)\sqrt{\gamma}n\leq{a}^{2}r^{2}t\sqrt{\gamma}n$ (4.17)
Together with $f_{1}\leq z^{(a+1)}_{t}(n)$ and $f_{i}\leq(t-1)|\tilde{U}_{i}|$ for $i\geq 2$, we derive from (4.9) that
$e(G)=t_{r}(k)n^{2}+z^{(a+1)}_{t}(n)+O(n).$
This concludes the proof of Theorem 5 (note that we have not used  ($E2$) or ($E3$)).
Proof of Theorem 6
We refine our earlier estimates and prove Theorem 6. Since $\sum_{i\neq j}e_{G}(Z_{j}^{i})\leq a^{2}r^{2}t\sqrt{\gamma}n$ and $f_{j}\leq(t-1)|\tilde{U}_{j}|$ for $j\in[2,r]$, we have the following bound better than (4.14),
$f_{1}\geq z_{t}(n)-3rC_{0}kn.$ (4.18)
Furthermore, we claim that
$\text{if}\quad b>0,\quad\text{then}\quad|W_{1}^{\prime}|\geq(1-\gamma)n\quad(\text{and thus}\quad 1\in L).$ (4.19)
Indeed, by Claim 4.8, if $|W_{1}^{\prime}|<(1-\gamma)n$, then we have $f_{1}\leq z_{t}^{(a+1)}(n,\dots,n,|W_{1}^{\prime}|)\leq z_{t}^{(a+1)}(n)-\delta n^{2-1/t}$ for some $\delta>0$ by ($E2$). This contradicts (4.18).
Next we show that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$.
Claim 4.11.
Suppose $v_{0}\in V(G)$ and $i\in[r]$ satisfy that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$. Then $v_{0}$ has less than $\varepsilon n$ neighbors in $U_{i}$. In particular, we have $Z_{i}^{\prime\prime}=\emptyset$ for all $i\in[r]$ and $W^{\prime\prime}=\emptyset$.
Proof.
The second part of the claim follows immediately from the definitions of $Z_{i}^{\prime\prime}$ and $W^{\prime\prime}$.
Suppose to the contrary, that there exist $v_{0}\in V(G)$ and $i\in[r]$ such that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$ and at least $\varepsilon n$ neighbors in $U_{i}$. Since $|Z_{j}^{j}|\geq(1-\sqrt{\gamma})an$ for all $j\in[r]$, there exist sets $N_{1},\dots,N_{r-1}$ each of size $\varepsilon n-\sqrt{\gamma}an$ such that $N_{j}\subseteq Z_{j}^{j}\cap N(v_{0})$ for $j\neq i$ and $N_{i}\subseteq(Z_{i}^{i}\cup W_{i})\cap N(v_{0})$. Recall that $W_{1}^{\prime}=W_{1}^{\prime}\cap V_{q_{1}}$. By averaging, there exists $N_{1}^{\prime}\subseteq N_{1}$ with $|N_{1}^{\prime}|\geq(\varepsilon n-\sqrt{\gamma}an-2r\gamma n)/(a+1)\geq\varepsilon n/(a+2)$ such that all vertices of $N_{1}^{\prime}$ are in $Z_{1}^{1}\cup W_{1}^{\prime}$ and from the same cluster, that is,
$N_{1}^{\prime}\subseteq Q,\text{ where }Q\in\{V_{1}\cap Z_{1}^{1},V_{2}\cap Z_{1}^{1},\dots,V_{a}\cap Z_{1}^{1},W_{1}^{\prime}\}.$
Note that $N_{1}^{\prime}\subseteq W_{1}^{\prime}$ is possible only if $i=1$. If $i\neq 1$, then let $N_{i}^{\prime}:=N_{i}\setminus((W_{i}\setminus W_{i}^{\prime})\cup V_{q_{1}})$ and for every $j\in[r]\setminus\{1,i\}$, let $N_{j}^{\prime}:=N_{j}$. By $(P1)$, $|W_{i}\setminus W_{i}^{\prime}|\leq 2r\gamma n$, and by (4.19), $|W_{i}\cap V_{q_{1}}|\leq\gamma n$. Thus, we have $|N_{j}^{\prime}|\geq\varepsilon n/(a+2)$ for all $j\in[r]$. Because the sets $N_{j}^{\prime}$ are small, we can not apply the degree conditions (Deg) to them and instead, we use (D0).
Recall that $B_{1}$ is given by (D0). Next we show that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}$. Suppose instead, there is such a copy of $K_{t-1,t}$, with parts denoted by $A$ and $B$, such that $|A|=t$ and $A\subseteq N_{1}^{\prime}\setminus B_{1}$ and $B\subseteq\tilde{U}_{1}\setminus B_{1}$. Recall that $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ and for each $j\in[r]\setminus\{1,i\}$, $N_{j}^{\prime}\subseteq Z_{j}^{j}$. Observe that for every $v\in\tilde{U}_{1}\setminus B_{1}$, we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$. Indeed, if $j\neq i$, then $N_{j}^{\prime}\subseteq Z_{j}^{j}$ and we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$ by (D0); otherwise note that $N_{i}^{\prime}\subseteq Z_{i}^{i}\cup(W_{i}^{\prime}\cap W_{i})$, and by (D0) and $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ we have $d(v,N_{i}^{\prime})\geq|N_{i}^{\prime}|-\sqrt{\gamma}n$. Therefore, we obtain that the vertices in $A\cup B$ have at least $|N_{j}^{\prime}|-(2t-1)\sqrt{\gamma}n\geq(1-\gamma^{1/3})|N_{j}^{\prime}|$ common neighbors in each $N_{j}^{\prime}$, $j\in[2,r]$. Because each pair $N_{j}^{\prime},N_{j^{\prime}}^{\prime}$ has a high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these common neighborhoods, which together with $A\cup B\cup\{v_{0}\}$ form a copy of $K_{r+1}(t)$, a contradiction.
Now given that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we shall give a refined estimate on $f_{1}$. Indeed, for each $V_{j}$, $j\in[a]$, we know that $G[N_{1}^{\prime}\setminus B_{1},(V_{j}\cap Z_{1}^{1})\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, implying that $e_{G}(N_{1}^{\prime}\setminus B_{1},(V_{j}\cap Z_{1}^{1})\setminus B_{1})=O(n^{2-1/(t-1)})$. Similarly we also have $e_{G}(N_{1}^{\prime}\setminus B_{1},W_{1}^{\prime}\setminus B_{1})=O(n^{2-1/(t-1)})$. Suppose $N_{1}^{\prime}\subseteq V_{q}$ for some $q\in[a]\cup\{q_{1}\}$, then we have
$E(G[\tilde{U}_{1}])=E(G[\tilde{U}_{1}\setminus(N_{1}^{\prime}\setminus B_{1})])\cup E(G[N_{1}^{\prime}\setminus B_{1},\tilde{U}_{1}\setminus(B_{1}\cup V_{q})])\cup E(G[N_{1}^{\prime}\setminus B_{1},B_{1}\cap\tilde{U}_{1})]).$
Recall that $|N_{1}^{\prime}|\geq\varepsilon n/(a+2)$ and $|B_{1}|\leq 2\sqrt{\gamma}n$. Therefore, we can bound $f_{1}\leq|E(G[\tilde{U}_{1}])|$ by
$f_{1}\leq z_{t}\left((1-\tfrac{\varepsilon}{a+2})n,n,\dots,n\right)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})<z_{t}(n)-3rC_{0}kn,$
where we used ($E2$) and $\gamma\ll\varepsilon$. This contradicts (4.18). ∎
When $b=0$, since $Z^{\prime\prime}_{i}=\emptyset$ for all $i\in[r]$, we can improve  (4.13) to $e(G)\leq t_{r}(k)n^{2}+z_{t}^{(a)}(n)+(r-1)\lfloor\frac{t-1}{2}an\rfloor$, proving Theorem 6 for $b=0$.
In the rest of the proof, we assume $b>0$. We start with the following claim.
Claim 4.12.
For $i\in[2,r]$ such that $|W_{i}|\geq 2\varepsilon n$, we have $\tilde{U}_{i}\subseteq U_{i}\cup V_{q_{i}}$.
Proof.
Suppose instead, for some $i_{0}\in[2,r]$ with $|W_{i_{0}}|\geq 2\varepsilon n$, there exists $v\in\tilde{U}_{i_{0}}\setminus(U_{i_{0}}\cup V_{q_{i_{0}}})$. By $(P4)$ and the fact that $v\in\tilde{U}_{i_{0}}\setminus U_{i_{0}}$, we infer that $d(v,Z_{j})\geq\varepsilon n$ for all $j\neq{i_{0}}$. Then, by Claim 4.11, we have $d(v,U_{i_{0}})<\varepsilon n$. Consequently, $d(v,W_{i_{0}}^{\prime}\cap W_{i_{0}})<\varepsilon n$, namely, $v$ has at least $2\varepsilon n-2\gamma n-\varepsilon n\geq(3/4)\varepsilon n$ non-neighbors in $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ (in $G$). Note that $v$ is adjacent to all the vertices of $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ in $T$. Since $G[W^{\prime}_{i_{0}}]\subseteq T[W^{\prime}_{i_{0}}]$, we infer that
$e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})\leq e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(3/4)\varepsilon n.$
Since $e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})$ and $\alpha_{i_{0}}=|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i}^{\prime}|+e_{T}(W_{i_{0}}^{\prime})+e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$, we have
$e_{G}(\tilde{U}_{i_{0}}\setminus Z_{i_{0}})$ $=e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})$
$\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})+e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(3/4)\varepsilon n$
$\leq\alpha_{i_{0}}+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})-(3/4)\varepsilon n.$ (4.20)
Combining (4.5) and (4.20) gives $\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})\leq\sum_{i\in[r]}\alpha_{i}+\sum_{i,j\neq i}e_{G}(Z_{j}^{i})-(3/4)\varepsilon n$. Since $\sum_{i,j\neq i}e_{G}(Z_{j}^{i})\leq a^{2}r^{2}t\sqrt{\gamma}n$ by (4.17) and $\gamma\ll\varepsilon$, it follows that
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})\leq\sum_{i\in[r]}\alpha_{i}-\varepsilon n/2$ (4.21)
Recall that $f_{1}\leq z_{t}(n)$ by (4.16). We next bound $f_{i}$ for $i\geq 2$, noting that
$f_{i}=e\left(Z_{i}^{i},G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]\right)+e_{G}\left(Z_{i}^{i},(\tilde{Z}_{i}\setminus Z_{i})\cup(W^{\prime}_{i}\cap V_{q_{1}})\right).$
By Claim 4.9 (2), $G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]$ has the maximum degree at most $t-1$, and thus,
$e\left(Z_{i}^{i},G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]\right)\leq e_{G}(Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}}))\leq\frac{t-1}{2}(|Z_{i}^{i}|+|W_{i}^{\prime}\setminus V_{q_{1}}|).$
By (4.12), $e\left(Z_{i}^{i},(\tilde{Z}_{i}\setminus Z_{i})\cup(W^{\prime}_{i}\cap V_{q_{1}})\right)\leq(t-1)\left(|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right)$. Putting these together,
$f_{i}\leq\frac{t-1}{2}(|Z_{i}^{i}|+|W^{\prime}_{i}\setminus V_{q_{1}}|)+(t-1)\left(|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right)=\frac{t-1}{2}\left(|\tilde{U}_{i}|+|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right).$
By $(P3)$ and (4.19), we have $|Z^{1}_{1}|\geq(1-\sqrt{\gamma})an$, $|W^{\prime}_{1}|\geq(1-\gamma)n$, and thus, $|\tilde{U}_{1}|\geq(a+1)n(1-\sqrt{\gamma})$. By $(P3)$ and (4.19), we also have $|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\leq(r-1)\sqrt{\gamma}n+\gamma n\leq r\sqrt{\gamma}n$. It follows that
$\sum_{i=2}^{r}f_{i}\leq\frac{t-1}{2}\left(kn-(a+1)n(1-\sqrt{\gamma})+(r-1)r\sqrt{\gamma}n\right)\leq(k-a-1)n\frac{t-1}{2}+\sqrt[3]{\gamma}n.$ (4.22)
Since $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, we rewrite (4.6) as $e(G)=e(G^{\prime})+\sum_{i\in[r]}f_{i}+e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})$. By Claim 4.4, it follows that $e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i}))$. Recall that $\sum_{i=1}^{r}\beta_{i}\leq 2(r-1)rC_{0}\sqrt{\gamma}n$ by (4.8). Together with (4.21) and(4.22), we derive that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(k-a-1)n\frac{t-1}{2}+\sqrt[3]{\gamma}n+2(r-1)rC_{0}\sqrt{\gamma}n-\varepsilon n/2<g_{2}(n,r,k,t),$
a contradiction. ∎
Next we upper bound $e(\tilde{U}_{i})$ and separate the discussions depending on whether $|W_{i}|\geq 2\varepsilon n$ or not. For $i\in[2,r]$, let $U_{i}^{*}:=\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$.
Claim 4.13.
For $i\in[2,r]$, we have
$e_{G}(\tilde{U}_{i})\leq\begin{cases}(t-1)\min\{|Z_{i}^{i}|,|W_{i}^{\prime}|\}\hfill\text{ if }a=1,\text{ and }|W_{i}|\geq 2\varepsilon n\\ |\tilde{U}_{i}|(t-1)/2\hfill\text{ if }a\geq 2\text{ and }|W_{i}|\geq 2\varepsilon n\\ (t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)+\alpha_{i}\hfill\text{ if }a=1,\text{ and }|W_{i}|<2\varepsilon n\\ \left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)\hfill\text{ if }a\geq 2\text{ and}|W_{i}|<2\varepsilon n,\end{cases}$
where $t^{\prime}:=(t-1)^{2}/(16(a-1))$. Moreover, when $|W_{i}|<2\varepsilon n$, $a\geq 2$ and $t<a$, we have $e(\tilde{U}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)$.
Proof.
Assume $|W_{i}|\geq 2\varepsilon n$. Since $|W^{\prime}_{1}\geq(1-\gamma)n$, we know that $q_{1}\neq q_{i}$. It follows from Claim 4.12 that $\tilde{U}_{i}\subseteq Z_{i}^{i}\cup(W^{\prime}_{i}\cap V_{q_{i}})$. By Claim 4.9 (2), $\tilde{U}_{i}$ is $K_{1,t}$-free and thus, $e_{G}(\tilde{U}_{i})\leq|\tilde{U}_{i}|(t-1)/2$. For $a=1$, since both $Z_{i}^{i}$ and $W_{i}^{\prime}\subseteq V_{q_{i}}$ are independent sets, it follows that $e_{G}(\tilde{U}_{i})\leq(t-1)\min\{|Z_{i}^{i}|,|W_{i}^{\prime}|\}$.
Now we consider the case when $|W_{i}|<2\varepsilon n$. Recall that
$e_{G}(\tilde{U}_{i})=f_{i}+e_{G}((Z^{\prime}_{i}\setminus Z_{i})\cup W^{\prime}_{i}).$
When $a=1$, we have $e(Z_{j}^{i})=0$ for all $j\in[r]$. Thus $f_{i}\leq(t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)$ by (4.12) and $e_{G}((Z^{\prime}_{i}\setminus Z_{i})\cup W^{\prime}_{i})\leq\alpha_{i}$ by (4.5). As a result, $e_{G}(\tilde{U}_{i})\leq(t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)+\alpha_{i}$ as desired.
Now assume $a\geq 2$. Let $Z^{*}_{i}:=(\tilde{Z}_{i}\setminus Z_{i})\setminus Z_{1}^{i}$, and $W^{*}_{i}:=W_{i}^{\prime}\setminus V_{q_{1}}$. For every $j\in[r]\setminus\{1,i\}$, consider the family ${\mathcal{F}}_{j}:=\{Z_{j}^{i}\cap V_{(j-1)a+1},\dots,Z_{j}^{i}\cap V_{ja},W^{*}_{i}\cap V_{q_{j}}\}$. Let $B_{j}$ be the smallest set in ${\mathcal{F}}_{j}$ and $A_{j}$ be an arbitrary set in ${\mathcal{F}}_{j}\setminus\{B_{j},W^{*}_{i}\cap V_{q_{j}}\}$. Further, define $A:=\bigcup_{j\in[r]\setminus\{1,i\}}A_{j}$ and $B:=\bigcup_{j\in[r]\setminus\{1,i\}}B_{j}$. By Claim 4.9, both $U^{*}_{i}\setminus A$ and $U^{*}_{i}\setminus B$ are $K_{1,t}$-free. This implies that $e_{G}(U^{*}_{i}\setminus B)\leq\left\lfloor|U^{*}_{i}\setminus B|\frac{t-1}{2}\right\rfloor$ and $e(B,G[U^{*}_{i}\setminus A])\leq|B|(t-1)$. Together with $e_{G}(A,B)\leq|A||B|$, we obtain that
$e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}\setminus B|\frac{t-1}{2}\right\rfloor+|B|(t-1)+|A||B|=\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+|B|\left(\frac{t-1}{2}+|A|\right).$
Since, by definition $|B|\leq(|Z^{*}_{i}|-|A|)/(a-1)=:x$, unless $\frac{t-1}{2}+|A|-|Z^{*}_{i}|<0$, we have
$|B|\left(\frac{t-1}{2}+|A|-|Z^{*}_{i}|\right)\leq x\left(\frac{t-1}{2}-(a-1)x\right)\leq\frac{(t-1)^{2}}{16(a-1)}=t^{\prime}.$
Trivially $|A|\leq|Z_{i}^{*}|$ and $|B|\leq|W^{*}_{i}|$. Altogether, we get
$e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+|Z^{*}_{i}||W^{*}_{i}|.$ (4.23)
Let $X_{i}:=Z_{1}^{i}\cup(W_{i}^{\prime}\cap V_{q_{1}})$ and we bound $e(X_{i},G[\tilde{U}_{i}])$ as follows. Since $\tilde{U}_{i}=\tilde{Z}_{i}\cup W^{\prime}_{i}=Z_{i}^{i}\cup X_{i}\cup Z^{*}_{i}\cup W^{*}_{i}$, we have
$e(X_{i},G[\tilde{U}_{i}])$ $=e_{G}(X_{i},Z_{i}^{i})+e(Z_{1}^{i},G[Z_{1}^{i}\cup Z^{*}_{i}\cup W^{\prime}_{i}])+e(W^{\prime}_{i}\cap V_{q_{1}},G[Z^{*}\cup W^{\prime}_{i}])$
$=e_{G}(X_{i},Z_{i}^{i})+e_{G}(Z_{1}^{i})+e_{G}(Z_{1}^{i},Z^{*}_{i}\cup W^{\prime}_{i})+e_{G}(W^{\prime}_{i}\cap V_{q_{1}},Z^{*}_{i}\cup W^{*}_{i}).$
Note that $e_{G}(Z_{1}^{i},Z^{*}_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$, $e_{G}(W^{\prime}_{i}\cap V_{q_{1}},W^{*}_{i})\leq e_{T}(W^{\prime}_{i})$, and
$e_{G}(Z^{i}_{1},W^{\prime}_{i})+e(W^{\prime}_{1}\cap V_{q_{1}},Z^{*}_{i})\leq|Z^{i}_{1}||W^{\prime}_{i}|+|W^{\prime}_{1}\cap V_{q_{1}}||Z^{*}_{i}|=|\tilde{Z}_{i}\setminus Z_{i}||W^{\prime}_{i}|-|Z^{*}||W^{*}|.$
Recall that $\alpha_{i}=|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|+e_{T}(W_{i}^{\prime})+e_{T}(\tilde{Z}_{i}\setminus Z_{i})$. It follows that
$e_{G}(Z_{1}^{i},Z^{*}_{i}\cup W^{\prime}_{i})+e_{G}(W^{\prime}_{i}\cap V_{q_{1}},W^{*}_{i}\cup Z^{*}_{i})$ $\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+e_{T}(W^{\prime}_{i})+|\tilde{Z}_{i}\setminus Z_{i}||W^{\prime}_{i}|-|Z^{*}||W^{*}|$
$=\alpha_{i}-|Z^{*}||W^{*}|.$
Applying $e_{G}(X_{i},Z_{i}^{i})\leq(t-1)|X_{i}|$ from (4.12) and $e_{G}(Z_{1}^{i})\leq\binom{a}{2}(t-1)|Z_{1}^{i}|$ from Claim 4.10, we derive that
$e(X_{i},G[\tilde{U}_{i}])$ $\leq(t-1)|X_{i}|+\binom{a}{2}(t-1)|Z_{1}^{i}|+\alpha_{i}-|Z^{*}_{i}||W^{*}_{i}|.$ (4.24)
Together with (4.23), this gives the desired upper bound
$e_{G}(\tilde{U}_{i})$ $\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+(t-1)|X_{i}|+\binom{a}{2}(t-1)|Z_{1}^{i}|+\alpha_{i}$
$\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+a^{2}t|X_{i}|+\alpha_{i}.$
At last, when $a\geq 2$ and $t<a$, we know that $U^{*}_{i}$ is $K_{1,t}$-free (Claim 4.9 (3)) and thus $e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor$. Together with (4.24), this gives $e_{G}(\tilde{U}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)$. ∎
For $i\in[r]$, if $|W_{i}|\geq 2\varepsilon n$, then $\tilde{Z}_{i}\setminus Z_{i}=\emptyset$ by Claim 4.12, and therefore, $\beta_{i}=0$. If $|W_{i}|<2\varepsilon n$, then $Z_{i}\setminus\tilde{Z}_{i}=\emptyset$ by $(Q2)$; for any $j\in L\setminus\{i\}$, we have $\tilde{Z}_{j}\setminus Z_{j}=\emptyset$ and $W^{\prime}_{j}\subseteq V_{q_{i}}$ again by Claim 4.12. Hence $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0$ because $|W^{\prime}_{j}|\leq n$. Together with (4.16), for $i\in[r]$,
$\beta_{i}=\begin{cases}0&\text{ if $|W_{i}|\geq 2\varepsilon n$ (including $i=1$)},\\ \sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0&\text{ otherwise}.\end{cases}$ (4.25)
We are ready to finish our proof. Write $z_{i}:=|\tilde{Z}_{i}|$ and $w_{i}:=|W_{i}^{\prime}|$ for $i\in[r]$. Further, write $n_{l}:=|Z_{1}^{1}\cap V_{l}|$ for $\in[a]$. Then $z_{1}=\sum_{l\in[a]}n_{l}$. By (4.16), we have $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(n_{1},\dots,n_{a},w_{1})$.
We separate the cases when $a\geq 2$ and when $a=1$.
The case when $a\geq 2$
Note that there are at most $r-b$ indices $i\in[r]$ such that $|W_{i}|<2\varepsilon n$. By Claim 4.13, and using that $\sum_{i=2}^{r}|U^{*}_{i}|=(k-a-1)n$, we have
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i})\leq f_{1}+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}+\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}$
Since $e(G)=e(G^{\prime})+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})$, it follows that
$e(G)\leq f_{1}+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}+e(G^{\prime})+\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}.$
Claim 4.4 and (4.25) together imply that
$e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(\beta_{i}-\alpha_{i})\leq t_{r}(k)n^{2}-\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}.$
Together with $f_{1}\leq z_{t}(n_{1},\dots,n_{a},w_{1})$, this gives
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}.$
If $an-z_{1}+n-w_{1}>0$, then by applying ($E3$) repeatedly, we have
$z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+(r-b)t^{\prime}\leq z_{t}(n).$
This implies that $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor$. Otherwise $an+n=z_{1}+w_{1}$, which implies that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}.$
This is the desired bound for $t\geq a$. When $t<a$, the bound for $e_{G}(\tilde{U}_{i})$ given by Claim 4.13 contains no $t^{\prime}$ term. We thus obtain $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor$, regardless whether $an-z_{1}+n-w_{1}>0$.
Now consider the case $a\geq 2$ and $b=1$. If $an-z_{1}+n-w_{1}>0$, then we choose a large constant $C^{*}>0$ such that
$z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+(r-b)t^{\prime}\leq z_{t}(n)-C^{*}$
and thus $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\frac{t-1}{2}(k-a-1)n-C^{*}<g_{2}(n,r,k,t)$. Otherwise $an+n=z_{1}+w_{1}$. Since $b=1$, this forces that, for all $i\geq 2$, we have $W_{i}^{\prime}=\emptyset$ and, by $(Q2)$,$\tilde{Z}_{i}=Z_{i}$. Now, by Claim 4.9, we have $\sum_{i=2}^{r}e_{G}(\tilde{U}_{i})\leq(r-1)\lfloor\frac{t-1}{2}an\rfloor$ and consequently, $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)\lfloor\frac{t-1}{2}an\rfloor$, as desired.
The case when $a=1$
Let $L_{1}\cup L_{2}\cup L_{3}$ be a partition of $[2,r]$ such that $i\in L_{1}$ if and only if $|\tilde{Z}_{i}|<n$, $i\in L_{2}$ if and only if $|\tilde{Z}_{i}|=n$, and $i\in L_{3}$ if and only if $|\tilde{Z}_{i}|>n$. The following properties hold for $L_{1}$, $L_{2}$ and $L_{3}$.
  1. $(R1)$
    If $i\in L_{1}$, then $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$, we have $i\in L$ and, by Claim 4.12, $\tilde{Z}_{i}=Z_{i}^{i}\subsetneq V_{i}$ and $W_{i}^{\prime}\subseteq V_{q_{i}}$.
  2. $(R2)$
    If $i\in L_{2}$, then $\tilde{Z}_{i}=Z_{i}^{i}=V_{i}$. Otherwise $\tilde{Z}_{i}\neq Z_{i}^{i}$, then $|Z_{i}^{i}|<n$ and $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$ and (4.12), we have $\tilde{Z}_{i}=Z_{i}^{i}$, a contradiction.
  3. $(R3)$
    If $i\in L_{3}$, then $\tilde{Z}_{i}\not\subseteq Z_{i}$. By Claim 4.12, we have $|W_{i}|<2\varepsilon n$, which implies that $Z_{i}^{j}=\emptyset$ for $j\neq i$ by $(Q2)$. Thus, $Z_{i}^{i}=V_{i}\subsetneq\tilde{Z}_{i}$.
By $(R1)$, for every $i\in L_{1}$, $G[\tilde{U}_{i}]$ is a bipartite graph of maximum degree at most $t-1$, and thus $e_{G}(\tilde{U}_{i})\leq(t-1)\min\{z_{i},w_{i}\}$. For $i\in L_{2}$, since $\tilde{Z}_{i}=V_{i}$ is an independent set, we have $e_{G}(\tilde{U}_{i})\leq(t-1)w_{i}+e_{G}(W_{i}^{\prime})\leq(t-1)w_{i}+\alpha_{i}$. For $i\in L_{3}$, using (4.5) and the independence of $Z_{j}^{i}$ for all $j\in[r]$, we have
$e_{G}(\tilde{U}_{i})$ $=e_{G}((\tilde{Z}_{i}\setminus V_{i})\cup W_{i}^{\prime})+e_{G}((\tilde{Z}_{i}\setminus V_{i})\cup W_{i}^{\prime},Z_{i}^{i})$
$\leq\alpha_{i}+(t-1)(|\tilde{Z}_{i}\setminus V_{i}|+|W_{i}^{\prime}|)=\alpha_{i}+(t-1)(z_{i}-n+w_{i}).$
Using $\sum_{i\in L_{1}\cup L_{3}}(z_{i}-n)=n-z_{1}$ and $\sum_{i=2}^{r}w_{i}=bn-w_{1}$, we derive that
$\sum_{i\in L_{1}}\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}}w_{i}+\sum_{i\in L_{3}}(z_{i}-n+w_{i})$
$=$ $bn-w_{1}+\sum_{i\in L_{1}}(\min\{z_{i},w_{i}\}-w_{i})+\sum_{i\in L_{3}}(z_{i}-n)$
$=$ $bn-w_{1}+\sum_{i\in L_{1}}(z_{i}-n+\min\{n-w_{i},n-z_{i}\})+\sum_{i\in L_{3}}(z_{i}-n)$
$=$ $bn-w_{1}+n-z_{1}+\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}.$
Therefore,
$\sum_{i=2}^{r}e_{G}(\tilde{U}_{i})$ $\leq\sum_{i\in L_{1}}(t-1)\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}}((t-1)w_{i}+\alpha_{i})+\sum_{i\in L_{3}}((t-1)(z_{i}-n+w_{i})+\alpha_{i})$
$=(t-1)(bn-w_{1}+n-z_{1})+\sum_{i\in L_{1}}(t-1)\min\{n-w_{i},n-z_{i}\}+\sum_{i\in L_{2}\cup L_{3}}\alpha_{i}.$
Recall that $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\geq 2}(\beta_{i}-\alpha_{i})$ by Claim 4.4 and (4.16). For $i\in L_{1}\cup L_{2}$, we have $\tilde{Z}_{i}=Z^{i}_{i}$ and thus $\beta_{i}=0$ by (4.25). It follows that (note that $L\cap L_{3}=\emptyset$)
$\sum_{i\geq 2}\beta_{i}=\sum_{i\in L_{3}}\beta_{i}=\sum_{i\in L_{3}}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}\sum_{i\in L_{3}\setminus\{j\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}(n-z_{j})(w_{j}-n)$
Note that $1\in L$ by (4.19) and $(n-z_{1})(w_{1}-n)\leq 0$ by Claim 4.8. Furthermore, since $n-z_{j}=0$ for $j\in L_{2}$, it follows that $\sum_{i\geq 2}\beta_{i}=\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)$. Consequently, $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)-\sum_{i\geq 2}\alpha_{i}$.
Recall that $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(z_{1},w_{1})$. By ($E3$), we have $z_{t}(z_{1},w_{1})+(t-1)(n-z_{1}+n-w_{1})\leq z_{t}(n)$. Thus, combining these estimates together, we get
$e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+y,$
where $y:=\sum_{i\in L_{1}}((t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i}))$. For each $i\in L_{1}$, let $y_{i}:=\min\{n-w_{i},n-z_{i}\}$ and $y_{i}^{\prime}:=\max\{n-w_{i},n-z_{i}\}$. Then $y_{i}\leq y_{i}^{\prime}$ and thus,
$(t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i})=y_{i}(t-1-y_{i}^{\prime})\leq\lfloor(t-1)^{2}/4\rfloor.$
Since $L_{1}\subseteq L\setminus\{1\}$, we have $|L_{1}|\leq b-1$. Consequently, $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+(b-1)\lfloor(t-1)^{2}/4\rfloor=g_{2}(n,r,k,t)$. ∎