### v2

# Turán number of complete multipartite graphs in multipartite graphs

Abstract.
In this paper we study a multi-partite version of the Erdős–Stone theorem. Given integers $r<k$ and $t\geq 1$, let $\operatorname{ex}_{k}(n,K_{r+1}(t))$ be the maximum number of edges of $K_{r+1}(t)$-free $k$-partite graphs with $n$ vertices in each part, where $K_{r+1}(t)$ is the complete $(r+1)$-partite graph with $t$ vertices in each part. We determine the exact value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$ for $t\leq 3$, $r<k\leq 2r$ and sufficiently large $n$. We also characterize all extremal graphs for $r,k$ such that $r$ divides $k$, analogous to a result of Erdős and Simonovits on forbidding $K_{r+1}(t)$ in general graphs.
Jie Han is partially supported by a Simons Collaboration Grant 630884 and National Natural Science Foundation of China (12371341). Yi Zhao is partially supported by NSF grant DMS 2300346 and Simons Collaboration Grant 710094.
1. Introduction
Generalizing Mantel’s theorem from 1907 [Mantel], Turán’s theorem from 1941 [MR0018405] started the systemetic study of Extremal Graph Theory. Given a graph $F$, let $\operatorname{ex}(n,F)$ denote the largest number of edges in a graph not containing $F$ as a subgraph (called $F$-free). Let $K_{r}$ denote the complete graph on $r$ vertices and $T_{r}(n)$ denote the complete $r$-partite graph on $n$ vertices with $\lfloor n/r\rfloor$ or $\lceil n/r\rceil$ in each part (known as the Turán graph); and $t_{r}(n)$ be the size of $T_{r}(n)$. Turán’s theorem [MR0018405] states that $\operatorname{ex}(n,K_{r+1})=t_{r}(n)$ for all $n\geq r\geq 1$ and in addition, $T_{r}(n)$ is the unique extremal graph.
Let $K_{t_{1},\dots,t_{r}}$ denote the complete $r$-partite graph with parts of size $t_{1},\dots,t_{r}$ and write $K_{r}(t)=K_{t,\dots,t}$ with $r$ parts. A celebrated result of Erdős and Stone [MR0018807] determines $\operatorname{ex}(n,K_{r+1}(t))$ asymptotically:
$\operatorname{ex}(n,K_{r+1}(t))=t_{r}(n)+o(n^{2})=\left(1-\frac{1}{r}\right)\frac{n^{2}}{2}+o(n^{2}).$
Erdős [MR0232703] and Simonovits [MR0233735] independently improved the error term above to $O(n^{2-1/t})$. Simonovits [MR0233735] also showed that any extremal graph for $K_{r+1}(t)$ can be obtained from $T_{r}(n)$ by adding or removing $O(n^{2-1/t})$ edges. Later Erdős and Simonovits [MR0292707] determined the structure of extremal graphs for $K_{r+1}(t)$ for $t\leq 3$ as follows.
Theorem 1.
[MR0292707] For $t\leq 3$, every extremal graph $G$ for $K_{r+1}(t)$ has a vertex partition $U_{1},\dots,U_{r}$ such that
  • $G[U_{i},U_{j}]$ is complete for all $i\neq j$,
  • $G[U_{i}]=n/r+o(n)$,
  • $G[U_{1}]$ is extremal for $K_{t,t}$, and
  • $G[U_{2}],\dots,G[U_{r}]$ are extremal for $K_{1,t}$.
The restriction $t\leq 3$ in Theorem 1 comes from our knowledge on $\operatorname{ex}(n,K_{t,t})$. A well-known open problem in Extremal Graph Theory is proving $\operatorname{ex}(n,K_{t,t})=\Omega(n^{2-1/t})$ and this is only known for $t\leq 3$.
Extremal problems with multipartite graphs as host graphs have been studied since 1951, when Zarankiewicz proposed the study of the largest number of edges in a $K_{s,t}$-free bipartite graph. Let $\mathcal{G}(n_{1},\dots,n_{k})$ denote the family of $k$-partite graphs with $n_{1},\dots,n_{k}$ vertices in its parts and write $\mathcal{G}_{k}(n)=\mathcal{G}(n,\dots,n)$ with $k$ parts. Given a graph $F$, define $\operatorname{ex}(n_{1},\dots,n_{k};F)$ as the largest number of edges in $F$-free graphs from $\mathcal{G}(n_{1},\dots,n_{k})$, and let $\operatorname{ex}_{k}(n,F)=\operatorname{ex}(n,\dots,n;F)$ (with $k$ parts). (Trivially $\operatorname{ex}_{k}(n,F)=\binom{k}{2}n^{2}$ if the chromatic number $\chi(F)>k$.) In 1975 Bollobás, Erdős, and Szemerédi [MR0389639] investigated several Turán-type problem for multipartite graphs. Applying a simple counting argument, they showed that, for any $n,k,r\in\mathbb{N}$ with $k>r$,
$\operatorname{ex}_{k}(n,K_{r+1})=t_{r}(k)n^{2}$ (1.1)
The main results of [MR0389639] were on the minimum degree that forces a copy of $K_{r}$ in the graphs of $\mathcal{G}_{k}(n)$. This problem has been intensively studied and resolved when $k=r$ (frequently in its complementary form concerning independent transversals) [MR2195582, MR1860440, MR1208805, MR2246152], and more recently for $k>r$ [MR4496024] (certain cases are still unsolved). There are several other extremal results for multipartite graphs. Bollobás, Erdős, and Straus [MR0379256] determined $\operatorname{ex}(n_{1},\dots,n_{r};K_{r})$ for all $n_{1},\dots,n_{r}$ and $r$. Let $kK_{r}$ denote $k$ vertex-disjoint copies of $K_{r}$. The problem $\operatorname{ex}(n_{1},\dots,n_{r};kK_{r})$ were studied more recently [MR3724116, MR4425157, MR4595892] and settled by Chen, Lu, and Yuan [CLY23] when $n_{1},\dots,n_{r}$ are sufficiently large ($k,r$ are arbitrary but fixed). The minimum pair density of multipartite graphs that forces a clique was studied by Bondy, Shen, Thomassé, and Thomassen [bondy] and Pfender [pfender].
In this paper we study $\operatorname{ex}_{k}(n,K_{r+1}(t))$, the multi-partite version of the Erdős–Stone theorem and Theorem 1. To give the precise value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$, we need the following definition. Given $a,t,n_{1},\dots,n_{a}\in\mathbb{N}$, let $z_{t}^{(a)}(n_{1},\dots,n_{a})$ be the $a$-partite Zarankiewicz number for $K_{t,t}$, that is, the maximum number of edges in a $K_{t,t}$-free bipartite graph with part sizes $n_{1},\dots,n_{a}$. For simplicity, we write $z_{t}^{(a)}(n)$ if $n_{1}=\cdots=n_{a}$. We also write $z_{t}(m,n)$ for $z_{t}^{(2)}(m,n)$.
Theorem 1 says that any extremal graph for $K_{r+1}(t)$ is the join of an extremal graph for $K_{t,t}$ and $r-1$ extremal graphs for $K_{1,t}$. Inspired by this, a natural guess of an extremal graph for its $r$-partite analogue is as follows. Suppose $k=ar+b$ with $0\leq b<r$. We start with an $n$-blowup of $T_{r}(k)$, which has $r$ classes and each class has either $a$ or $a+1$ parts. We add a $K_{t,t}$-free graph to one class with the most number of parts and $\{K_{1,t},K_{2,2}\}$-free graphs to the remaining $r-1$ classes. If $r<k\leq 2r$, then it is easy to see (as in the proof of Theorem 2) that this graph is $K_{r+1}(t)$-free and has $t_{r}(k)n^{2}+z_{t}(n,n)+(k-r-1)(t-1)n$ edges, and therefore
$\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n,n)+(k-r-1)(t-1)n.$
In this paper we first improve this lower bound when $t\geq 2$ and $r<k\leq 2r$.
Theorem 2.
Suppose $t\geq 2$, $r<k\leq 2r$ and $n\geq 8t^{2}$. Then $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq g(n,r,k,t)$, where
$g(n,r,k,t):=t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(k-r-1)n+\min\{k-r-1,2r-k\}\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
The following is our main result, in which we prove a matching upper bound when $t\in\{2,3\}$ and $n$ is sufficiently large.
Theorem 3.
For any $t\in\{2,3\}$ and integers $r$, there exists $n_{0}=n_{0}(r)\in\mathbb{N}$ such that for $n\geq n_{0}$ and $r<k\leq 2r$, we have $\operatorname{ex}_{k}(n,K_{r+1}(t))=g(n,r,k,t)$.
In fact, we conjecture that $\operatorname{ex}_{k}(n,K_{r+1}(t))=g(n,r,k,t)$ holds for all $t\geq 2$, $r<k\leq 2r$, and sufficiently large $n$.
A natural question is whether our result can be extended to larger values of $t$ and $k$. For larger value of $t$, although we can use $z_{t}(n,n)$ without knowing its precise value, we need several properties of this function in our proof. Kővári, Sós, Turán [MR0065617] showed that $z_{t}(n,n)=O(n^{2-1/t})$ for $t\geq 2$ and proving a matching lower bound is a well-known open problem:
  1. ($Z$)
    $z_{t}(n,n)=\Omega(n^{2-1/t})$ for $t\geq 2$.
It was shown [MR0200182, MR0223262] that ($Z$) holds for $t=2,3$. In addition, we will need the following properties.
  1. ($E1$)
    For any $a\in\mathbb{N}$, there exists $\delta>0$ such that for large $n$, $z_{t}^{(a+1)}(n)-z_{t}^{(a)}(n)\geq{\delta n^{2-1/t}}.$
  2. ($E2$)
    for any $\varepsilon\in(0,1]$, there exists $\delta>0$ such that for large $n$,
    $z_{t}(n,n)-z_{t}((1-\varepsilon)n,n)\geq{\delta n^{2-1/t}}.$
  3. ($E3$)
    $z_{t}(m,n)-z_{t}(m-1,n)\geq t-1$.
We can easily verify ($E1$)($E3$) for $t=2,3$. First, a proof of ($E1$) from ($Z$) for all $t$ is given in Section 3.3. Second, when $t=2,3$, ($E2$) follows from $z_{t}(m,n)\leq(1+o(1))mn^{1-1/t}$ by Füredi [MR1395691] and $z_{t}(n,n)\geq(1-o(1))n^{2-1/t}$ for $t=2,3$. Third, ($E3$) holds trivially because adding a vertex with $t-1$ edges to a $K_{t,t}$-free graph will not create a copy of $K_{t,t}$.
We do not know whether similar properties hold when $k>2r$: in ($E2$) we must deal with the $\lceil k/r\rceil$-partite Zarankiewicz number; we also need to replace $t-1$ by $\Omega(a^{2}t)$ in ($E3$), which seems out of reach at present.
Theorems 2 and 3 show that our problem is more complex than its non-partite counterpart, Theorem 1. Finally, we show that, when $r$ divides $k$, this additional complexity does not exist, and we give an analogue of Theorem 1, modulo the existence of a set of constantly many exceptional vertices.
Theorem 4.
For $r,k\in\mathbb{N}$ with $r\mid k$ and $t=2,3$, there exist $C_{0},n_{0}\in\mathbb{N}$ such that the following holds for $n\geq n_{0}$. Let $G$ be a $K_{r+1}(t)$-free $k$-partite graph with $n$ vertices in each part and $\operatorname{ex}_{k}(n,K_{r+1}(t))$ edges. Then there is a vertex partition of $G$ into $U_{1},\dots,U_{r}$, each consisting of exactly $k/r$ parts of $G$, and a vertex set $Z\subseteq V(G)$ with $|Z|\leq C_{0}$ such that
  • $G[U_{i}\setminus Z,U_{j}\setminus Z]$ is almost complete for all $i\neq j$,
  • $G[U_{1}\setminus Z]$ is $K_{t,t}$-free, and
  • $G[U_{2}\setminus Z],\dots,G[U_{r}\setminus Z]$ are $K_{1,t}$-free.
Showing $Z=\emptyset$ in Theorem 4 requires an $a$-partite analogue of ($E2$), which is unknown for $a\geq 3$.
For the rest of this paper we only consider $r\geq 2$, as it is easy to see that Theorems 24 hold for $r=1$.
Notation. Given integers $n\geq m\geq 1$, let $[n]=\{1,\dots,n\}$ and $[m,n]=\{m,m+1,\dots,n\}$. We omit floors and ceilings unless they are crucial, e.g., we may choose a set of $\varepsilon n$ vertices even if our assumption does not guarantee that $\varepsilon n$ is an integer.
When $X,Y\subseteq V(G)$ intersect, $E_{G}(X,Y)$ is defined as the collection of ordered pairs in $(x,y)\in X\times Y$ such that $\{x,y\}\in E(G)$. Write $e_{G}(X,Y)=|E_{G}(X,Y)|$. For a vertex $v$ in $G$, let $N(v,X)=N(v)\cap X$ and $d(v,X)=|N(v,X)|$. Moreover, given $X\subseteq V(G)$, let $e(X,G)$ be the number of edges of $G$ incident to the vertices of $X$. Given two graphs $G$ and $H$ on a common vertex set $V$, $G\cap H$ denotes a graph on $V$ with $E(G\cap H)=E(G)\cap E(H)$. Given a $k$-partition $\{V_{1},V_{2},\dots,V_{k}\}$, a set $S$ is called crossing if $|S\cap V_{i}|\leq 1$, $i\in[k]$.
When we choose constants $x,y>0$, $x\ll y$ means that for any $y>0$ there exists $x_{0}>0$ such that for any $x<x_{0}$ the subsequent statement holds. Hierarchies of other lengths are defined similarly. Furthermore, all constants in the hierarchy are positive and for a constant appearing in the form $1/s$, we always mean to choose $s$ as an integer.
2. Proof of Theorem 2
In this section we prove Theorem 2, that is, $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq g(n,r,k,t)$ for $r<k\leq 2r$. Our proof needs a $t$-regular $K_{2,2}$-free bipartite graph with $n$ vertices in each part. It is well known (see [MR101250]) that such graph exists for infinitely many $n\in\mathbb{N}$ with $n\geq t^{2}$. The following proposition from [MR2519934]*Section 2 allows $n$ to be any integer that is at least $8t^{2}$.
Proposition 2.1.
[MR2519934] For $t\geq 1$ and $n\geq 8t^{2}$, there exists a $t$-regular $K_{2,2}$-free bipartite graph with $n$ vertices in each part.
Now we prove our lower bound on $\text{ex}_{k}(n,K_{r+1}(t))$ stated in Theorem 2.
Proof of Theorem 2.
First assume $r<k\leq 2r$ and $t\geq 2$. Let $V_{1},\dots,V_{k}$ be disjoint sets of size $n$ and, if $k<2r$, let $V_{k+1},\dots,V_{2r}$ be empty sets. Let $G^{\prime}$ be a complete $r$-partite graph with parts $V_{1}\cup V_{r+1},V_{2}\cup V_{r+2},\dots,V_{r}\cup V_{2r}$. Moreover, we add to $G^{\prime}$ a maximum $K_{t,t}$-free bipartite graph with bipartition $V_{1}\cup V_{r+1}$, and a $(t-1)$-regular $K_{2,2}$-free bipartite graph on $V_{i}\cup V_{i+r}$ for $2\leq i\leq k-r$ (the existence of such graph is guaranteed by Proposition 2.1). The resulting graph is $K_{r+1}(t)$-free because a copy of $K_{r+1}(t)$ has at most $12t-$ 1 vertices in $V_{1}\cup V_{r+1}$ and at most $t$ vertices in $V_{i}\cup V_{i+r}$ for $2\leq i\leq k-r$. This graph has $t_{r}(k)n^{2}+z_{t}(n)+(t-1)(k-r-1)n$ edges, and thus, $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(k-r-1)n$. This proves the theorem when $t=2$ or $k=2r$.
Now assume $r<k<2r$ and $t\geq 2$. Let $b=k-r$. Our goal is to give a better construction that shows
$\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(b-1)n+\min\{b-1,r-b\}\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
Let $V_{i,j}$, $(i,j)\in[r]\times[2]$ be vertex sets, where $V_{b+1,2},\dots,V_{r,2}$ are empty sets and other sets have size $n$. Let $G=K(V_{1,1}\cup V_{1,2},\dots,V_{r,1}\cup V_{r,2})$ be the complete $r$-partite graph with parts $V_{1,1}\cup V_{1,2},\dots,V_{r,1}\cup V_{r,2}$. Thus $e(G)=t_{r}(k)n^{2}$.
We first revise the partition as follows. Let $t^{\prime}:=\lceil(t-1)/2\rceil$ and $b^{\prime}:=\min\{b-1,r-b\}$. Let $\{V_{i,j}^{\prime},(i,j)\in[r]\times[2]\}$ be obtained from $\bigcup V_{i,j}$ by moving a set $S_{i,1}$ of $t^{\prime}$ vertices from $V_{i,1}$ to $V_{i+b-1,1}$, and moving a set $S_{i,2}$ of $t^{\prime}$ vertices from $V_{i,2}$ to $V_{i+b-1,2}$, for every $i\in[2,b^{\prime}+1]$. For $i\in[r]$, let $U_{i}:=V_{i,1}^{\prime}\cup V_{i,2}^{\prime}$ and $H:=K(U_{1},\dots,U_{r})\cap K(V_{1,1},\dots,V_{r,1},V_{1,2},\dots,V_{r,2})$. Let $H^{\prime}$ be obtained from $H$ by adding
  • a $K_{t,t}$-free bipartite graph on $U_{1}$ of size $z_{t}(n,n)$, and
  • a maximum $\{K_{1,t},K_{2,2}\}$-free bipartite graph on $U_{i}$ for $i\in[2,b]$.
  • $2t^{\prime}$ vertex-disjoint copies of $K_{1,t-1}$ on each $U_{i}$, $i\in[b+1,b+b^{\prime}]$, where the centers of stars are the $2t^{\prime}$ vertices of $S_{i-b+1,1}\cup S_{i-b+1,2}$; then add a copy of $K_{t^{\prime},t^{\prime}}$ on $S_{i-b+1,1}\cup S_{i-b+1,2}$. See Figure 1.
Figure 1. The lower bound construction for $K_{r+1}(t)$ with $r=5$, $k=8$ and $t=3$. The left figure is the standard construction similar to the one given in Theorem 1; the right figure is the construction presented in our proof of Theorem 3.
For $i\in[2,b]$, Proposition 2.1 implies that each $H^{\prime}[U_{i}]$ is $(t-1)$-regular and thus for $i\in[2,b^{\prime}+1]$, $H^{\prime}[U_{i}]$ has $(n-t^{\prime})(t-1)$ edges, and for $i\in[b^{\prime}+2,b]$ it has $n(t-1)$ edges; for $i\in[b+1,b+b^{\prime}]$, each $H^{\prime}[U_{i}]$ has $2t^{\prime}(t-1)+(t^{\prime})^{2}$ edges. Therefore, the number of edges in $H^{\prime}\setminus H$ is
$z_{t}(n,n)+(n-t^{\prime})(t-1)b^{\prime}+n(t-1)(b-b^{\prime}-1)+b^{\prime}(2t^{\prime}(t-1)+(t^{\prime})^{2})$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+b^{\prime}(t^{\prime})^{2}+b^{\prime}t^{\prime}(t-1)$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+2b^{\prime}(t^{\prime})^{2}+b^{\prime}t^{\prime}(t-1-t^{\prime})$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+2b^{\prime}(t^{\prime})^{2}+\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
We claim that $H$ contains $t_{r}(k)n^{2}-2b^{\prime}(t^{\prime})^{2}$ edges. Indeed, for every $i\in[2,b^{\prime}+1]$, the vertices of $S_{i,1}$ moved from $V_{i,1}$ to $V_{i+b-1,1}$ lose $n$ edges to $V_{i+b-1,1}$ and gain $n-t^{\prime}$ edges to $V_{i,2}$, thus having a net loss $(t^{\prime})^{2}$ edges between $U_{i}$ and $U_{i+b-1}$; the same holds for $S_{i,2}$. Thus, our claim holds after summing over all $b^{\prime}$ such rows, which implies that $e(H^{\prime})=t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(b-1)n+b^{\prime}\lfloor\frac{(t-1)^{2}}{4}\rfloor$.
At last, we show that $H^{\prime}$ is $K_{r+1}(t)$-free. Recall that by construction, every $U_{i}$ is triangle-free, $U_{1}$ is $K_{t,t}$-free, $U_{2},\dots,U_{b}$ are $\{K_{1,t},K_{2,2}\}$-free, and each of $U_{b+1},\dots,U_{U_{b+b^{\prime}}}$ induces vertex-disjoint copies of $K_{1,t-1}$ whose centers are joined by copies of $K_{t^{\prime},t^{\prime}}$. Suppose $K$ is a copy of $K_{r+1}(t)$ in $H^{\prime}$. By construction, $K$ contains at most $12t-$ 1 vertices in $U_{1}$. We claim that $|V(K)\cap(U_{i}\cup U_{i+b-1})|\leq 2t$ for $i\in[2,b^{\prime}+1]$ and $|V(K)\cap U_{j}|\leq t$ for $i\in[b^{\prime}+2,b]\cup[b+b^{\prime}+1,r]$, which will lead to a contradiction with $|V(K)|=(r+1)t$. Indeed, for $i\in[2,r]$, if $K$ contains at least $t+1$ vertices in $U_{i}$, then these vertices induce either a copy of $K_{1,t}$ or $K_{2,2}$. By construction, this is only possible for $i\in[b+1,b+b^{\prime}]$ and in that case $V(K)\cap U_{i}$ must intersect both $S_{i-b+1,1}$ and $S_{i-b+1,2}$. Furthermore, there exists $v\in V(K)\cap S_{i-b+1,1}$ and $v^{\prime}\in V(K)\cap S_{i-b+1,2}$ such that $v$ and $v^{\prime}$ are in different color classes of $K$. Since $v$ and $v^{\prime}$ have no common neighbor in $U_{i}\cup U_{i-b+1}$, $K$ induces at most two classes on $U_{i}\cup U_{i-b+1}$. Thus, $|V(K)\cap(U_{i}\cup U_{i-b+1})|\leq 2t$ and we are done. ∎
Remark. Note that for $t=2$, both our constructions give the same value, which means that the extremal graph is not unique. Indeed, it is easy to see that one can construct $b^{\prime}+1$ different ones – as we can move vertices for a subset of the $b^{\prime}$ rows.
3. Stability and proof outline
Let us first consider extremal graph for $\operatorname{ex}_{k}(n,K_{r+1})$. Given $r,k\in\mathbb{N}$ with $k>r$, write $k=ar+b$ for $10\leq b\leq r-$ 1. By Turán’s theorem, the Turán graph $T_{r}(k):=K_{a,\dots,a,a+1,\dots,a+1}$ (with $b$ parts of size $a+1$ and $r-b$ parts of size $a$) is the unique largest $K_{r+1}$-free graph on $k$ vertices. The following definition shows that there are many extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$.
Definition 3.1.
Let ${\mathcal{T}}_{r,k}(n)$ be the collection of $k$-partite graphs with parts $V_{1},\dots,V_{k}$ of size $n$ defined as follows. If $b>0$, we arbitrarily divide $V_{ar+1},\dots,V_{k}$ into $r$ sets $W_{1},\dots,W_{r}$ (some of them may be empty) such that each $W_{i}$ is a subset of $V_{j}$ for some $j$; if $b=0$, then let $W_{1},\dots,W_{r}$ be empty sets. Now let $T$ be the $r$-partite graph with parts $U_{1},\dots,U_{r}$ such that
$U_{i}=W_{i}\cup Z_{i},\text{ where }Z_{i}:=V_{(i-1)a+1}\cup\cdots\cup V_{ia},$
obtained from the complete $r$-partite graph $K(U_{1},\dots,U_{r})$ by removing edges between $W_{i}$ and $W_{i^{\prime}}$, $i\neq i^{\prime}$, whenever $W_{i},W_{i^{\prime}}\subseteq V_{j}$ for some $j$ (in other words, $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$).
Since $T$ is $r$-partite, it is $K_{r+1}$-free. Let $U=\bigcup_{i\in[r]}U_{i}$ and $W=\bigcup_{i\in[r]}W_{i}$. Note that $e_{T}(U)=\binom{r}{2}a^{2}n^{2}$ while the number of edges of $T$ incident to $W$ is equal to $|W|(k-a-1)n=b(k-a-1)n^{2}$. Since $t_{r}(k)=\binom{r}{2}a^{2}+b(k-a-1)$, it follows that $e(T)=t_{r}(k)n^{2}$. By (1.1), $T$ is an extremal graph for $K_{r+1}$. [^1] Indeed, Chen, Lu, and Yuan [CLY23] showed that ${\mathcal{T}}_{r,k}(n)$ are all the extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$ and described all extremal graphs for $\operatorname{ex}(n_{1},\dots,n_{k};K_{r+1})$.
3.1. A stability theorem
We need the following stability result for $\operatorname{ex}_{k}(n,K_{r+1}(t))$. Given two graphs $G,H\in\mathcal{G}_{k}(n)$ on the same parts $V_{1},\dots,V_{k}$, we say that $G$ and $H$ are $\gamma$-close if $|E(G)\triangle E(H)|\leq\gamma n^{2}$.
Theorem 5.
For any $k,r,t\in\mathbb{N}$ and any $\gamma>0$, there exist $\varepsilon>0$ and $n_{0}\in\mathbb{N}$ such that the following holds for every integer $n\geq n_{0}$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}(t)$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Then $G$ is $\gamma$-close to a member of ${\mathcal{T}}_{r,k}(n)$. In particular, we have $e(G)\leq(t_{r}(k)+\gamma)n^{2}$.
In the earlier version of this paper, we gave a self-contained proof of Theorems 5. Here we derive it from a stronger result of Chen, Lu and Yuan [CLY23]*Theorem 1.5, in which they provide more structural information. Their definition and result are more general by allowing the parts $V_{i}$ of $G$ to have different sizes. Here we only state their definition and result that we need for the balanced case.
Definition 3.2 (Stable partition).
Let $k\geq r\geq 2$ be integers. Let $\mathcal{P}=\{P_{1},\dots,P_{r}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{k}\}$ where $|V_{i}|=n$ for all $i\in[k]$ be partitions of a common vertex set. For $i\in[r],j\in[k]$, a set $W=P_{i}\cap V_{j}$ is called an integral part if $W=V_{j}$, and called a partial part otherwise. We say that $\mathcal{P}$ is stable to $\mathcal{V}$, if each of $P_{1},\dots,P_{r}$ has the same number of integral parts and at most one partial part.
Definition 3.3 ($(X,\varepsilon)$-stable).
Let $n,r,k$ be integers with $n\geq k\geq r\geq 2$. For a given spanning subgraph $G$ of $K_{k}(n)$, let $\mathcal{P}=\{P_{1},\dots,P_{r}\}$ be a partition of $V(G)$ and $\mathcal{V}=\{V_{1},\dots,V_{k}\}$ be the natural $k$-partition. Given $\varepsilon\in(0,1)$ and a set $X\subseteq V(G)$ of size at most $\varepsilon n$, we say that $\mathcal{P}$ is an $(X,\varepsilon)$-stable partition if
  • $G-X$ is $\varepsilon$-close to $K(P_{1},\dots,P_{r})-X$,
  • $\{P_{1}\setminus X,\dots,P_{r}\setminus X\}$ is stable to $\{V_{1}\setminus X,\dots,V_{k}\setminus X\}$.
Theorem 6.
[CLY23]*Theorem 1.5 Let $F$ be a graph with chromatic number $r+1\geq 3$. For every $\varepsilon>0$ there exists $\delta>0$ and integer $n_{0}>0$ such that the following holds for $n\geq n_{0}$. Let $G$ be an $F$-free subgraph of $K_{k}(n)$ with $k>r$ such that $e(G)\geq\operatorname{ex}_{k}(n,F)-\delta n^{2}$. Then, $G$ has an $(X,\varepsilon)$-stable partition $\{P_{1},\dots,P_{t-1}\}$ for some set $X\subseteq V(G)$ of size at most $\varepsilon n$.
Now Theorem 5 follows from Theorem 6 by setting $F=K_{r+1}(t)$ and noticing that i) a partition is stable if and only if it is a partition required in Definition 3.1 and ii) as $|X|\leq\varepsilon n$, $X$ is incident to at most $\varepsilon n^{2}$ edges that we have no control on.
We remark that although Theorem 6 is stronger, the additional structural information is not immediately useful to us. For example, the set $X$ introduced in Theorem 6 is the set of atypical vertices of $G$ (see also the proof outline in next section). However, we need a stronger control and indeed we need to distinguish two kinds of atypical vertices. Identifying them from $X$ is almost equivalent to identifying them from $V(G)$. So we choose to present and use the easier and more classical version, Theorem 5, in this paper.
3.2. Outline of the proofs.
Now we give an outline of our proofs. Let $G\in\mathcal{G}_{k}(n)$ be $K_{r+1}(t)$-free and has the maximum number of edges. Since $e(G)>t_{r}(k)n^{2}$, we can assume that $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$. Since $e(G)$ is maximum, we can easily derive a minimum degree condition by symmetrization arguments.
Next we define atypical vertices. Roughly speaking, there are two types of atypical vertices: the first type of vertices, denoted by $Z^{\prime\prime}\cup W^{\prime\prime}$, are the “wrong” ones that do not exist in ${\mathcal{T}}_{r,k}(n)$; the second type of vertices, denoted by $(W^{\prime}_{i}\setminus W_{i})\cup\bigcup_{j\neq i}Z_{j}^{i}$ for $i\in[r]$, are the vertices that are not in $U_{i}$ but behave like the vertices of $U_{i}$, in other words, they are in the wrong place. We temporarily ignore the first type of atypical vertices because there are only a constant number of them (see $(P1)$ and $(P3)$) and they contribute only $O(n)$ to $e(G)$. For the second type of atypical vertices, there are only $o(n)$ of them (see $(P1)$ and $(P3)$) and we move them to appropriate rows and redefine our partition as $\tilde{U}_{1},\dots,\tilde{U}_{r}$ (see (4.2)). A key observation is that $Z_{j}^{i}\neq\emptyset$ (namely, there is a vertex in $Z_{j}$ but behaves as a vertex in $Z_{i}$) only if $|W_{j}|\geq(1-o(1))n$.
Now we estimate $e(G)$. We split $E(G)$ into $E_{G}(\tilde{U}_{1}),\dots,E_{G}(\tilde{U}_{r})$, and $E(G^{\prime})$, where $G^{\prime}:=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have a relatively good estimate of $e(G^{\prime})$ (see Claim 4.3) taking into account that the partition is no longer balanced. In contrast, due to the second type of atypical vertices, we can only show that each $G[\tilde{U}_{i}]$ is “almost” $K_{t,t}$-free (see Claim 4.4). Similarly, we show that all but at most one rows are “almost” $K_{1,t}$-free (see Claim 4.6). Assuming that $e_{G}(\tilde{U}_{1})$ is the largest among all $e_{G}(\tilde{U}_{i})$, $i\in[r]$, we can use these properties and give an upper bound of $e(G)$. Next we show that $\tilde{U}_{1}$ has no atypical vertices (Claim 4.8), and thus $e_{G}(\tilde{U}_{1})\leq z_{t}^{(a+1)}(n)$. We further refine our estimate on $\tilde{U}_{i}$, $i>1$, and show that each second type atypical vertex contributes at most a constant number of edges to $E(G)\setminus E(G^{\prime})$ (Claim 4.9). In summary, $\tilde{U}_{1}$ is indeed $K_{t,t}$-free, $e_{G}(\tilde{U}_{i})=O(n)$ for $i>1$, and $|Z^{\prime\prime}\cup W^{\prime\prime}|=O(1)$, from which we conclude the proof of Theorem 4.
To prove Theorem 3, we refine earlier estimates as follows. We first show that $|W_{1}^{\prime}|=(1-o(1))n$ and $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, where we use ($E2$). The rest of the proofs are further refinements of our estimates. In particular, we show that if $|W_{i}^{\prime}|$ is not too small, then $\tilde{U}_{i}$ essentially contains no vertex from other rows.
3.3. Two quick proofs
Proof of ($Z$) $\Rightarrow$ ($E1$).
Given $a+1$ sets $V_{1},\dots,V_{a+1}$ of size $n$, we define an $(a+1)$-partite graph $G$ on $V_{1},\dots,V_{a+1}$ as follows. Let $V_{2}^{\prime}$ be a set of $n$ vertices consisting of $\lfloor n/2\rfloor$ vertices from $V_{1}$ and $\lceil n/2\rceil$ vertices from $V_{2}$. We place an extremal graph $G^{\prime}$ for $z_{t}^{(a)}(n)$ on $V_{2}^{\prime},V_{3},\dots,V_{a+1}$, in other words, $G^{\prime}$ is an $a$-partite $K_{t,t}$-free graph with $z_{t}^{(a)}(n)$ edges. Next we add a maximum bipartite $K_{t,t}$-free graph $G^{\prime\prime}$ on the remaining vertices of $V_{1}$ and $V_{2}$. By ($Z$), $e(G^{\prime\prime})\geq z_{t}^{(2)}(\lfloor n/2\rfloor)\geq\delta n^{2-1/t}$ for some $\delta>0$. Thus $G=G^{\prime}\cup G^{\prime\prime}$ is $K_{t,t}$-free and $e(G)=e(G^{\prime})+e(G^{\prime\prime})\geq z_{t}^{(a)}(n)+\delta n^{2-1/t}$. This gives ($E1$). ∎
We need the following simple proposition.
Proposition 3.4.
Given $r,t\in\mathbb{N}$ and reals $\gamma,\varepsilon>0$ such that $\varepsilon^{2}>3r^{2}t^{2}\gamma$, and let $n$ be sufficiently large. Suppose $G$ is a $K_{r+1}(t)$-free graph with vertex partition $V=U_{1}\cup\cdots\cup U_{r}$ such that $|U_{i}|\geq n$ for $i\in[r]$ and $d(U_{i},U_{j})\geq 1-\gamma$, $i,j\in[r]$, $i\neq j$. Let $X\subseteq V$ be the set of vertices $v$ satisfies that $d(v,U_{i})\geq\varepsilon|U_{i}|$ for all $i\in[r]$. Then $|X|\leq 2(t-1)\varepsilon^{-rt}$.
Proof.
We call a copy of $K_{r}{(t)}$ in $G$ useful if it consists of exactly $t$ vertices from each of $U_{1},\dots,U_{r}$. We first show that for every $v\in X$, $N(v)$ contains many useful copies of $K_{r}{(t)}$. Indeed, $d(U_{i},U_{j})\geq 1-\gamma$ for every $i,j\in[r]$, $i\neq j$ implies that $G[U_{i},U_{j}]$ has at most $\gamma|U_{i}||U_{j}|$ non-edges. Since $d(v,U_{i})\geq\varepsilon|U_{i}|$ for all $i\in[r]$, take $W_{i}\subseteq N(v)\cap U_{i}$ of size exactly $\varepsilon|U_{i}|$. We can find $\prod_{i\in[r]}\binom{\varepsilon|U_{i}|}{t}$ $rt$-sets which consists of $t$ vertices from each $U_{i}$, amongst which, at most
$\sum_{i,j\in[r],i\neq j}\gamma|U_{i}||U_{j}|\cdot\left(\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}\right)\cdot\frac{t}{\varepsilon|U_{i}|}\frac{t}{\varepsilon|U_{j}|}\leq\frac{r^{2}t^{2}\gamma}{\varepsilon^{2}}\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}$
of them contain crossing non-edges. Therefore, $N(v)$ contains at least
$\left(1-\frac{r^{2}t^{2}\gamma}{\varepsilon^{2}}\right)\prod_{i^{\prime}\in[r]}\binom{\varepsilon|U_{i^{\prime}}|}{t}\geq\frac{\varepsilon^{rt}}{2}\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t}$
useful copies of $K_{r}{(t)}$, where we used that $r^{2}t^{2}\gamma\varepsilon^{-2}<1/3$. Since $G$ is $K_{r+1}(t)$-free, each useful copy $K$ of $K_{r}{(t)}$ is in $N(v)$ for at most $t-1$ choices of $v\in X$. Double counting on the number of pairs $(v,K)$ such that $K\subseteq N(v)$ is useful, we obtain that
$|X|\frac{\varepsilon^{rt}}{2}\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t}\leq(t-1)\prod_{i^{\prime}\in[r]}\binom{|U_{i^{\prime}}|}{t},$
which gives $|X|\leq 2(t-1)\varepsilon^{-rt}$. ∎
4. Main Proofs
Given integers $1\leq s\leq t$ and sufficiently large $m,n$, Kővári, Sós, Turán [MR0065617] showed that $z(m,n,s,t)\leq Cmn^{1-1/s}$ for some $C=C(t)>0$, that is, a bipartite graph $G$ with parts of size $m$ and $n$ has at most $Cmn^{1-1/s}$ edges if $G$ has no copy of $K_{s,t}$ where the part of size $s$ is in the part of $G$ of size $m$. We start the proof with the general setting, that is, for $k=ar+b$ with $0\leq b<r$. After we conclude the proof of Theorem 4, we focus on the case $k\leq 2r$.
Proofs of Theorems 3 and 4.
Suppose $t=2,3$ and thus ($Z$) holds, that is, $z_{t}^{(2)}(n)\geq cn^{2-1/t}$ for some $c>0$. Take $C=C(t)$ as in the Kővári–Sós–Turán result in the previous paragraph. We choose constants
$1/n\ll\gamma\ll\varepsilon\ll\varepsilon^{\prime}\ll 1/k,1/t,c,C.$
Suppose $G$ is $K_{r+1}(t)$-free and has the maximum number of edges, that is, $e(G)=\operatorname{ex}_{k}(n,K_{r+1}(t))$. Suppose further that $e(G)>g(n,r,k,t)>t_{r}(k)n^{2}$. By Theorem 5, $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$ such that $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$, where $U_{1},\dots,U_{r}$ is a partition of $V(T)$ satisfying the following properties:
  • $U_{i}=W_{i}\cup Z_{i}$ such that $Z_{i}=V_{(i-1)a+1}\cup\cdots\cup V_{ia}$, and
  • $W_{i}=\emptyset$ if $b=0$ and $W_{i}$ is a subset of $V_{q_{i}}$ for some $q_{i}$ with $ar<q_{i}\leq k$ otherwise.
For simplicity, we write $z_{t}(n)=z_{t}^{(\lceil k/r\rceil)}(n)$.
The fact that $G$ is $\gamma$-close to $T$ gives the following observation.
  1. (D0)
    for any $i\in[r]$, there exists $B_{i}\subseteq U_{i}$ of size at most $2\sqrt{\gamma}n$ such that for any $v\in U_{i}\setminus B_{i}$ and $A\subseteq\bigcup_{j\in[r]\setminus\{i\}}U_{j}$ satisfying that none of the vertices of $A$ is in the same cluster as $v$ is, we have $\overline{d}(v,A)\leq\sqrt{\gamma}n$.
To see it, fix $i\in[r]$ and write $U^{*}:=\bigcup_{j\neq i}U_{j}$. Since $G$ is $\gamma$-close to $T$, we have
$e_{G}(Z_{i},U^{*})\geq|Z_{i}||U^{*}|-\gamma n^{2},\quad\text{and}\quad e_{G}\left(W_{i},U^{*}\setminus V_{q_{i}}\right)\geq|W_{i}||U^{*}\setminus V_{q_{i}}|-\gamma n^{2}.$
Let $B_{i}^{\prime}\subseteq Z_{i}$ be the set of vertices $v$ such that $\overline{d}(v,U^{*})>\sqrt{\gamma}n$, and $B_{i}^{\prime\prime}\subseteq W_{i}$ be the set of vertices $w$ such that $\overline{d}(w,U^{*}\setminus V_{q_{i}})>\sqrt{\gamma}n$. The displayed line above implies that $|B_{i}^{\prime}|\leq\sqrt{\gamma}n$ and $|B_{i}^{\prime\prime}|\leq\sqrt{\gamma}n$. Now (D0) holds by setting $B_{i}=B_{i}^{\prime}\cup B_{i}^{\prime\prime}$.
Minimum degree
For $i\in[k]$, let $N_{i}:=N_{T}(u_{i})$ for some $u_{i}\in V_{i}$. Note that this is well-defined as the vertices of $V_{i}$ share the same neighborhood in $T$. Using the maximality of $e(G)$, we derive that for every $u\in V_{i}$, $i\in[k]$
$d_{G}(u)\geq d_{T}(u)-2t\gamma n.$
Indeed, since $G$ is $\gamma$-close to $T$, that is, $|E(G)\triangle E(T)|\leq\gamma n^{2}$, for each $i\in[k]$, we can greedily pick distinct $u_{1},\dots,u_{t}\in V_{i}$, such that $|N_{G}(u_{j})\triangle N_{i}|\leq\gamma n^{2}/(n-j)\leq 2\gamma n$, for $j\in[t]$. Let $N_{i}^{\prime}:=\bigcap_{j\in[t]}N_{G}(u_{j})$ and note that $|N_{i}^{\prime}\triangle N_{i}|\leq 2t\gamma n$. In particular, $|N_{i}^{\prime}|\geq|N_{i}|-2t\gamma n$. Now for a contradiction suppose there is $u\in V_{i}$ such that $d_{G}(u)<d_{T}(u)-2t\gamma n=|N_{i}|-2t\gamma n$. Then we replace $N_{G}(u)$ by $N_{i}^{\prime}$, that is, we disconnect all the edges of $u$ in $G$ and connect $u$ to the vertices of $N_{i}^{\prime}$. Thus, we obtain a $k$-partite graph on the same vertex set as $G$ and has more edges than $G$. Therefore, by the maximality of $G$, this new graph contains a copy of $K_{r+1}(t)$, denoted by $K$. Clearly, $K$ must contain the vertex $u$, as $G$ is $K_{r+1}(t)$-free. Moreover, $K$ must miss at least one vertex from $u_{1},\dots,u_{t}$, say $u_{j}$, because the set $\{u,u_{1},\dots,u_{t}\}$ is independent in $G$ and $K$ has independence number $t$. However, as the neighborhood of $u$ $N_{i}^{\prime}$ is a subset of $N_{G}(u_{j})$, we can replace $u$ by $u_{j}$ and still get a copy of $K_{r+1}(t)$, which is in $G$, a contradiction.
Therefore, comparing with the degrees in $T$, we derive that for any vertex $u$,
$d_{G}(u)\geq\begin{cases}(k-a)n-|W_{i}|-2t\gamma n,\,\,\,\,\hfill\text{ if $u\in Z_{i}$ for $i\in[r]$,}\\ (k-1-a)n-2t\gamma n,\,\,\,\,\hfill\text{ if }u\in W.\end{cases}$ (4.1)
Atypical vertices
In this step we identify a set of atypical vertices, that is, those behave differently from the majority of the vertices. Let $W:=\bigcup_{i\in[r]}W_{i}=V_{ar+1}\cup\cdots\cup V_{k}$. We define $W^{\prime\prime}:=\{v\in W:d(v,Z_{j})\geq\varepsilon n,\text{ for all }j\in[r]\}$ and $W_{i}^{\prime}:=\{v\in W:d(v,Z_{i})<\varepsilon n\}$. Then we have $W=W^{\prime\prime}\cup W_{1}^{\prime}\cup\cdots\cup W_{r}^{\prime}$. Next, for $i\in[r]$, let $Z^{\prime\prime}:=\bigcup_{i\in[r]}Z_{i}^{\prime\prime}$, where
$Z_{i}^{\prime\prime}:=\{v\in Z_{i}:d(v,Z_{j})\geq\varepsilon n,\text{ for all }j\in[r]\setminus\{i\}\quad\text{and}\quad d(v,U_{i})\geq\varepsilon n\}.$
Furthermore, let $Z_{i}^{\prime}:=Z_{i}\setminus Z_{i}^{\prime\prime}$ and write $Z_{i}^{\prime}$ as $\bigcup_{j\in[r]}Z_{i}^{j}$, where $Z_{i}^{j}$, $j\neq i$, consists of the vertices $v\in Z_{i}$ such that $d(v,Z_{j})<\varepsilon n$, and $Z_{i}^{i}$ consists of the vertices $v$ such that $d(v,U_{i})<\varepsilon n$. The following are some useful properties of these sets.
Claim 4.1.
The following properties hold for all $i\in[r]$.
  1. $(P1)$
    $|W_{i}^{\prime}\setminus W_{i}|\leq 2\gamma n$ and $|W^{\prime\prime}|\leq C_{0}:=2t\varepsilon^{-rt}$.
  2. $(P2)$
    $W=W^{\prime\prime}\cup W_{1}^{\prime}\cup\cdots\cup W_{r}^{\prime}$ is a partition of $W$.
  3. $(P3)$
    $|Z_{i}^{\prime\prime}|\leq C_{0}$, $|Z_{i}^{j}|\leq\sqrt{\gamma}n$ for $j\neq i$, and $|Z_{i}^{i}|\geq(1-\sqrt{\gamma})an$.
  4. $(P4)$
    $\bigcup_{j\in[r]}Z_{i}^{j}$ is a partition of $Z_{i}^{\prime}$.
Proof.
Recall the definition of $W^{\prime\prime}$ and that $d(Z_{i},Z_{j})\geq 1-\gamma$ for distinct $i,j\in[r]$. Applying Proposition 3.4 to the graph $G[W^{\prime\prime}\cup Z]$ with vertex partition $(U_{1},\dots,U_{r})$, we obtain that $|W^{\prime\prime}|\leq C_{0}:=2t\varepsilon^{-rt}$. We next show that $|W_{i}^{\prime}\setminus W_{i}|\leq 2\gamma n$ for each $i\in[r]$. Indeed, because $G$ is $\gamma$-close to $T$, we have $e_{G}(Z_{i},W_{i}^{\prime}\setminus W_{i})\geq an|W_{i}^{\prime}\setminus W_{i}|-\gamma n^{2}$. On the other hand, by definition, $e_{G}(Z_{i},W_{i}^{\prime}\setminus W_{i})<|W_{i}^{\prime}\setminus W_{i}|\cdot\varepsilon n$. Thus, we get $|W_{i}^{\prime}\setminus W_{i}|<\gamma n/(a-\varepsilon)<2\gamma n$, verifying $(P1)$.
To see $(P2)$, suppose there is a vertex $v\in W_{i}^{\prime}\cap W_{j}^{\prime}$. By definition, $d(v)\leq(k-1)n-2(a-\varepsilon)n<(k-1-a)n-\sqrt{\gamma}n$, contradicting (4.1).
Next we show $(P3)$. Fix $i\in[r]$. Since $G$ is $\gamma$-close to $T$, we have $d(Z_{j},Z_{j^{\prime}})\geq 1-\gamma$ and $d(U_{i},Z_{j})\geq 1-\gamma$ for distinct $j,j^{\prime}\in[r]\setminus\{i\}$. Thus, we can apply Proposition 3.4 on $G[U_{i}\cup\bigcup_{j\neq i}Z_{j}]$ (with the obvious $r$-partition) and obtain $|Z_{i}^{\prime\prime}|\leq C_{0}$. Moreover, for $i\neq j$, from $d(Z_{i},Z_{j})\geq 1-\gamma$ we infer $|Z_{i}^{j}|\leq(\gamma/\varepsilon)n\leq\sqrt{\gamma}n$, as $\gamma\ll\varepsilon$. Therefore, we also get $|Z_{i}^{i}|\geq|Z_{i}|-|Z_{i}^{\prime\prime}|-\sum_{j\neq i}|Z_{i}^{j}|\geq an-C_{0}-(r-1)\gamma n/\varepsilon\geq(1-\sqrt{\gamma})an$.
Now we show $(P4)$. By definition, if $v\in Z_{i}^{i}$, then $d(v,U_{i})<\varepsilon n$; if $v\in Z_{i}^{j}$ for $j\neq i$, then $d(v,Z_{j})<\varepsilon n$. Thus, we have $Z_{i}^{\prime}\subseteq\bigcup_{j\in[r]}Z_{i}^{j}$ by definition. A vertex $v\in Z_{i}^{i}\cap Z_{i}^{j}$, $j\neq i$, satisfies that $d(v)<kn-(|U_{i}|-\varepsilon n)-(a-\varepsilon)n\leq(k-a)n-|W_{i}|-(1-2\varepsilon)n$, contradicting (4.1). A vertex $v\in Z_{i}^{j}\cap Z_{i}^{j^{\prime}}$ for distinct $j,j^{\prime}\in[r]\setminus\{i\}$ satisfies that $d(v)<(k-1)n-2(a-\varepsilon)n\leq(k-a-2)n+2\varepsilon n$, contradicting (4.1) as well. Thus, $\bigcup_{j\in[r]}Z_{i}^{j}$ is a partition of $Z_{i}^{\prime}$. ∎
For $i\in[r]$, our refined partition is defined by
$\tilde{U}_{i}:=\tilde{Z}_{i}\cup W_{i}^{\prime},\text{ where }\tilde{Z}_{i}:=\bigcup_{j\in[r]}Z_{j}^{i}.$ (4.2)
Then $V(G)=Z^{\prime\prime}\cup W^{\prime\prime}\cup\bigcup_{i\in[r]}\tilde{U}_{i}$. Note that for any $v\in\tilde{U}_{i}$, we have $d(v,Z_{i}^{i})\leq d(v,Z_{i})\leq\varepsilon n$, and thus $d(v,\tilde{Z}_{i})\leq\varepsilon n+(r-1)\sqrt{\gamma}n$ by $(P3)$.
For every $i\in[r]$, note that $(P1)$ implies that $|W_{i}\setminus W_{i}^{\prime}|\leq C_{0}+(r-1)2\gamma n\leq 2r\gamma n$, and similarly $(P3)$ implies that $|Z_{i}\setminus\tilde{Z}_{i}|\leq C_{0}+(r-1)\sqrt{\gamma}n\leq r\sqrt{\gamma}n$.
We now derive a more handy minimum degree condition. For convenience, define $\overline{d}(v,A)=|A|-d(v,A)$. For $v\in Z_{i}^{i}$, we have $\overline{d}(v,\tilde{U}_{i})\geq\overline{d}(v,U_{i})-|U_{i}\setminus\tilde{U}_{i}|$. Since $\overline{d}(v,U_{i})>an+|W_{i}|-\varepsilon n$ and $|U_{i}\setminus\tilde{U}_{i}|\leq|Z_{i}\setminus\tilde{Z}_{i}|+|W_{i}\setminus W^{\prime}_{i}|\leq\varepsilon n/2$, we have $\overline{d}(v,\tilde{U}_{i})\geq an+|W_{i}|-\varepsilon n-\varepsilon n/2$. By (4.1), $\overline{d}(v)\leq an+|W_{i}|+\sqrt{\gamma}n$. It follows that $\overline{d}(v,V\setminus\tilde{U}_{i})\leq 2\varepsilon n$. Now consider $v\in\tilde{U}_{i}\setminus Z^{i}_{i}$. The definition of $\tilde{U}_{i}$ implies that $d(v,Z_{i})<\varepsilon n$ and $\overline{d}(v,Z_{i})>an-\varepsilon n$. Assume $v\in V_{j}$. Then $V_{j}\cap Z_{i}=\emptyset$ and trivially $\overline{d}(v,V_{j})=n$. It follows that $\overline{d}(v,Z_{i}\cup V_{j})>(a+1)n-\varepsilon n$. Hence $\overline{d}(v,\tilde{Z}_{i}\cup V_{j})\geq\overline{d}(v,Z_{i}\cup V_{j})-|Z_{i}\setminus\tilde{Z}_{i}|>(a+1)n-\frac{3}{2}\varepsilon n$. On the other hand, either case of (4.1) implies that $\overline{d}(v)\leq(a+1)n+\sqrt{\gamma}n$. Consequently, $\overline{d}(v,V\setminus(\tilde{Z}_{i}\cup V_{j}))\leq 2\varepsilon n$. In summary, for $i\in[r]$ and $j\in[k]$,
  1. (Deg)
    If $v\in Z_{i}^{i}$, then $\overline{d}(v,V\setminus\tilde{U}_{i})\leq 2\varepsilon n$; if $v\in(\tilde{U}_{i}\setminus Z^{i}_{i})\cap V_{j}$, then $\overline{d}(v,V\setminus(\tilde{Z}_{i}\cup V_{j}))\leq 2\varepsilon n$.
Next we prove further properties on $Z_{i}^{j}$ and $\tilde{Z}_{j}$.
Claim 4.2.
If $Z_{i}^{j}\neq\emptyset$ for some $i\neq j$, then the following holds.
  1. $(Q1)$
    For $v\in Z_{i}^{j}$ and $A\subseteq V(G)\setminus(Z_{i}\cup Z_{j})$, we have $d(v,A)\geq|A|-\varepsilon n-\sqrt{\gamma}n$.
  2. $(Q2)$
    $|W_{i}|\geq(1-\varepsilon-\sqrt{\gamma})n$.
  3. $(Q3)$
    If $|\tilde{Z}_{j}\setminus Z_{j}|\geq t$, then $|W_{j}|\leq 2t\varepsilon n$.
Proof.
Note that $d(v,Z_{j})\leq\varepsilon n$ and $d(v,Z_{i})\leq(a-1)n$, that is, $v$ has at least $n+(an-\varepsilon n)=(a+1)n-\varepsilon n$ non-neighbors in $Z_{i}\cup Z_{j}$. On the other hand, (4.1) says that $v$ has at most $an+|W_{i}|+\sqrt{\gamma}n$ non-neighbors in $G$. Combining these two we get that $v$ has at most $|W_{i}|-n+\varepsilon n+\sqrt{\gamma}n\leq\varepsilon n+\sqrt{\gamma}n$ non-neighbors outside $Z_{i}\cup Z_{j}$, and thus $(Q1)$ holds. The fact that $|W_{i}|-n+\varepsilon n+\sqrt{\gamma}n\geq 0$ implies $(Q2)$.
For $(Q3)$, suppose to the contrary, $|\tilde{Z}_{j}\setminus Z_{j}|\geq t$ and $|W_{j}|>2t\varepsilon n$. By $(Q1)$ with $A=W_{j}$, arbitrary $t$ vertices in $\tilde{Z}_{j}\setminus Z_{j}$ have at least $|W_{j}|-t(\varepsilon+\sqrt{\gamma})n\geq t$ common neighbors in $W_{j}$. We thus obtain a copy of $K_{t,t}$ with one part in $\tilde{Z}_{j}\setminus Z_{j}$ and the other part in $W_{j}$ – denote its vertex set by $B$. For any $i^{\prime}\in[r]\setminus\{j\}$ such that $B\cap Z_{i^{\prime}}^{j}\neq\emptyset$, we have $|W_{i^{\prime}}|\geq(1-\varepsilon-\sqrt{\gamma})n$ by $(Q2)$. Since $|W_{j}|>2t\varepsilon n$, $W_{i^{\prime}}$ and $W_{j}$ do not belong to the same cluster, and thus no vertex of $B$ is in the same cluster that contains $W_{i^{\prime}}$, which implies that the vertices of $B$ have at least $|W_{i^{\prime}}|-2t(2\varepsilon n)\geq n/2$ common neighbors in $W_{i^{\prime}}$ by (Deg). For any $i^{\prime\prime}\in[r]\setminus\{j\}$ such that $B\cap Z_{i^{\prime\prime}}^{j}=\emptyset$ (and thus $B\cap Z_{i^{\prime\prime}}=\emptyset$), by (Deg) we have that the vertices of $B$ have at least $n/2$ common neighbors in $Z_{i^{\prime\prime}}$. Because $G$ is $\gamma$-close to $T$, these common neighborhoods, each of size at least $n/2$, have densities close to one between each pair, and thus contain a copy of $K_{r-1}{(t)}$. Together with $B$, they form a copy of $K_{r+1}(t)$ in $G$, a contradiction. ∎
In particular, when $b=0$ (and thus $W_{i}=\emptyset$ for all $i$), $(Q2)$ implies that $Z_{i}^{j}=\emptyset$ whenever $i\neq j$. Consequently,
$\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}\quad\text{for all $i\in[r]$ when $b=0$}.$ (4.3)
Let $L\subseteq[r]$ be the set of indices $i$ such that $|W_{i}|\geq(1-\varepsilon-\sqrt{\gamma})n$. $(Q2)$ and $(Q3)$ imply that
  • for $i\in[r]\setminus L$, we have $Z_{i}^{j}=\emptyset$ for $j\neq i$.
  • for $i\in L$, $|\tilde{Z}_{i}\setminus Z_{i}|\leq t-1$ and thus $|\tilde{Z}_{i}|\leq an+t-1$.
First Estimate on $e(G)$
Let $G^{\prime}=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)$. Since $G^{\prime}$ is $r$-partite, it is $K_{r+1}$-free. As $G^{\prime}$ is a subgraph of $G\in\mathcal{G}_{k}(n)$, we have $e(G^{\prime})\leq t_{r}(k)n^{2}$ (but this is not good enough when $b>0$). Below we give an upper bound for $e(G^{\prime})$, which will be used throughout the proof. Recall that $T=K(V_{1},\dots,V_{k})\cap K(U_{1},\dots,U_{r})$ has precisely $t_{r}(k)n^{2}$ edges.
Claim 4.3.
We have $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(\beta_{i}-\alpha_{i})$, where
$\beta_{i}:=$ $\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|\right)\text{and }$
$\alpha_{i}:=$ $|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|+e_{T}(W_{i}^{\prime})+e_{T}(\tilde{Z}_{i}\setminus Z_{i}).$
Proof.
We first obtain $G^{(0)}:=K(Z_{1}\cup W_{1}^{\prime},\dots,Z_{r}\cup W_{r}^{\prime})\cap K(V_{1},\dots,V_{k})$ from $T$. During this process, we lose the edges of $T$ between $W_{i}$ and $W_{j}$, $j\neq i$, if both ends of the edges are placed in $W^{\prime}_{i}$. Thus
$e(G^{(0)})=t_{r}(k)n^{2}-\sum_{i\in[r]}e_{T}(W_{i}^{\prime}).$ (4.4)
We imagine a dynamic process of obtaining $G^{\prime}$ from $G^{(0)}$ by recursively moving vertices. To estimate $e(G^{\prime})$, we track the changes of the edges with respect to complete $r$-partite graphs (but also respecting the $k$-partition of $G$). More precisely, for $l>0$, let
$G^{(l)}:=K(Z_{1}^{(l)}\cup W_{1}^{\prime},\dots,Z_{r}^{(l)}\cup W_{r}^{\prime})\cap K(V_{1},\dots,V_{k})$
such that the $r$-partition of $G^{(l)}$ can be obtained by moving exactly one vertex from the partition of $G^{(l-1)}$. The process terminates after $m:=\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}|$ steps and thus $G^{\prime}$ is a subgraph of $G^{(m)}$. Furthermore, throughout the process, we only move vertices from the color classes in $L$ to other color classes. Therefore, we can give a linear ordering to the members of $L$, and for $i\in L$ we move vertices from $Z_{i}$ only after we have moved the vertices in color classes $j$ prior to $i$ (denoted by $j<_{L}i$). Now, in the $l$-th step, suppose we move $v$ from $Z_{j}^{(l-1)}$ to $Z_{i}^{(l-1)}$, namely, $v\in Z_{j}^{i}$, then the change is
$e(G^{(l)})-e(G^{(l-1)})=|Z_{j}^{(l-1)}\setminus V_{p}|+|W_{j}^{\prime}|-|\tilde{Z}_{i}^{(l-1)}|-|W_{i}^{\prime}|,$
where $V_{p}\ni v$ and $\tilde{Z}_{i}^{(l-1)}=Z_{i}^{(l-1)}\setminus V_{p}$.
Note that we have $|Z_{j}^{(l-1)}\setminus V_{p}|\leq(a-1)n+|\tilde{Z}_{j}\setminus Z_{j}|$. Moreover for any $j^{\prime}<_{L}j$, we have $Z_{j^{\prime}}^{i}\subseteq Z_{i}^{(l-1)}$. Therefore, we have $|\tilde{Z}_{i}^{(l-1)}|\geq an-|Z_{i}\setminus\tilde{Z}_{i}|+\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|$. Putting all these together, we get
$e(G^{(l)})-e(G^{(l-1)})\leq|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|-|W_{i}^{\prime}|.$
Recalling that we moved $v$ from $Z_{j}^{(l-1)}$ to $Z_{i}^{(l-1)}$ at the $l$-th step, we obtain
$e(G^{\prime})-e(G^{(0)})\leq\sum_{l=1}^{m}\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|-|W_{i}^{\prime}|\right),$
where $i,j$ depends on $l$. Since $m=\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}|$, we have
$\sum_{l=1}^{m}(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-|W_{i}^{\prime}|)$
$=$ $\sum_{i\in[r]}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|-|W_{i}^{\prime}|)$
$=$ $\sum_{i\in[r]}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|)-\sum_{i\in[r]}|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|.$
Moreover, it is not hard to see that
$\sum_{l=1}^{m}\sum_{j^{\prime}<_{L}j}|Z_{j^{\prime}}^{i}|=\sum_{i\in[r]}\sum_{\{j_{1},j_{2}\}\in\binom{L\setminus\{i\}}{2}}|Z_{j_{1}}^{i}||Z_{j_{2}}^{i}|=\sum_{i\in[r]}e_{T}(\tilde{Z}_{i}\setminus Z_{i}).$
Now the claim follows by combining these estimates with (4.4). ∎
What remains is to estimate the number of edges in each $\tilde{U}_{i}$. For $i\in[r]$, we have $e(G[\tilde{U}_{i}])=e(Z_{i}^{i},G[\tilde{U}_{i}])+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})$. To bound $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}((\tilde{Z}_{i}\setminus Z_{i})\cup W_{i}^{\prime})$, we note that $e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})\leq|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|$ and $e_{G}(W_{i}^{\prime})\leq e_{T}(W_{i}^{\prime})$. However, we may not have $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$ because each $Z_{j}^{i}$ is an independent set in $T$, but may not be independent in $G$ when $a\geq 2$. Thus, $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+\sum_{j\neq i}e_{G}(Z_{j}^{i})$. Putting these together, for each $i\in[r]$, we have
$e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})+e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq\alpha_{i}+\sum_{j\neq i}e_{G}(Z_{j}^{i}).$ (4.5)
Let $f_{i}:=e(Z_{i}^{i},G[\tilde{U}_{i}])$. By Claim 4.3, (4.5) and $e(G)=e(G^{\prime})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})$, we derive that
$e(G)$ $\leq t_{r}(k)n^{2}+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}\left(f_{i}+\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})\right)$ (4.6)
$\leq t_{r}(k)n^{2}+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}\left(f_{i}+\beta_{i}+\sum_{j\neq i}e_{G}(Z_{j}^{i})\right)$ (4.7)
We now focus on the structure of each $\tilde{U}_{i}$. We first show that $G[\tilde{U}_{i}]$ is “almost” $K_{t,t}$-free.
Claim 4.4.
The following holds for all $i\in[r]$.
  1. $(K1)$
    Both $G[\tilde{Z}_{i}]$ and $G[Z_{i}^{i}\cup W_{i}^{\prime}]$ are $K_{t,t}$-free.
  2. $(K2)$
    If $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$, then $|W_{i}^{\prime}\setminus V_{q_{i}}|\leq t-1$.
  3. $(K3)$
    If $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$, then $G[\tilde{Z}_{i}\cup(W_{i}^{\prime}\cap V_{q_{i}})]$ is $K_{t,t}$-free.
Proof.
For $(K1)$, suppose there is a copy of $K_{t,t}$ in $\tilde{U}_{i}$, with vertex set denoted by $B$, contained in $\tilde{Z}_{i}$ or in $Z_{i}^{i}\cup W_{i}^{\prime}$. Let $N_{B}$ be the set of common neighbors of these $2t$ vertices of $B$. First assume that $B\subseteq\tilde{Z}_{i}$. Then for any $j\in L\setminus\{i\}$, by (Deg) we have $|N_{B}\cap W_{j}^{\prime}|\geq|W_{j}^{\prime}|-4t\varepsilon n$, and thus by $(P1)$ $|N_{B}\cap W_{j}\cap W_{j}^{\prime}|\geq|W_{j}^{\prime}|-4t\varepsilon n-2\gamma n\geq n/2$. For any $j\notin L\cup\{i\}$, because $B\cap Z_{j}=\emptyset$ by $(Q2)$, we have $|N_{B}\cap Z_{j}|\geq an-4t\varepsilon n\geq n/2$ by (Deg). Note that every set in $\{N_{B}\cap Z_{j}:j\notin L\}\cup\{N_{B}\cap W_{j}\cap W_{j}^{\prime}:j\in L\}$ has size at least $n/2$ and every pair of them has density at least $1-4\gamma$. Therefore we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, which gives rise to a copy of $K_{r+1}(t)$ together with $B$, a contradiction.
Second we assume that $B\subseteq Z_{i}^{i}\cup W_{i}^{\prime}$. In this case we note that for any $j\neq i$, we have $B\cap Z_{j}=\emptyset$ and thus by (Deg), we have $|N_{B}\cap Z_{j}^{j}|\geq(1-\sqrt{\gamma})an-4t\varepsilon n\geq n/2$. Then as these sets have high pairwise densities, as in the previous case, we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, yielding a copy of $K_{r+1}(t)$ together with $B$, a contradiction. Now $(K1)$ is proved.
Now we turn to $(K2)$, and suppose $|W_{i}^{\prime}|>2t\varepsilon n+2\gamma n$ and thus $|W_{i}\cap W_{i}^{\prime}|>2t\varepsilon n$ by $(P1)$. First, if $W_{i}^{\prime}$ contains at least $t$ vertices which are not from $V_{q_{i}}$ (the cluster containing $W_{i}$), then by (Deg), each of these vertices have at most $2\varepsilon n$ non-neighbors in $W_{i}\cap W_{i}^{\prime}$, and thus we can find a copy of $K_{t,t}$ in $W_{i}^{\prime}$, contradicting $(K1)$. So we have $|W_{i}^{\prime}\setminus V_{q_{i}}|\leq t-1$.
For $(K3)$, suppose there is a copy of $K_{t,t}$ as stated in the claim, whose vertex set is denoted by $B$. As in the previous paragraph, we have $|W_{i}|>2t\varepsilon n$ by $(P1)$. Now observe crucially that if $B\cap Z_{j}^{i}\neq\emptyset$, then by $(Q2)$ $|W_{i}|+|W_{j}|>n$, and thus, $W_{i}$ and $W_{j}$ are not from the same cluster. So by (Deg), for any $j\in[r-1]\setminus\{i\}$, if $B\cap Z_{j}^{i}=\emptyset$, then the vertices of $B$ have large common neighborhoods in $Z_{j}^{j}$; if $B\cap Z_{j}^{i}\neq\emptyset$, then the vertices of $B$ have large common neighborhoods in $W_{j}\cap W_{j}^{\prime}$ (note that $|W_{j}|\geq(1-\varepsilon-\sqrt{\gamma})n$ by $(Q2)$). Since each of these common neighborhoods have size at least $n/2$ and each pair of them has high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these sets, yielding a copy of $K_{r+1}(t)$ together with $B$, a contradiction. ∎
We now derive a lower bound for $\sum f_{i}$ from Claims 4.3 and 4.4. For $i\in[r]$, we have $\beta_{i}\leq\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|)$ as $|W_{j}^{\prime}|-n\leq|W_{j}^{\prime}\setminus V_{q_{j}}|$. Fix $j\in L\setminus\{i\}$. Note that $|W_{j}|\geq(1-2\varepsilon)n$. We have $|\tilde{Z}_{j}\setminus Z_{j}|\leq t-1$ by $(Q3)$, and $|W_{j}^{\prime}\setminus V_{q_{j}}|\leq t-1$ by $(K2)$. If $|W_{i}|>n/2$, then $|Z_{j}^{i}|\leq t-1$ by $(Q3)$. Furthermore, since $|Z_{i}\setminus\tilde{Z}_{i}|\leq(r-1)\sqrt{\gamma}n+C_{0}$ by $(P3)$, it follows that
$|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|\right)\leq(t-1)(t-1+t-1+(r-1)\sqrt{\gamma}n+C_{0})\leq(t-1)r\sqrt{\gamma}n.$
Otherwise $|W_{i}|\leq n/2$, and by $(Q2)$, we have $Z_{i}^{i^{\prime}}=\emptyset$ for any $i^{\prime}\neq i$. This implies $|Z_{i}\setminus\tilde{Z}_{i}|=|Z_{i}^{\prime\prime}|\leq C_{0}$. Using $|Z_{j}^{i}|\leq\sqrt{\gamma}n$, $(Q3)$, and $(K2)$, we derive that
$|Z_{j}^{i}|\left(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|\right)\leq\sqrt{\gamma}n(2(t-1)+C_{0})\leq 2C_{0}\sqrt{\gamma}n.$
Summarizing these two cases for all $j\in L\setminus\{i\}$, we obtain that $\beta_{i}\leq(r-1)2C_{0}\sqrt{\gamma}n$, and consequently,
$\sum_{i\in[r]}\beta_{i}\leq 2(r-1)rC_{0}\sqrt{\gamma}n.$ (4.8)
On the other hand, for all $i\neq j$, the graph $G[Z_{j}^{i}]$ is $K_{t,t}$-free by $(K1)$ and thus, by $(P3)$, $\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})\leq r(r-1)C\left(\sqrt{\gamma}n\right)^{2-1/t}$. Applying this with (4.7), (4.8), and the fact that $e(Z^{\prime\prime}\cup W^{\prime\prime},G)\leq(r+1)C_{0}kn$, we obtain that
$e(G)$ $\leq t_{r}(k)n^{2}+(r+1)C_{0}kn+\sum_{i\in[r]}f_{i}+2(r-1)rC_{0}\sqrt{\gamma}n+\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})$
$\leq t_{r}(k)n^{2}+\sum_{i\in[r]}f_{i}+r^{2}C\sqrt{\gamma}n^{2-1/t},$
as $\gamma\ll 1$. Using the assumption $e(G)\geq g(n,r,k,t)\geq t_{r}(k)n^{2}+z_{t}(n)$, we infer that
$\sum_{i\in[r]}f_{i}\geq z_{t}(n)-r^{2}C\sqrt{\gamma}n^{2-1/t}\geq\frac{c}{2}n^{2-1/t}$ (4.9)
by using ($Z$), $z_{t}(n)\geq z_{t}^{(2)}(n)\geq cn^{2-1/t}$, and $\gamma\ll 1$.
We next study the existence of $K_{1,t}$ in each color class. To do so, we consider a copy of $K_{3}(t)$ in $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ for some $i\neq j$.
Claim 4.5.
For any $i\neq j$, if $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ contains a copy $K$ of $K_{3}{(t)}$, then there exists $l\notin\{i,j\}$ such that $V(K)$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$.
Proof.
We may assume that $r>2$ as otherwise the claim is trivial. Suppose to the contrary that there is a copy $K$ of $K_{3}{(t)}$ in, say, $\tilde{U}_{1}$ and $\tilde{U}_{2}$, such that for every $l\in[3,r]$, there is a cluster in $U_{l}$ which does not intersect $B:=V(K)$. Let $V_{i_{l}}$ be a cluster in $Z_{l}$ such that $B\cap V_{i_{l}}=\emptyset$, and if there is no such cluster in $Z_{l}$, then we choose $V_{i_{l}}=V_{q_{l}}$. Note that in the former case, we have $|\tilde{U}_{l}\cap V_{i_{l}}|=|Z_{l}^{l}\cap V_{i_{l}}|\geq(1-\sqrt{\gamma}a)n$. In the latter case, we have $Z_{l}^{1}\neq\emptyset$ or $Z_{l}^{2}\neq\emptyset$, which implies that $|W_{l}|\geq(1-2\varepsilon)n$ by $(Q2)$, and thus $|\tilde{U}_{l}\cap V_{i_{l}}|=|W_{l}^{\prime}\cap V_{i_{l}}|\geq(1-3\varepsilon)n$. Now, by (Deg), every vertex in $B$ has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{l}\cap V_{i_{l}}$ for each $l\in[3,r]$. Since for every $l$ we have $|\tilde{U}_{l}\cap V_{i_{l}}|\geq 0.9n$, one can find large common neighborhoods (e.g. of size $n/2$) of all vertices of $B$ in each $\tilde{U}_{l}\cap V_{i_{l}}$, and then find a copy of $K_{r-2}{(t)}$ in these sets. Altogether we obtain a copy of $K_{r+1}(t)$, a contradiction.
Therefore, for such a copy $K$ of $K_{3}(t)$, there exists $l\notin\{i,j\}$ such that $K$ must intersect all clusters of $U_{l}$. Since $V(K)\cap Z_{l}\neq\emptyset$, we have $Z_{l}^{i}\neq\emptyset$ or $Z_{l}^{j}\neq\emptyset$. Then by $(Q2)$, $|W_{l}|\geq(1-2\varepsilon)n$ and in particular, $V_{q_{l}}\neq\emptyset$. Therefore $V(K)\cap V_{q_{l}}\neq\emptyset$. ∎
Claim 4.6.
For all but exactly one $j\in[r]$, we have $d(v,Z_{j}^{j})\leq t-1$ for all $v\in\tilde{U}_{j}$.
Proof.
First assume that there exists $j\in[r]$ such that $G[\tilde{U}_{j}]$ contains a copy of $K_{1,t}$, with vertex set denoted by $\{v,u_{1},\dots,u_{t}\}$, $v\in\tilde{U}_{j}$ and $u_{1},\dots,u_{t}\in Z_{j}^{j}$. Fix $i\in[r]\setminus\{j\}$ and let $N^{\prime}$ be the set of common neighbors of $u_{1},\dots,u_{t}$ in $\tilde{U}_{i}\cap U_{i}$. Suppose $v\in V_{p}$ and let $N$ be the set of common neighbors of these $t+1$ vertices in $\tilde{U}_{i}\cap U_{i}$. In particular, $N\subseteq N^{\prime}$ and $N$ is almost equal to the union of $a$ or $a+1$ clusters in $\tilde{U}_{i}$. Suppose there is a copy of $K_{t-1,t}$ with parts $S_{1}$ of size $t-1$ and $S_{2}$ of size $t$ such that $S_{1}\subseteq N^{\prime}$ and $S_{2}\subseteq N$. Then by Claim 4.5, there exists $l\in[r]\setminus\{i,j\}$ such that $B\cap Z_{l}\neq\emptyset$ and $B\cap V_{q_{l}}\neq\emptyset$, where $B$ denotes the vertex set of the copy of $K_{3}{(t)}$. This is impossible since $v$ is the only possible vertex in $B\cap(Z_{l}\cup V_{q_{l}})$ and can not satisfy both. Therefore, letting $N^{*}=N\cup(N^{\prime}\cap V_{p})$, we infer that $e_{G}(N^{*})=e_{G}(N)+e_{G}(N,N^{\prime}\setminus N)=O(n^{2-1/(t-1)})$.
By $(P1)$$(P3)$ and (Deg), we have $|\tilde{U}_{i}\setminus N^{*}|\leq 3(t+1)\varepsilon n$. Let $E^{i}$ be the set of the edges incident to $\tilde{U}_{i}\setminus N^{*}$ and counted in $f_{i}$. We split it to $E^{i}\cap E_{G}(Z_{i}^{i})$ and $E^{i}\cap E_{G}(\tilde{U}_{i}\setminus Z_{i}^{i},Z_{i}^{i})$. Note that by $(K1)$, each of the terms can be split further into at most $k$ $K_{t,t}$-free bipartite graphs, each with one part of size at most $3(t+1)\varepsilon n$ and the other part of size at most $(1+(r-2)\sqrt{\gamma})an$. Therefore, we obtain that
$f_{i}=O(\varepsilon n^{2-1/t})+O(n^{2-1/(t-1)})=O(\varepsilon n^{2-1/t}).$ (4.10)
Now assume there exist distinct $j_{1},j_{2}\in[r]$ such that each $G[\tilde{U}_{j_{i}}]$ contains a copy of $K_{1,t}$ whose part of size $t$ is in $Z_{j_{i}}^{j_{i}}$. The arguments above imply that (4.10) holds for all $i\in[r]$, and consequently, $\sum_{i\in[r]}f_{i}=O(\varepsilon n^{2-1/t})$, contradicting (4.9).
On the other hand, if $d(v,Z_{j}^{j})\leq t-1$ for all $j\in[r]$ and all $v\in\tilde{U}_{j}$, then $\sum_{j\in[r]}f_{j}\leq(t-1)kn$, again contradicting (4.9). ∎
By Claim 4.6, without loss of generality, we assume that,
for $\text{for }i\geq 2,\quad d(v,Z_{i}^{i})\leq t-1\ \text{for all}\ v\in\tilde{U}_{i},\quad\text{and thus},\quad f_{i}\leq\begin{cases}(t-1)|\tilde{U}_{i}|\hfill\text{ if }a\geq 2,\\ (t-1)|\tilde{U}_{i}\setminus Z_{i}^{i}|\hfill\text{ if }a=1.\end{cases}$ (4.11)
If $b=0$, then $\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}$ for all $i$ by (4.3). In this case $\tilde{U}_{1}$ is $K_{t,t}$-free by $(K1)$ and $\tilde{U}_{i}$ is $K_{1,t}$-free for all $i\geq 2$ by (4.11). Since $G$ is $\gamma$-close to $K_{r}(an)$, $G[U_{i}\setminus Z^{\prime\prime},U_{j}\setminus Z^{\prime\prime}]$ is almost complete for all $i\neq j$. This completes the proof of Theorem 4 with $Z:=Z^{\prime\prime}$. ∎
By (4.9) and (4.11), we get
$f_{1}\geq z_{t}(n)-\varepsilon n^{2-1/t}.$ (4.12)
In particular, we claim that
$|W_{1}|>3t\varepsilon n\quad\text{ if }b>0$ (4.13)
(which we will refine a moment later). Indeed, the edges counted in $f_{1}$ can be covered by $G[Z_{1}^{1}]$, $G[Z_{1}^{1},W_{1}\cap W_{1}^{\prime}]$, and at most $k$ $K_{t,t}$-free bipartite graphs, each with a part of size at most $\sqrt{\gamma}n$ and a part of size at most $an$. If $|W_{1}|\leq 3t\varepsilon n$, then $e_{G}(Z_{1}^{1},W_{1}\cap W_{1}^{\prime})=O(\varepsilon n^{2-1/t})$. Together with $e_{G}(Z_{1}^{1})\leq z_{t}^{(a)}(n)$, we have
$f_{1}\leq z_{t}^{(a)}(n)+O(\varepsilon n^{2-1/t})<z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}$
by ($E1$), contradicting (4.12).
Now we can give a much cleaner structure, shown in a series of claims below. A key step is to show that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$. From now on we only consider $k\leq 2r$.
Claim 4.7.
Suppose $v_{0}\in V(G)$ and $i\in[r]$ satisfy that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$. Then $v_{0}$ has less than $\varepsilon n$ neighbors in $U_{i}$. In particular, we have $Z^{\prime\prime}=\emptyset$ and $W^{\prime\prime}=\emptyset$.
Proof.
The second part of the claim follows immediately from the definitions of $Z^{\prime\prime}$ and $W^{\prime\prime}$.
Suppose to the contrary, that there exist $v_{0}\in V(G)$ and $i\in[r]$ such that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$ and at least $\varepsilon n$ neighbors in $U_{i}$. Since $|Z_{j}^{j}|\geq(1-\sqrt{\gamma})an$ for all $j\in[r]$, there exist sets $N_{1},\dots,N_{r-1}$ each of size $\varepsilon n-\sqrt{\gamma}n$ such that $N_{j}\subseteq Z_{j}^{j}\cap N(v_{0})$ for $j\neq i$ and $N_{i}\subseteq(Z_{i}^{i}\cup W_{i})\cap N(v_{0})$. Recall that $W_{1}^{\prime}=W_{1}^{\prime}\cap V_{q_{1}}$. By averaging, there exists $N_{1}^{\prime}\subseteq N_{1}$ with $|N_{1}^{\prime}|\geq(\varepsilon n-\sqrt{\gamma}n-2r\gamma n)/2\geq\varepsilon n/3$ such that all vertices of $N_{1}^{\prime}$ are in $Z_{1}^{1}\cup W_{1}^{\prime}$ and from the same cluster, that is,
$N_{1}^{\prime}\subseteq Q,\text{ where }Q\in\{Z_{1}^{1},W_{1}^{\prime}\}.$
Note that $N_{1}^{\prime}\subseteq W_{1}^{\prime}$ is possible only if $i=1$ and $a=1$. If $i\neq 1$, then let $N_{i}^{\prime}:=N_{i}\setminus((W_{i}\setminus W_{i}^{\prime})\cup V_{q_{1}})$ and for every $j\in[r]\setminus\{1,i\}$, let $N_{j}^{\prime}:=N_{j}$. By $(P1)$, $|W_{i}\setminus W_{i}^{\prime}|\leq 2r\gamma n$, and by (4.16), $|W_{i}\cap V_{q_{1}}|\leq\gamma n$. Thus, we have $|N_{j}^{\prime}|\geq\varepsilon n/3$ for all $j\in[r]$. Because the sets $N_{j}^{\prime}$ are small, we can not apply the degree conditions (Deg) to them and instead, we use (D0).
Recall that $B_{1}$ is given by (D0). Next we show that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}$. Suppose instead, there is such a copy of $K_{t-1,t}$, with parts denoted by $A$ and $B$, such that $|A|=t$, $A\subseteq N_{1}^{\prime}\setminus B_{1}$ and $B\subseteq\tilde{U}_{1}\setminus B_{1}$. Recall that $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ and for each $j\in[r]\setminus\{1,i\}$, $N_{j}^{\prime}\subseteq Z_{j}^{j}$. Observe that for every $v\in\tilde{U}_{1}\setminus B_{1}$, we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$. Indeed, if $j\neq i$, then $N_{j}^{\prime}\subseteq Z_{j}^{j}$ and we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$ by (D0); otherwise note that $N_{i}^{\prime}\subseteq Z_{i}^{i}\cup(W_{i}^{\prime}\cap W_{i})$, and by (D0) and $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ we have $d(v,N_{i}^{\prime})\geq|N_{i}^{\prime}|-\sqrt{\gamma}n$. Therefore, we obtain that the vertices in $A\cup B$ have at least $|N_{j}^{\prime}|-(2t-1)\sqrt{\gamma}n\geq(1-\gamma^{1/3})|N_{j}^{\prime}|$ common neighbors in each $N_{j}^{\prime}$, $j\in[2,r]$. Because each pair $N_{j}^{\prime},N_{j^{\prime}}^{\prime}$ has a high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these common neighborhoods, which together with $A\cup B\cup\{v_{0}\}$ form a copy of $K_{r+1}(t)$, a contradiction.
Now given that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we give a refined estimate on $f_{1}$. Indeed, since $G[N_{1}^{\prime}\setminus B_{1},Z_{1}^{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we get $e_{G}(N_{1}^{\prime}\setminus B_{1},Z_{1}^{1}\setminus B_{1})=O(n^{2-1/(t-1)})$. Similarly $e_{G}(N_{1}^{\prime}\setminus B_{1},W_{1}^{\prime}\setminus B_{1})=O(n^{2-1/(t-1)})$. Suppose $N_{1}^{\prime}\subseteq V_{q}$ for some $q\in\{1,q_{1}\}$, then we have
$E(G[\tilde{U}_{1}])=E(G[\tilde{U}_{1}\setminus(N_{1}^{\prime}\setminus B_{1})])\cup E(G[N_{1}^{\prime}\setminus B_{1},\tilde{U}_{1}\setminus(B_{1}\cup V_{q})])\cup E(G[N_{1}^{\prime}\setminus B_{1},B_{1}\cap\tilde{U}_{1})]).$
Recall that $|N_{1}^{\prime}|\geq\varepsilon n/3$ and $|B_{1}|\leq 2\sqrt{\gamma}n$. Therefore, (regardless of $a=1$ or $(a,b)=(2,0)$) we can bound $f_{1}\leq|E(G[\tilde{U}_{1}])|$ by
$f_{1}\leq z_{t}\left((1-\tfrac{\varepsilon}{3})n,n\right)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})<z_{t}(n)-3rC_{0}kn,$
where we used ($E2$) and $\gamma\ll\varepsilon$. This contradicts (4.12). ∎
When $a=2$ and $b=0$ (i.e., $k=2r$), since $Z^{\prime\prime}_{i}=\emptyset$ and $W_{i}=\emptyset$ for all $i\in[r]$, by (4.3), we get $\tilde{U}_{i}=Z_{i}$ for all $i\in[r]$. Therefore $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)(t-1)n$ by $(K1)$ and (4.11), proving Theorem 3 for $k=2r$.
For the remaining of the proof, we only need to consider $a=1$ (and thus $b>0$). Moreover, now for $i,j\in[r]$ each $Z_{i}^{j}\subseteq Z_{i}$ is an independent set and thus $e_{G}(Z_{i}^{j})=0$. So we can first update our bounds on $e(G)$ and $f_{1}$. Recall the bounds (4.7), (4.8) and (4.11) and we have
$e(G)$ $\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i})$
$\leq t_{r}(k)n^{2}+f_{1}+(r-1)(t-1)(1+\sqrt{\gamma})n+2(r-1)rC_{0}\sqrt{\gamma}n,$
yielding
$f_{1}\geq z_{t}(n,n)-C_{0}n$ (4.14)
Claim 4.8.
Suppose $b>0$. Then $\tilde{U}_{1}=Z_{1}^{1}\cup W_{1}^{\prime}$ and $W_{1}^{\prime}\subseteq V_{q_{1}}$.
Proof.
Suppose to the contrary, there is a vertex $v$ in $\tilde{U}_{1}\setminus(Z_{1}^{1}\cup W_{1}^{\prime})$ or $W_{1}^{\prime}\setminus V_{q_{1}}$, namely, $v\in Z_{i}^{1}$ for some $2\leq i\leq r$ or $v\in W_{1}^{\prime}\setminus V_{q_{1}}$. Suppose $v\in V_{l}$. Then $l\neq q_{1}$. Moreover, if $i$ is defined, then $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus(\tilde{Z}_{i}\cup V_{l})$; otherwise, $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus V_{l}$. By (Deg), we have $\overline{d}(v,W_{1}^{\prime}\cap V_{q_{1}})\leq 2\varepsilon n$. Let $N:=W_{1}^{\prime}\cap V_{q_{1}}\cap N(v)$. We have $|(W_{1}^{\prime}\cap V_{q_{1}})\setminus N|\leq 2\varepsilon N$. Since $|W_{1}^{\prime}\setminus V_{q_{1}}|\leq|W_{1}^{\prime}\setminus W_{1}|\leq 2\gamma n$, it follows that $|W^{\prime}_{1}\setminus N|\leq 2\varepsilon n+2\gamma n\leq 3\varepsilon n$.
Recall (4.13), $|W_{1}|>3t\varepsilon n$. By $(K3)$ (if $v\in\tilde{Z}_{1}\setminus Z_{1}$) or $(K1)$ (if $v\in W_{1}^{\prime}\setminus V_{q_{1}}$), we know that $G[Z_{1}^{1},N]$ contains no $K_{t-1,t}$ with the part of size $t$ in $N$. This implies that $e_{G}(Z_{1}^{1},N)=O(n^{2-1/(t-1)})$. Furthermore, by $(P3)$ and $(K1)$, $G[\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1}]$ is a $K_{t,t}$-free bipartite graph with one part of size at most $(r-1)\sqrt{\gamma}n$ and the other part of size at most $n$. Thus, $e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})\leq C(r-1)\sqrt{\gamma}n^{2-1/t}$. By the similar arguments, we have $e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})\leq C(3\varepsilon n)n^{1-1/t}$.
Putting these bounds together (and note that $Z_{1}^{1}$ is an independent set, we get
$f_{1}$ $=e_{G}(Z_{1}^{1},N)+e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})+e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})$
$=O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})+O(\varepsilon n^{2-1/t}).$
By ($E1$), this contradicts (4.14). ∎
Claim 4.8 shows that $\tilde{U}_{1}$ has no atypical vertices and is thus $K_{t,t}$-free by $(K1)$. Furthermore, since $\tilde{U}_{1}=Z_{1}^{1}\cup W_{1}^{\prime}$ and $W_{1}^{\prime}\subseteq V_{q_{1}}$, it follows that
$\alpha_{1}=\beta_{1}=0,\quad\text{and}\quad e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(|Z_{1}^{1}|,|W^{\prime}_{1}|).$ (4.15)
Therefore, if $|W_{1}^{\prime}|\leq(1-\gamma)n$, then we have $f_{1}\leq z_{t}(n,|W_{1}^{\prime}|)\leq z_{t}(n,n)-\delta n^{2-1/t}$ for some $\delta>0$ by ($E2$). This contradicts (4.14). So we obtain
$\text{if }a=1,\text{ then }|W_{1}^{\prime}|\geq(1-\gamma)n\quad(\text{and thus}\quad 1\in L).$ (4.16)
Next we study $G[\tilde{U}_{i}]$ for $i\geq 2$. A key observation is that copies of $K_{1,t}$ in $G[\tilde{U}_{i}]$ together with copies of $K_{t-1,t}$ in $\tilde{U}_{1}$ may form copies of $K_{3}{(t)}$, which are restricted by Claim 4.5.
Claim 4.9.
Suppose $i\in[2,r]$.
  1. (1)
    If there is a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$, then there exists $l\in[r]\setminus\{i\}$ such that the vertex set of $K_{1,t}$ intersects both $V_{q_{l}}$ and $Z_{l}$.
  2. (2)
    Both $\tilde{Z}_{i}\setminus Z_{1}$ and $Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ are $K_{1,t}$-free.
Proof.
For Part (1), let $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$. Since $B\cap(Z_{1}\cup V_{q_{1}})=\emptyset$ and $\tilde{U}_{1}\subseteq Z_{1}\cup V_{q_{1}}$, by (Deg), all vertices of $B$ have at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}$. Letting $N:=\tilde{U}_{1}\cap\bigcap_{w\in B}N(w)$, we have $|N|\geq|\tilde{U}_{1}|-(t+1)2\varepsilon n$.
First assume that $N$ is $K_{t-1,t}$-free and thus $e_{G}(N)=O(n^{2-1/(t-1)})$. Note that, since $|\tilde{U}_{1}\setminus N|\leq(t+1)2\varepsilon n$, the edges in $\tilde{U}_{1}$ incident to $\tilde{U}_{1}\setminus N$ can be split into two bipartite $K_{t,t}$-free graphs each with one part of size at most $(t+1)2\varepsilon n$ and the other part of size at most $n$. Thus, the number of such edges is $O(\varepsilon n^{2-1/t})$. This gives $f_{1}=O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})$, contradicting (4.12).
We thus assume $N$ contains a copy of $K_{t-1,t}$. Together with $B$, they form a copy of $K_{3}{(t)}$ in $G[\tilde{U}_{1}\cup\tilde{U}_{i}]$ and we denote its vertex set by $B^{\prime}$. By Claim 4.5, there exists $l\notin\{1,i\}$ such that $B^{\prime}$ intersects $V_{q_{l}}$ and $Z_{l}$. By Claim 4.8, $\tilde{U}_{1}\cap U_{l}=\emptyset$, so $B^{\prime}\cap Z_{l}=B\cap Z_{l}$ and $B$ indeed intersects $Z_{l}$. Since $\tilde{U}_{i}\cap Z_{l}\supseteq B\cap Z_{l}\neq\emptyset$, we infer that $|W_{l}|\geq(1-2\varepsilon)n$ from $(Q3)$, which implies that $q_{l}\neq q_{1}$ because of (4.13). It follows that $W_{1}\cap V_{q_{l}}=\emptyset$ and thus $B\cap V_{q_{l}}=B^{\prime}\cap V_{q_{l}}\neq\emptyset$, as desired.
For Part (2), let $A_{i}:=Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ and $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{Z}_{i}\setminus Z_{1}$ or in $A_{i}$. Then, by the first part of the claim, there exists $l\in[r]\setminus\{i\}$ such that $B$ intersects $V_{q_{l}}$ and $Z_{l}$. This is impossible if $B\subseteq A_{i}$ because $A_{i}\cap Z_{l}=\emptyset$ for any $l\notin\{1,i\}$, and also impossible if $B\subseteq\tilde{Z}_{i}\setminus Z_{1}$ because in which case $B\cap W=\emptyset$ and thus $B\cap V_{q_{l}}=\emptyset$ for any $l\notin\{1,i\}$. ∎
The following claim shows a clean structure for the $\tilde{U}_{i}$ such that $W_{i}^{\prime}$ is not too small.
Claim 4.10.
For $i\in[2,r]$ such that $|W_{i}|\geq 2\varepsilon n$, we have $\tilde{U}_{i}\subseteq U_{i}\cup V_{q_{i}}$.
Proof.
Suppose instead, for some $i_{0}\in[2,r]$ with $|W_{i_{0}}|\geq 2\varepsilon n$, there exists $v\in\tilde{U}_{i_{0}}\setminus(U_{i_{0}}\cup V_{q_{i_{0}}})$. By $(P4)$ and the fact that $v\in\tilde{U}_{i_{0}}\setminus U_{i_{0}}$, we infer that $d(v,Z_{j})\geq\varepsilon n$ for all $j\neq{i_{0}}$. Then, by Claim 4.7, we have $d(v,U_{i_{0}})<\varepsilon n$. Consequently, $d(v,W_{i_{0}}^{\prime}\cap W_{i_{0}})<\varepsilon n$, namely, $v$ has at least $2\varepsilon n-2\gamma n-\varepsilon n\geq(1/2)\varepsilon n$ non-neighbors in $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ (in $G$). Note that $v$ is adjacent to all the vertices of $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ in $T$. Since $G[W^{\prime}_{i_{0}}]\subseteq T[W^{\prime}_{i_{0}}]$, we infer that
$e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})\leq e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(1/2)\varepsilon n.$
Since $e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$ and $\alpha_{i_{0}}=|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i}^{\prime}|+e_{T}(W_{i_{0}}^{\prime})+e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$, we have
$e_{G}(\tilde{U}_{i_{0}}\setminus Z_{i_{0}})$ $=e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})$
$\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(1/2)\varepsilon n$
$\leq\alpha_{i_{0}}-(1/2)\varepsilon n.$ (4.17)
Combining (4.5) and (4.17) gives
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z_{i})\leq\sum_{i\in[r]}\alpha_{i}-(1/2)\varepsilon n$ (4.18)
Recall that $f_{i}\leq(t-1)|\tilde{U}_{i}\setminus Z_{i}|$ ($i\geq 2$) by (4.11), $|W^{\prime}_{1}|\geq(1-\gamma)n$ by (4.16), and $|Z_{i}^{i}|\geq(1-\sqrt{\gamma})n$ by $(P3)$. Therefore, as $\sum_{i=1}^{r}|\tilde{U}_{i}\setminus Z_{i}|\leq(b-1)n+r\sqrt{\gamma}n+{\gamma}n$, we obtain
$\sum_{i=2}^{r}f_{i}\leq(t-1)\left((b-1)n+r\sqrt{\gamma}n+{\gamma}n\right)\leq(t-1)(b-1)n+\sqrt[3]{\gamma}n.$ (4.19)
Since $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, (4.7) becomes $e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i})-\alpha_{i})$. Recall that $\sum_{i=1}^{r}\beta_{i}\leq 2r^{2}C_{0}\sqrt{\gamma}n$ by (4.8). Together with (4.18) and (4.19), we derive that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+\sqrt[3]{\gamma}n+2r^{2}C_{0}\sqrt{\gamma}n-\varepsilon n/2<g(n,r,k,t),$
as $\gamma\ll\varepsilon$. This is a contradiction. ∎
Let $L_{1}\cup L_{2}\cup L_{3}$ be a partition of $[2,r]$ such that $i\in L_{1}$ if and only if $|\tilde{Z}_{i}|<n$, $i\in L_{2}$ if and only if $|\tilde{Z}_{i}|=n$, and $i\in L_{3}$ if and only if $|\tilde{Z}_{i}|>n$. The following properties hold for $L_{1}$, $L_{2}$ and $L_{3}$.
  1. $(R1)$
    If $i\in L_{1}$, then $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$, we have $i\in L$ and, by Claim 4.10, $\tilde{Z}_{i}=Z_{i}^{i}\subsetneq V_{i}$ and $W_{i}^{\prime}\subseteq V_{q_{i}}$.
  2. $(R2)$
    If $i\in L_{2}$, then $\tilde{Z}_{i}=Z_{i}^{i}=V_{i}$. Indeed, otherwise $\tilde{Z}_{i}\neq Z_{i}^{i}$, then $|Z_{i}^{i}|<n$ and $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$ and Claim 4.10, we have $\tilde{Z}_{i}=Z_{i}^{i}$, a contradiction.
  3. $(R3)$
    If $i\in L_{3}$, then $\tilde{Z}_{i}\not\subseteq Z_{i}$ (otherwise $|\tilde{Z}_{i}|\leq n$). By Claim 4.10, we have $|W_{i}|<2\varepsilon n$, which implies that $Z_{i}^{j}=\emptyset$ for $j\neq i$ by $(Q2)$. Thus, $Z_{i}^{i}=Z_{i}=V_{i}\subsetneq\tilde{Z}_{i}$.
Now we derive our final bound on $e(G)$. Write $z_{i}:=|\tilde{Z}_{i}|$ and $w_{i}:=|W_{i}^{\prime}|$ for $i\in[r]$.
By Claim 4.3 and the fact that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, we have
$e(G)=e(G^{\prime})+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}\left(\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i})\right).$
Moreover, as $a=1$, (4.5) becomes $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})\leq\alpha_{i}$. It follows that $e_{G}(\tilde{U}_{i})=f_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i})\leq f_{i}+\alpha_{i}$. For $i\in\{1\}\cup L_{1}\cup L_{2}$, we simply use $f_{i}$ as the upper bound and thus we get
[…]
Additional work is needed for $i\in L_{3}$. We let $\lambda=\max\{0,|L_{1}|-|L_{3}|\}$ and for $i\in L_{3}$, let $\lambda_{i}$ be the number of indices $j\in L_{1}$ such that $Z_{j}^{i}\neq\emptyset$. By $(R1)$$(R3)$, we know that if $Z_{j}^{i}\neq\emptyset$, then $i\in L_{3}$ and $j\in L_{1}$. This implies that $\lambda_{i}\geq 1$ for every $i\in L_{3}$, and $\sum_{i\in L_{3}}\lambda_{i}\geq|L_{1}|$, yielding that
$\sum_{i\in L_{3}}(\lambda_{i}-1)\geq\lambda.$ (4.20)
Recall that $G[tZ_{i}\setminus Z_{1}^{i}]$ is $K_{1,t}$-free by Claim 4.9 (2). Since $Z_{i}^{i}=Z_{i}$ is an independent set, it follows that $e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})\leq(t-1)|\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})|$. Together with (4.11), this gives
$e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})+e_{G}(Z_{1}^{i},Z_{i})\leq(t-1)|\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})|+(t-1)|Z_{1}^{i}|=(t-1)(z_{i}-n).$
Therefore, for $i\in L_{3}$, writing $\varrho_{i}:=e_{T}(Z_{1}^{i},\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i}))$, we have
$e_{G}(\tilde{Z}_{i})=e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})+e_{G}(Z_{1}^{i},Z_{i})+e_{G}(Z_{1}^{i},\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})\leq(t-1)(z_{i}-n)+\varrho_{i}.$
Moreover, the definition of $\alpha_{i}$ implies that
$e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})-\alpha_{i}$ $\leq e_{T}(W_{i}^{\prime})+|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|-\alpha_{i}=-e_{T}(\tilde{Z}_{i}\setminus Z_{i})$
$=-\varrho_{i}-e_{T}(\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}))$
$\leq-\varrho_{i}-\binom{\lambda_{i}}{2}\leq-\varrho_{i}+1-\lambda_{i}.$
Finally, by (4.11), we have $e_{G}(Z_{i},W_{i}^{\prime})\leq(t-1)w_{i}$ for $i\in L_{3}$. Combining all these inequalities together, we obtain that, for $i\in L_{3}$,
$e_{G}(\tilde{U}_{i})-\alpha_{i}$ $=e_{G}(\tilde{Z}_{i})+e_{G}(Z_{i},W_{i}^{\prime})+e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})-\alpha_{i}$
$\leq(t-1)(z_{i}-n+w_{i})+(1-\lambda_{i}).$
It follows that $\sum_{i\in L_{3}}(e_{G}(\tilde{U}_{i})-\alpha_{i})\leq\sum_{i\in L_{3}}(t-1)(z_{i}-n+w_{i})-\lambda$ by using (4.20).
Using $\sum_{i=2}^{r}(z_{i}-n)=n-z_{1}$ and $\sum_{i=2}^{r}w_{i}=bn-w_{1}$, we derive that
$\sum_{i\in L_{1}}\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}\cup L_{3}}(z_{i}-n+w_{i})=$ $\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}+\sum_{i=2}^{r}(z_{i}-n+w_{i})$
$=$ $\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}+bn-w_{1}+n-z_{1}.$
Therefore,
$\sum_{i=2}^{r}(e_{G}(\tilde{U}_{i})-\alpha_{i})$ $\leq\sum_{i\in L_{1}}(t-1)\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}\cup L_{3}}(t-1)(z_{i}-n+w_{i})-\lambda$
$=(t-1)(bn-w_{1}+n-z_{1})+\sum_{i\in L_{1}}(t-1)\min\{n-w_{i},n-z_{i}\}-\lambda.$
Finally, we work on the $\beta_{i}$’s and recall that $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|)$. For $i\in\{1\}\cup L_{1}\cup L_{2}$, as $\tilde{Z}_{i}\setminus Z_{i}=\emptyset$, $\beta_{i}=0$. For $i\in L_{3}$, as $|W_{i}|<2\varepsilon n$, we have $Z_{i}\setminus\tilde{Z}_{i}=\emptyset$; for any $j\in L\setminus\{i\}$, we have $\tilde{Z}_{j}\setminus Z_{j}=\emptyset$ and $W^{\prime}_{j}\subseteq V_{q_{j}}$ again by Claim 4.10. Hence $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0$ because $|W^{\prime}_{j}|\leq n$. It then follows that (noting that $L\cap L_{3}=\emptyset$)
$\sum_{i\geq 1}\beta_{i}$ $=\sum_{i\in L_{3}}\beta_{i}=\sum_{i\in L_{3}}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)$
$=\sum_{j\in L}\sum_{i\in L_{3}\setminus\{j\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}(n-z_{j})(w_{j}-n).$
Note that $1\in L$ by (4.16) and $(n-z_{1})(w_{1}-n)\leq 0$ by Claim 4.8. Furthermore, since $n-z_{j}=0$ for $j\in L_{2}$, it follows that $\sum_{i\geq 1}\beta_{i}=\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)$.
Recall that $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(z_{1},w_{1})$. By ($E3$), we have $z_{t}(z_{1},w_{1})+(t-1)(n-z_{1}+n-w_{1})\leq z_{t}(n)$. Thus, combining these estimates together, by (4.7), we get
$e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(e_{G}(\tilde{U}_{i})-\alpha_{i}+\beta_{i})\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+y-\lambda,$ (4.21)
where $y:=\sum_{i\in L_{1}}((t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i}))$. For each $i\in L_{1}$, let $y_{i}:=\min\{n-w_{i},n-z_{i}\}$ and $y_{i}^{\prime}:=\max\{n-w_{i},n-z_{i}\}$. Then $y_{i}\leq y_{i}^{\prime}$ and thus,
$(t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i})=y_{i}(t-1-y_{i}^{\prime})\leq\lfloor(t-1)^{2}/4\rfloor\leq 1.$
Since $L_{1}\subseteq L\setminus\{1\}$, we have $|L_{1}|\leq b-1$. Moreover, by Claim 4.10, we have $w_{i}\leq|W_{i}|+|W_{i}^{\prime}\setminus W_{i}|\leq 2\varepsilon n+\gamma n\leq 3\varepsilon n$ for each $i\in L_{3}$. If $|L_{1}\cup L_{2}|\leq b-2$, then
$bn=\sum_{i\in[r]}|W_{i}|\leq n+(b-2)n+(r-b+1)\cdot 3\varepsilon n<bn,$
a contradiction. This implies $|L_{1}\cup L_{2}|\geq b-1$, and $|L_{3}|\leq r-b$. Since $|L_{1}|-\lambda=\min\{|L_{1}|,|L_{3}|\}\leq|L_{3}|\leq r-b$, it follows that $|L_{1}|-\lambda\leq\min\{b-1,r-b\}$. Consequently, as $\lfloor(t-1)^{2}/4\rfloor\leq 1$, we get
$y-\lambda\leq|L_{1}|\lfloor(t-1)^{2}/4\rfloor-\lambda\leq\min\{b-1,r-b\}\lfloor(t-1)^{2}/4\rfloor,$
Together with (4.21), it gives the desired bound $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+\min\{b-1,r-b\}\lfloor(t-1)^{2}/4\rfloor=g(n,r,k,t)$. This completes the proof of Theorem 3 for $k<2r$. ∎
Acknowledgments
The authors would like to thank Junxue Zhang for creating a figure and helpful comments.